### v1

# Detection of Gravitational Anomaly at Low Acceleration from a Highest-quality Sample of 36 Wide Binaries with Accurate 3D Velocities

Abstract
We set out to accurately measure gravity in the low-acceleration range $(10^{-11},10^{-9})$ m s-2 from 3D motions of isolated wide binary stars. Gaia DR3 provides precise measurements of the four sky-plane components of the 3D relative displacement and velocity ($\mathbf{r},\mathbf{v}$) for a wide binary, but not comparably precise line-of-sight (radial) separation and relative velocity $v_{r}$. Based on our new observations and the public databases/publications, we assemble a sample of 36 nearby (distance <150<150pc) wide binaries in the low-acceleration regime with accurate values of $v_{r}$ (uncertainty <100<100 m s-1). Kinematic contaminants such as undetected stellar companions are well under control using various observational diagnostics such as Gaia’s ruwe parameter, the color-magnitude diagram, multi-epoch observations of radial velocities, Speckle interferometric follow-up observations, and requiring Hipparcos-Gaia proper motion consistency. For the parameter $\Gamma\equiv\log_{10}\sqrt{\gamma}$ with $\gamma\equiv G/G_{\rm N}$ (where $G$ is a parameter generalizing Newton’s constant $G_{\rm N}$ in elliptical orbits), we find $\Gamma=0.102_{-0.021}^{+0.023}$, inconsistent with standard gravity at $4.9\sigma$, giving a gravity boost factor of $\gamma=1.600_{-0.141}^{+0.171}$ under the assumption of elliptical orbits and angular momentum conservation. Four wide binaries have 3D relative velocities exceeding their estimated Newtonian escape velocities with $1<v_{\rm obs}/v_{\rm escN}\leq 1.2$. These systems are unlikely to be chance associations and are expected in a nonstandard paradigm such as Milgromian dynamics (MOND). The hypothesis that Newtonian gravity can be extrapolated to the low-acceleration limit is falsified by this independent study with accurate 3D velocities. Future radial velocity monitoring and Speckle interferometric imaging for larger samples will be useful to refine the present result.
[…]
thanks: corresponding author: chae@sejong.ac.kr
kyuhyunchae@gmail.com
I Introduction
Einstein (Einstein, 1916) invented general relativity as a relativistic theory of gravity whose nonrelativistic limit is Poisson’s equation, Newtonian gravitation. For the past century this Newton-Einstein standard gravity has been the basis for the modern cosmological paradigm and the requirement of dark matter from astronomical observations. However, Newtonian gravity applicable in the nonrelativistic regime was based on the empirical laws of planetary motions (discovered by Kepler four centuries ago) in the high-acceleration $g_{\rm N}>10^{-7}$m s-2 regime. Hereafter, $g_{\rm N}$ refers to Newton’s constant ($G_{\rm N}$) times mass over distance squared for the system under consideration, i.e., $g_{\rm N}=G_{\rm N}M_{\rm tot}/r^{2}$ for a binary with total mass $M_{\rm tot}$ and 3D separation $r$.
From an empirical point of view, Poisson’s equation was shown to be broken by Le Verrier’s work as gravity gets sufficiently strong, already in the 19th century even before any relativistic theory existed. While a relativistic theory is successful in the strong-gravity regime, the hypothesis that Poisson’s equation can be extrapolated indefinitely into the low-acceleration limit requires a direct empirical verification. General relativity’s success in strong-gravity regimes does not guarantee its correctness in the low-acceleration limit (perhaps, as the classically perfect theory of Maxwell’s electrodynamics has fundamental limitations). Any piece of scientific truth is eventually determined by experimental/observational facts. The same holds for any hypothetically dominant dark matter component as introduced to force agreement with standard gravity, hence the ongoing worldwide campaigns attempting a direct detection of any such component (see, e.g., the report by Billard et al. 2022 and Section 27 of Navas et al. 2024), without any positive signal up to now (e.g., Carlin et al. 2025).
Milgrom (Milgrom, 1983) first noticed/suggested that galactic rotation curves start to deviate significantly from Newton’s $r^{-2}$ law only around a critical acceleration $a_{0}$ near $10^{-10}$ m s-2. Milgrom’s proposal, known as modified Newtonian dynamics (MOND), posits a modification of standard gravitational dynamics in the low acceleration regime to break the strong equivalence principle (while keeping Einstein’s equivalence principle and Galileo’s universality of free fall) and makes a number of salient predictions for gravitational dynamics in astrophysical systems (Sanders and McGaugh, 2002; Famaey and McGaugh, 2012; Banik and Zhao, 2022; Merritt, 2020), including the idiosyncratic external field effects (EFE; Milgrom, 1983; Chae and Milgrom, 2022) in the internal dynamics of freely falling systems under external fields.
Wide binary stars provide a unique probe to test standard gravity and MOND, as first proposed in Hernandez et al. (2012), with more recent examples such as Scarpa et al. (2017); Pittordis and Sutherland (2018); Banik and Zhao (2018). From a theoretical point of view, binaries of point-like masses are the simplest possible systems to test gravity, and wide binaries can have sufficiently low internal acceleration to probe the ‘dark matter’ regime, while any conceivable effect of hypothetical dark matter is negligible, given the stringent limits on this component locally, coming from observed vertical kinematics of disk stars (e.g., Read 2014). From an observational point of view, these binaries can be found in the solar neighborhood, so that their distances are directly measurable (an unusual advantage in astronomy) and their velocities can be measured with a sufficiently good precision.
The advent of the Gaia Data Release 3 (DR3; Vallenari et al., 2023) database ignited intensive wide binary gravity tests in recent years. Because Gaia DR3 does not generally provide sufficiently precise line-of-sight (radial) velocities, several statistical methods have been developed based only on sky-plane (tangential) velocities $v_{p}$ (for a summary of the methods, see Table 3 of Chae 2025c). Because gravity is sensitive to the 3D relative velocity $\mathbf{v}$ between the pair at the 3D displacement $\mathbf{r}$, use of $v_{p}$ does not allow any meaningful gravitational inference from individual wide binaries whose very long orbital periods preclude anything but essentially instantaneous observations. Thus, statistical methods requiring large numbers of wide binaries have been introduced to deal with the issues of projection effects (of $\mathbf{v}$ onto $v_{p}$), orbital orientation, phase occupancy, and false binaries (i.e. gravitationally-unbound fly-bys or chance associations).
All statistical analyses of binaries (Chae, 2023, 2024b, 2024a; Hernandez, 2023; Hernandez et al., 2024b; Hernandez and Kroupa, 2025; Yoon et al., 2025) covering a sufficiently broad dynamic range to include both high ($g_{\rm N}>10^{-8}$ m s-2) and low ($g_{\rm N}<10^{-9}$ m s-2) acceleration regimes show that the median value of $v_{p}$ (or related quantity) is boosted about 20% in the low-acceleration regime as long as the fraction of hierarchical systems $f_{\rm multi}$ (with stellar multiplicity $\geq 3$) is calibrated/checked using the binaries in the high-acceleration regime, regardless of the sample choices that largely dictate $f_{\rm multi}$. See Hernandez et al. (2024a) for a critical review of some divergent results that are not based on the calibration of $f_{\rm multi}$ and do not include internal validation checks through the presence of a high acceleration Newtonian region.
Although statistical methods with $v_{p}$ have been popular, it is clearly desirable to use directly measured 3D velocity $\mathbf{v}$ for gravity tests because individual systems can be analyzed and understood in greater detail, leading to much more accurate gravity inferences. Recently, Chae (2025c, b) developed a Bayesian 3D modeling methodology to infer probability density functions (PDFs) of gravity in individual systems (individually not highly restrictive broad distributions) and then to statistically consolidate them to derive the effective strength of gravity in a common acceleration regime. Chae (2025c) first carried out an extensive study of 312 wide binaries using a simplified 3D model based on the Gaia DR3 3D velocities, which include relatively less precise radial velocities (RVs), and found a dichotomy such that 125 wide binaries in the strong acceleration regime $g_{\rm N}\gtrsim 10^{-8}$ m s-2 agree well with Newton, while 111 wide binaries in the transition and low-acceleration regimes show a $4.2\sigma$ anomaly. Chae (2025b) carried out a pilot study of 32 wide binaries with a fully general 3D model based on accurate 3D velocities including precise HARPS RVs from Saglia et al. (2025), and found a moderate indication of dichotomy that 8 wide binaries with $g_{\rm N}<10^{-9}$ m s-2 show a gravity boost while 24 wide binaries with $g_{\rm N}>10^{-9}$ m s-2 agree well with Newton.
The key to a reliable inference of gravity with 3D motions of wide binaries is the construction of a sample of pure binaries free of undetected kinematic contaminants, such as unresolved companion stars or resolvable but too faint stars. In this work we employ an unprecedented combination of observational diagnostics to have a full control of potential kinematic contaminants. They include not only well-known diagnostics such as imposing a color-magnitude diagram (CMD) exclusion region and Gaia’s ruwe limit, but also multi-epoch observations of radial velocities over more than several years and detailed Speckle interferometric imaging of the stars of many wide binaries. These diagnostics work together to flag close/unresolved contaminants (within tens of au from the star) and more distant contaminants.
For the first time we carry out Speckle observations of 391 wide binaries selected from recent samples used by two of us (e.g., Hernandez et al. 2024b; Chae 2024b). These observations form part of an ongoing campaign to image wide binaries at the diffraction limit of the 2.1 m telescope at the Observatorio Astronomico Nacional (SPM) using Speckle interferometry. These speckle observations will not only flag certain individual binaries but also give us the measured probability that the stringently selected samples will have resolvable faint companions. On the other hand, we collect as many high-precision radial velocities measured with different telescopes and, more importantly, at different epochs separated by at least several years. The collection includes our new observations of 60 wide binaries with the Las Cumbres Observatory (LCO) Network of Robotic Echelle Spectrographs (NRES) and 6 wide binaries with the GEMINI-North Observatory MAROON-X spectrograph (hereafter MAROON-X). It also includes wide binaries with radial velocities selected from HARPS (Saglia et al., 2025), SDSS4 DR17 APOGEE (hereafter APOGEE), [^1] https://www.sdss4.org/dr17/irspec/radialvelocities/ and Scarpa et al. (2017). In addition, we use the comparison of Hipparcos and Gaia proper motions to have a control in the selection process.
Since many previous studies (e.g., Chae 2024b, a; Yoon et al. 2025; Chae 2025c, b) of wide binaries including the recent 3D analyses already confirm that Newton is verified at least for $g_{\rm N}>10^{-8}$ m s-2, we focus on the low-acceleration regime only. We construct an extremely curated sample of 36 pure wide binaries in the low-acceleration regime that have accurate and precise relative RVs with measurement error <100<100 m s-1. Our new sample is $\approx 4$ times as large as the Saglia et al. (2025) sample of wide binaries having radial velocity measurements in the low-acceleration regime, with comparable data qualities, and thus our inferred constraints on gravity can be expected to have twice the precision than the pilot study by Chae (2025b).
In Section II we briefly describe theoretical motivations and the Bayesian 3D modeling methodology. In Section III, we describe the process of selecting the statistical sample of wide binaries to be used for gravity inference. We present the results on inferred gravity and relevant discussions in Section IV. We give our thoughts on the meanings and implications of the results in Section V and conclude in Section VI with a future outlook. The Python codes used in this work can be found at Chae (2025a), and the observational data for our sample and the Bayesian outputs will be available on Zenodo under an open-source Creative Commons Attribution license. In Appendices A, B, C, and D, we describe the observations or collections of RVs. In Appendix E, we describe the Speckle interferometric observations.
II Theoretical motivation and the 3D modeling methodology
In this work we set out to measure gravity in the low-acceleration regime in the context of distinguishing between standard and nonstandard theories of gravity. Although MOND is not yet an established physical theory of relativistic gravity (not to mention quantum gravity), it is useful to consider nonrelativistic MOND models as modifications of standard gravity in the low-acceleration regime. MOND predicts that the $r^{-2}$ law with Newton’s constant $G_{\rm N}$ will no longer hold in the acceleration regime $g_{\rm N}\lesssim 10^{-9}$m s-2. According to MOND, how gravity will behave depends on the details of the EFE. For a truly isolated system without an external field, the gravity law will gradually switch to a $r^{-1}$ behavior between $10^{-8}\lesssim g_{\rm N}\lesssim 10^{-10}$m s-2. In reality, common dynamical systems such as binary stars and galaxies can be subject to external fields of various strength, and consequently, in the regime of low internal acceleration, gravity is predicted by MOND to become pseudo-Newtonian, showing an $r^{-2}$ behavior with a rescaled gravitational parameter $G(>G_{\rm N})$ that depends on the external field (see, e.g., Chae and Milgrom 2022 for specific numerical examples).
In the case of wide binaries in the solar neighborhood, the external field due to the Galaxy is $\approx 1.8a_{0}$, which is so strong in the context of MOND that the gravity boost factor is expected to be about $G/G_{\rm N}\approx 1.4$ for a test particle (e.g., Banik and Zhao 2018; Chae and Milgrom 2022). Complete numerical solutions for a two-body dynamics under an external field (due to a third body) find a lower boost factor of 1.21.31.2-1.3 (Pflamm-Altenburg, 2025) using the QUMOND formalism (Milgrom, 2010).
To test Newtonian and MOND predictions for wide binaries of the solar neighborhood, it is convenient to approximate instantaneous orbits as elliptical even if global orbits deviate from closed ellipses as in the MOND case (Pflamm-Altenburg, 2025). While this assumption is perfectly valid in testing Newtonian gravity, any result deviating from Newton needs to be interpreted correctly in the context of testing MOND models such as AQUAL (Bekenstein and Milgrom, 1984) and QUMOND (Milgrom, 1983), or any other alternative model of gravity of interest. Regardless of the details of any modified gravity theory, an approach, in which a $G\to\gamma G_{N}$ model is tested, will reveal the presence of a gravitational anomaly if the inferred value of $\gamma\neq 1$.
We use the 3D modeling methodology of Chae (2025c, b) that infers PDFs $p_{i}(\Gamma)$ ($i=1,\cdots,N_{\rm binary}$) of the parameter
$\Gamma\equiv\log_{10}\sqrt{\gamma}\equiv\log_{10}\sqrt{G/G_{\rm N}}$ (1)
for individual systems (along with orbit and orientation parameters: see Figure 1) and then statistically consolidate them through a normalized product of $p_{i}$.
Figure 1: (Adapted from Chae (2025b)) A general 3D geometry of an elliptical orbit. Here $z^{\prime}$ represents the line-of-sight (radial) direction in observer’s frame with the ++ sign indicating the direction pointing to the observer. The relative radial velocity refers to $v_{z^{\prime}}(\equiv v_{r})$. See Chae (2025b) for the definition of all the other parameters.
Here we briefly describe the essential points of the methodology and refer the reader to Chae (2025c, b) for the details. Each binary has the six measured quantities of $\mathbf{r}=\{x^{\prime},y^{\prime},z^{\prime}\}$ and $\mathbf{v}=\{v_{x^{\prime}},v_{y^{\prime}},v_{z^{\prime}}\}$, where $x^{\prime}$ and $y^{\prime}$ are (nearly) exact, $v_{x^{\prime}}$, $v_{y^{\prime}}$, and $v_{z^{\prime}}$ are very precise, and $z^{\prime}$ is usually not so precise. Taking advantage of the fact that $x^{\prime}$ and $y^{\prime}$ are fixed by the data, we have a reduced set of free parameters $\mathbf{\Theta}=\{e,i,\phi_{0},\Delta\phi(\equiv\phi-\phi_{0}),\log_{10}f_{M},\Gamma\}$, where $e$ is the eccentricity with the range $(0,1)$, $i$ is the inclination with the range $(0^{\circ},180^{\circ})$, $\phi_{0}$ is the phase of the periastron with the range $(0^{\circ},360^{\circ})$, and $\Delta\phi$ is the phase relative to $\phi_{0}$ known as the true anomaly with the range $(0^{\circ},360^{\circ})$. [^2] We note that the angle parameter $\theta$ shown in Figure 1, which is related to the argument of the ascending node, can be calculated a a function of the free parameters. Here $f_{M}$ represents the total mass of the binary system normalized by the observational mass. Inclusion of the parameter $f_{M}$ in the Bayesian modeling means that we are allowing a probability distribution of mass (hence including a confidence interval on this parameter for each binary) with its prior set at the observationally inferred value.
The posterior probability of the parameters $p(\mathbf{\Theta})$ is defined by
$\ln p(\mathbf{\Theta})=\ln\mathcal{L}+\sum_{l}\ln f_{\rm pr}(\Theta_{l}),$ (2)
where $\mathcal{L}$ is the likelihood function (connecting the free parameters with $\mathbf{r}$ and $\mathbf{v}$) whose details can be found in Chae (2025b), and $f_{\rm pr}(\Theta_{l})$ ($l=1,\cdots,6$) is the prior probability for the parameter $\Theta_{l}$. The imposed priors are as follows: $f_{\rm pr}(i)$ = $\sin(i)$ (isotropic orientation), $f_{\rm pr}(\phi_{0})$ = uniform (random orientation), and
$f_{\rm pr}(\Delta\phi)=\frac{(1-e^{2})^{3/2}}{2\pi}\frac{1}{[1+e\cos(\Delta\phi)]^{2}},$ (3)
a phase occupancy probability which is inversely proportional to the scalar velocity at any given phase. Equation (3) ensures that each instantaneous motion being observed occurs at a random time during its orbital period. We also impose priors on $e$, $f_{M}$, and $\Gamma$ as follows: $f_{\rm pr}(e)=(1+\alpha)e^{\alpha}$ taking $\alpha=1$ (the thermal probability distribution) as the nominal choice but considering also a full range of possibilities given by $0\leq\alpha\leq 1.3$ (as inferred by Hwang et al. (2022) for wide binaries of the solar neighborhood), a normal probability distribution of $\log_{10}f_{M}$ with $(\mu,\sigma)=(0,0.021)$ (i.e. 5% scatter in the observational total mass), and a uniform distribution of $\Gamma$ in the range $1<\Gamma<1$. We note that only the flat prior is considered for $\Gamma$ when we seek to measure it from the binary dynamics data.
We will mainly focus on measuring gravity at low acceleration by deriving the PDF of $\Gamma$ while treating the rest of the parameters as essentially nuisance parameters. To ensure an unbiased inference of $\Gamma$, it is necessary to use proper priors on the nuisance parameters as given above (in particular, the priors on inclination $i$ and orbit true anomaly $\Delta\phi$). For example, observational identification of a wide binary is blind to inclination or orbit true anomaly, so a sufficiently large sample of wide binaries should follow the distributions expected from randomness. Thus, the posterior PDFs of the nuisance parameters will provide internal self-consistency checks of the results.
As an auxiliary analysis, we will also carry out Bayesian modeling of the observed 3D motions at fixed gravity models with fixed values of $\Gamma$ including $\Gamma=0$ (i.e., the Newtonian case). The primary purpose of this auxiliary analysis will be to estimate the Newtonian escape velocity for each binary at a fixed Newtonian or pseudo-Newtonian gravity model. We will use the estimated Newtonian escape velocities in conjunction with various observational diagnostics to assemble gravitationally-bound pure binary systems, as will be described in the next section. The pilot study by Chae (2025b) based on the benchmark sample by Saglia et al. (2025) will provide a guidance in this regard. In particular, the Saglia et al. (2025) (sub)sample of 8 (or 9) wide binaries in the low-acceleration regime ($10^{-9}$ m s-2) includes one system where the observed 3D velocity exceeds the Newtonian escape velocity by about 12%. That system is, however, extremely unlikely to be a chance association in the 3D space based on various observational diagnostics and statistical properties of the solar neighborhood. It will be interesting to see how our enlarged sample turns out to be.
III Assembling a Carefully Curated Wide Binary Sample with Accurate 3D velocities
To probe gravity in the low-acceleration regime through 3D modeling of wide binaries, we need kinematically uncontaminated pure binaries with accurate and precise values of the 6 components of $\mathbf{r}$ and $\mathbf{v}$ (Figure 1). Since the currently available observation facilities do not permit a sufficiently precise measurement of radial separation $z^{\prime}$, it is unavoidable to work with only 5 precise components along with generally imprecise $z^{\prime}$. Because Gaia DR3 provides sufficiently precise values for the 4 sky-projected quantities $x^{\prime}$, $y^{\prime}$, $v_{x^{\prime}}$, and $v_{y^{\prime}}$, we focus on collecting wide binaries with precise $v_{r}(\equiv v_{z^{\prime}})=-({\rm RV}_{B}-{\rm RV}_{A})$ where ${\rm RV}_{A}$ and ${\rm RV}_{B}$ are, respectively, the RVs of the brighter and fainter components relative to the Sun in the usual sense of the sign (i.e., the positive sign representing the direction of moving away from the Sun). Throughout, the brighter and fainter components are denoted by $A$ and $B$, respectively.
To collect pure wide binaries with accurate and precise $v_{r}$ in the low-acceleration regime, we follow three steps. In the first step, we collect as many as possible wide binaries with relatively precise RVs for both components and carry out Bayesian modeling for all the systems. In the second step, we select wide binaries that have low internal acceleration ($10^{-9}$ m s-2 based on the Bayesian modeling results) and pass the basic observational criteria required for pure binaries. In the final step, we select the clean sample that is most likely to be free from any kinematic contamination based on various observational diagnostics and the Bayesian modeling results.
III.1 Step 1: Construction of a large sample of wide binaries with relatively precise radial velocities
The collection includes our new observations carried out during 2024-2025 as well as public database/publications. New RVs are available for 60 wide binaries from LCO (Appendix A) and for 6 wide binaries from MAROON-X (Appendix B). For the case of LCO observations, RVs were measured independently at two epochs separated by a few months for 18 of them. We also collect wide binaries with precise RVs from public databases and individually published results: 195 wide binaries within 300 pc from APOGEE (Appendix D), 32 wide binaries with HARPS RVs from Saglia et al. (2025), and 24 wide binaries from Scarpa et al. (2017) after excluding obvious chance-alignment and kinematically contaminated cases (see Appendix C). Because some wide binaries are included more than once in various samples, we have 306 unique wide binaries with relatively precise RVs (compared to Gaia DR3 RVs). This combined sample will be our raw or scratch sample.
Figure 2 shows the distribution of the reported nominal uncertainties of $v_{r}$ for the systems in the raw sample. All systems satisfy $\sigma_{v_{r}}<350$ m s-1 by selection, and the majority have $\sigma_{v_{r}}<100$ m s-1. Table 1 lists wide binaries for which independent measurements of $v_{r}$ were made with different instruments at different epochs. Time baselines between different observations considered in this work are summarized in Figure 3. For all but one system (the fourth) two independent values are consistent with each other. The first (obs1) values may be more accurate (and are more precise in most cases), and so only the first values are included in the raw sample. The fourth system (Gaia DR3 5607190344506642432 & 5607189485513198208) with the HARPS value of $v_{r}=-0.080\pm 0.003$ km s-1 from Saglia et al. (2025) will be used despite the inconsistency with the Scarpa value of $v_{r}=-0.441\pm 0.032$ km s-1 from Scarpa et al. (2017), because the HARPS measurements are generally reliable and the HARPS value is consistent with the Gaia DR3333https://gea.esac.esa.int/archive/ value of $v_{r}=-0.290\pm 0.184$ km s-1.
Figure 2: The distribution of nominal uncertainties of $v_{r}(\equiv{\rm RV}_{A}-{\rm RV}_{B})$ for the raw sample of 306 unique wide binaries assembled from various observations and database/publications.
[…]
Table 1: Wide binaries with two measurements of $v_{r}$ from independent observations with different instruments
[…]
Gaia DR3 identifier Gaia DR3 identifier ssaaSky-projected separation between the pair from El-Badry et al. (2021). $v_{r}$bbRelative radial velocity between the pair $v_{r}\equiv{\rm RV}_{A}-{\rm RV}_{B}$ $v_{r}$bbRelative radial velocity between the pair $v_{r}\equiv{\rm RV}_{A}-{\rm RV}_{B}$ obs1/obs2ccTwo independent observations at different epochs. Only the value from obs1 will be used in this work. time baselineddApproximate median time baseline between obs1 and obs2.
Star A Star B obs1 obs2
[kau] [km s-1] [km s-1] [yr]
2776055105362407680 2776054899203977728 4.83 $0.468\pm 0.017$ $0.605\pm 0.150$ HARPS/LCO 3
4940794866807373952 4940794488850252928 4.94 $0.449\pm 0.004$ $0.387\pm 0.069$ HARPS/LCO 3
5060104351007433472 5060105897197110144 9.08 $0.254\pm 0.035$ $0.377\pm 0.024$ HARPS/Scarpa 5
5607190344506642432 5607189485513198208 12.33 $0.080\pm 0.003$ $0.441\pm 0.032$ HARPS/Scarpa 5
3285218186904332288 3285218255623808640 1.42 $0.590\pm 0.017$ $0.572\pm 0.020$ HARPS/Scarpa 5
4249652990144051840 4249652783985617920 2.74 $140\pm 0.004 0.0$$061\pm 0.026$ HARPS/Scarpa 5
2201661297490051968 2201661091331626752 8.37 $018\pm 0.006 0.0$$002\pm 0.110$ MAROON-X/LCO 0.3
1172915990414659328 1172920487244742912 24.68 $043\pm 0.033 0.0$$027\pm 0.077$ APOGEE/LCO 6
1282815063829295360 1282817022334383232 16.77 $0.283\pm 0.041$ $0.170\pm 0.021$ LCO/Scarpa 11
3230677565443833088 3230677874682668672 4.97 $0.076\pm 0.057$ $0.135\pm 0.022$ LCO/Scarpa 11
3550081879381593728 3550084490721711872 7.87 $0.595\pm 0.082$ $0.410\pm 0.029$ LCO/Scarpa 11
Figure 3: Summary of individual (given in parentheses at the top row and the left column) and pairwise time baselines for the various observations of radial velocities used. All given entries represent approximate median values. Hipparcos epoch is for observations of proper motions.
We note that the acquired values of $v_{r}$, except for those from Saglia et al. (2025), do not include corrections for gravitational redshifts or convective flows in stellar atmospheres. Following the recommendation of Saglia et al. (2025), we will quadratically add $40$ m  s-1 to all nominal uncertainties of $v_{r}$ when using them for modeling.
Individual Bayesian 3D modeling is carried out for each binary from the raw sample either with a fixed value of $\Gamma$ (including the Newtonian case) or allowing it to vary within the range $1<\Gamma<1$. These Bayesian outputs will be used in the selection of wide binaries in the low-acceleration regime and in the derivation of the value of $\Gamma$ through statistical consolidation. Since the Bayesian outputs provide orbit solutions, we can estimate various quantities for each system, including the Newtonian gravitational acceleration defined above and the Newtonian escape velocity $v_{\rm escN}\equiv\sqrt{2G_{\rm N}M_{\rm tot}/r}$ where $r$ is estimated from the Bayesian outputs.
Figure 4: This figure shows the color-magnitude diagram of stars in the raw sample of 306 wide binaries using Gaia’s BP-RP color and absolute magnitude $M_{G}$ in Gaia’s $G$ band. All stars in the clean sample are fainter than the color-dependent $M_{G}$ cut line.
III.2 Step 2: Selection of the basic-cut sample of wide binaries in the low-acceleration regime
In the second step, we select a statistical sample of wide binaries in the low-acceleration regime that have sufficiently precise $v_{r}$ and are likely to be free from any kinematic contaminants based on various observational diagnostics available at present. We first require basic selection and quality cuts of the following:
  • Newtonian gravitational acceleration $g_{\rm N}<10^{-9}$m s-2 where $g_{\rm N}$ is derived from Bayesian reconstructed orbits in generalized gravity.
  • Distance of the binary system from the Sun is less than 150 pc.
  • Relative RV $v_{r}$ and relative 3D velocity $v_{\rm obs}=|\mathbf{v}_{\rm obs}|$ have measurement errors less than $100$ m s-1.
  • Both stars have Gaia’s ruwe <1.25<1.25. (This threshold follows Saglia et al. (2025) but a slightly different threshold such as <1.20<1.20 or <1.30<1.30 yields consistent results.)
  • Both stars are below a color-dependent luminosity cut line in the main sequence of the color-magnitude diagram as shown in Figure 4. This threshold helps remove unresolved close binaries, i.e., photometric binaries.
The above selection criteria retain only wide binaries of low internal acceleration that are relatively nearby and thus have highest data qualities. The criteria also help remove cases with potential kinematic contaminants towards the goal of selecting pure binary individuals. Only 75 systems (about 25%) of the raw sample pass the above cuts. This basic-cut sample is listed in Table 2. In the table, we provide the measured relative radial velocity $v_{r}$ along with its observational source. In the table, we also provide the ratio of the relative 3D velocity to the Newtonian escape velocity $v_{\rm obs}/v_{\rm escN}$, and selection merits for the final clean sample to be described below. The distribution of $v_{\rm obs}/v_{\rm escN}$ for the basic-cut sample can be found in Figure 5. It shows that this sample includes cases with >1>1 and even with >1.2>1.2.
[…]
Table 2: Measured $v_{r}$ and Bayesian inferred $v_{\rm obs}/v_{\rm escN}$ and $\Gamma$ in the basic-cut sample
[…]
Star A/B [pc] [km s-1] [km s-1]
1 2642922251741817216/2642922286101533952 56.01 $0.503\pm 0.657$ $0.156\pm 0.040$ $0.33\pm 0.09$ $0.342_{-0.162}^{+0.255}$ APOGEEggHere APOGEE refers to either APOGEE-1 or APOGEE-2(N or S). Note that most RVs used happen to come from APOGEE-2. 4
2 3840226230398314368/3840219358450646656 118.00 $4.099\pm 3.087$ $0.140\pm 0.085$ $0.78\pm 0.44$ $0.698_{-0.145}^{+0.391}$ APOGEE None
3 3861880665230888960/3858878375716469120 131.31 $0.527\pm 0.327$ $0.386\pm 0.051$ $1.16\pm 0.09$ $0.400_{-0.142}^{+0.234}$ APOGEE 4
4 315289980082496000/315289980082496256 82.89 $1.707\pm 2.100$ $0.869\pm 0.061$ $1.17\pm 0.08$ $0.338_{-0.117}^{+0.228}$ APOGEE None
5 2627346157705716352/2627346535662838272 147.70 $0.929\pm 1.605$ $0.145\pm 0.057$ $53\pm 0.08 0.0$$025_{-0.117}^{+0.240}$ APOGEE 4
6 2581807955200716800/2581806619466249728 145.88 N/A $0.011\pm 0.072$ $1.26\pm 0.10$ $0.425_{-0.129}^{+0.225}$ APOGEE None
7 2594556109625136128/2594555491149842432 110.38 $0.805\pm 5.848$ $0.128\pm 0.092$ $0.47\pm 0.25$ $0.331_{-0.260}^{+0.312}$ APOGEE None
8 783007245705158400/783007280064896128 134.55 $0.006\pm 2.106$ $0.437\pm 0.051$ $1.26\pm 0.12$ $0.339_{-0.121}^{+0.232}$ APOGEE None
9 1459718891236328192/1459715764500136192 137.77 $4.574\pm 3.297$ $0.019\pm 0.065$ $65\pm 0.12 0.0$$078_{-0.145}^{+0.262}$ APOGEE None
10 1476646250703083264/1476646285062821888 139.05 $0.142\pm 3.252$ $0.172\pm 0.092$ $0.55\pm 0.11$ $0.132_{-0.086}^{+0.263}$ APOGEE None
11 1534089651581519488/1534086421765263360 120.28 $1.159\pm 2.037$ $0.433\pm 0.081$ $1.45\pm 0.22$ $0.402_{-0.154}^{+0.240}$ APOGEE None
12 1555391177542038400/1555388222603962752 105.41 $0.204\pm 3.896$ $0.119\pm 0.079$ $0.75\pm 0.06$ $0.173_{-0.123}^{+0.266}$ APOGEE None
13 1580778076392231424/1580778213830765824 138.25 $1.838\pm 2.721$ $0.142\pm 0.085$ $84\pm 0.13 0.0$$062_{-0.103}^{+0.245}$ APOGEE None
14 1331462302965590400/1331462302965590144 117.89 $5.305\pm 2.730$ $0.986\pm 0.082$ $1.21\pm 0.09$ $0.355_{-0.115}^{+0.244}$ APOGEE None
15 1403853152105526272/1403852430551017984 54.69 $3.987\pm 2.311$ $0.015\pm 0.087$ $0.44\pm 0.10$ $0.149_{-0.081}^{+0.185}$ APOGEE None
16 1240752559313435520/1240752559313476224 108.20 $1.124\pm 3.164$ $0.004\pm 0.072$ $66\pm 0.07 0.0$$061_{-0.103}^{+0.261}$ APOGEE 4
17 5262754514488149632/5262754273969986688 138.99 $309\pm 0.224 0.0$$049\pm 0.048$ $61\pm 0.07 0.0$$059_{-0.119}^{+0.260}$ APOGEE None
18 4757024829820200704/4757026135488693760 89.24 $1.003\pm 0.526$ $0.633\pm 0.084$ $1.27\pm 0.06$ $0.563_{-0.104}^{+0.186}$ APOGEE None
19 690460424271103360/690272545221668096 109.64 $0.074\pm 1.335$ $0.126\pm 0.062$ $60\pm 0.12 0.0$$019_{-0.129}^{+0.263}$ APOGEE None
20 694284491350980864/694284628789933568 37.84 $0.259\pm 0.552$ $0.276\pm 0.100$ $66\pm 0.12 0.0$$089_{-0.141}^{+0.249}$ APOGEE 4
21 4763619318293079168/4763618910272182400 77.02 $2.327\pm 2.076$ $0.180\pm 0.075$ $51\pm 0.11 0.0$$085_{-0.136}^{+0.255}$ APOGEE None
22 1444305829863259648/1444305937237949184 136.58 $0.282\pm 1.320$ $0.349\pm 0.038$ $61\pm 0.10 0.0$$008_{-0.154}^{+0.267}$ APOGEE None
23 972821779152541824/972815560039899264 113.05 $0.828\pm 1.575$ $0.160\pm 0.048$ $65\pm 0.15 0.0$$063_{-0.159}^{+0.277}$ APOGEE 4
24 1172915990414659328/1172920487244742912 53.42 $006\pm 0.360 0.0$$043\pm 0.033$ $0.39\pm 0.05$ $0.172_{-0.113}^{+0.234}$ APOGEE 1, 2, 4, 5
25 5305981470567619456/5305981745445719680 93.79 $0.132\pm 0.257$ $0.337\pm 0.086$ $0.93\pm 0.07$ $0.243_{-0.117}^{+0.231}$ LCO 2
26 3749791158495959552/3749791055416743552 69.89 $0.401\pm 0.247$ $0.383\pm 0.073$ $54\pm 0.11 0.0$$064_{-0.148}^{+0.260}$ LCO 2
27 1282815063829295360/1282817022334383232 55.00 $0.327\pm 0.174$ $0.283\pm 0.041$ $0.75\pm 0.11$ $0.120_{-0.132}^{+0.252}$ LCO 1, 2, 5, 6
28 3793107930900527616/3793106419072038272 86.00 $0.748\pm 0.266$ $0.260\pm 0.094$ $0.62\pm 0.22$ $0.224_{-0.239}^{+0.304}$ LCO None
29 1258410612976538368/1258410750415492864 44.12 $0.192\pm 0.189$ $0.632\pm 0.082$ $0.88\pm 0.08$ $0.316_{-0.139}^{+0.231}$ LCO None
30 3170300942420466176/3170394607068638336 121.69 $750\pm 0.320 0.0$$055\pm 0.058$ $47\pm 0.08 0.0$$087_{-0.120}^{+0.256}$ LCO 2
31 5651775953326498688/5651752515687994368 55.09 $0.131\pm 0.245$ $0.119\pm 0.081$ $0.36\pm 0.08$ $0.205_{-0.123}^{+0.241}$ LCO 2
32 5741345125461459200/5741344919302959360 105.03 $0.631\pm 0.254$ $0.327\pm 0.042$ $0.94\pm 0.07$ $0.240_{-0.117}^{+0.248}$ LCO 1, 2
33 3116331104937277952/3116324881524878208 73.86 $0.181\pm 0.245$ $0.345\pm 0.060$ $1.20\pm 0.16$ $0.313_{-0.130}^{+0.239}$ LCO 2
34 3230677565443833088/3230677874682668672 63.74 $328\pm 0.189 0.0$$076\pm 0.057$ $0.37\pm 0.08$ $0.225_{-0.112}^{+0.307}$ LCO 2, 5
35 5185524920830592128/5185536774940328448 38.40 $0.103\pm 0.301$ $0.276\pm 0.086$ $0.52\pm 0.17$ $0.164_{-0.167}^{+0.268}$ LCO None
36 1967283042361454848/1967282939282261120 82.84 $0.514\pm 0.236$ $0.561\pm 0.082$ $0.96\pm 0.11$ $0.226_{-0.125}^{+0.223}$ LCO 1
37 6570796871887419648/6570797524722448896 78.18 $0.259\pm 0.215$ $0.105\pm 0.069$ $0.33\pm 0.10$ $0.299_{-0.122}^{+0.223}$ LCO None
38 3158926322836178816/3158878734598549376 78.03 $065\pm 0.606 0.0$$023\pm 0.066$ $0.43\pm 0.07$ $0.117_{-0.111}^{+0.230}$ LCO 1
39 1938247517245907456/1938247654684985216 59.47 $067\pm 0.187 0.0$$055\pm 0.046$ $0.88\pm 0.04$ $0.247_{-0.124}^{+0.224}$ LCO 6
40 1815165535636339072/1815165883534980992 91.65 $0.075\pm 0.274$ $0.853\pm 0.047$ $1.47\pm 0.09$ $0.408_{-0.103}^{+0.215}$ LCO None
41 1719835231806217472/1719835407900844544 99.62 $0.192\pm 0.361$ $0.297\pm 0.060$ $63\pm 0.09 0.0$$040_{-0.130}^{+0.249}$ LCO None
42 3550081879381593728/3550084490721711872 33.55 $0.692\pm 0.233$ $0.595\pm 0.082$ $1.09\pm 0.14$ $0.277_{-0.131}^{+0.239}$ LCO 5
43 1760471948915107200/1760477618271932672 75.42 $0.011\pm 0.212$ $0.947\pm 0.047$ $1.09\pm 0.06$ $0.338_{-0.124}^{+0.238}$ LCO None
44 4763357363943383808/4763357260864168320 90.68 $0.194\pm 0.276$ $0.283\pm 0.092$ $0.93\pm 0.06$ $0.206_{-0.098}^{+0.268}$ LCO None
45 1003223614961194752/1003223584897948160 103.02 $0.054\pm 0.315$ $0.430\pm 0.077$ $0.74\pm 0.09$ $0.123_{-0.123}^{+0.247}$ LCO 1
46 1142787168495168000/1142786996696476288 105.48 $0.101\pm 0.349$ $0.317\pm 0.075$ $54\pm 0.10 0.0$$027_{-0.138}^{+0.267}$ LCO None
47 5038962632187714432/5038962219872444160 115.58 $0.142\pm 0.263$ $0.212\pm 0.088$ $0.40\pm 0.17$ $0.352_{-0.179}^{+0.293}$ LCO None
48 2932313231147240960/2932313196787509376 137.36 $023\pm 0.231 0.0$$073\pm 0.085$ $49\pm 0.09 0.0$$056_{-0.120}^{+0.272}$ LCO None
49 3487243037508315648/3487237024554098560 79.73 $0.037\pm 0.214$ $0.170\pm 0.060$ $52\pm 0.11 0.0$$085_{-0.137}^{+0.248}$ LCO None
50 4599984642025088128/4599984504586131456 69.75 $215\pm 0.269 0.0$$086\pm 0.062$ $51\pm 0.09 0.0$$023_{-0.128}^{+0.253}$ LCO 1
51 4430185068482324864/4430185034123000960 100.24 $108\pm 0.276 0.0$$025\pm 0.072$ $0.48\pm 0.10$ $0.100_{-0.106}^{+0.268}$ LCO 1
52 5899243585161684864/5899244375435693952 80.48 $0.904\pm 0.217$ $0.566\pm 0.056$ $0.85\pm 0.09$ $0.191_{-0.133}^{+0.258}$ LCO None
53 2776055105362407680/2776054899203977728 108.61 $0.596\pm 0.286$ $0.468\pm 0.017$ $0.93\pm 0.06$ $0.314_{-0.154}^{+0.254}$ HARPS 3, 5
54 4940794866807373952/4940794488850252928 116.26 $0.354\pm 0.215$ $0.449\pm 0.004$ $0.90\pm 0.02$ $0.222_{-0.108}^{+0.233}$ HARPS 2, 3, 5
55 4722135642226902656/4722111590409480064 12.04 $0.505\pm 0.177$ $0.464\pm 0.028$ $70\pm 0.04 0.0$$079_{-0.108}^{+0.254}$ HARPS 3
56 5060104351007433472/5060105897197110144 35.90 $0.421\pm 0.169$ $0.254\pm 0.035$ $65\pm 0.12 0.0$$031_{-0.146}^{+0.247}$ HARPS 3, 5
57 4666907551119833984/4666907516760096512 55.00 $0.319\pm 0.179$ $0.197\pm 0.013$ $0.45\pm 0.08$ $0.139_{-0.130}^{+0.272}$ HARPS 3
58 5607190344506642432/5607189485513198208 38.08 $284\pm 0.188 0.0$$080\pm 0.003$ $0.35\pm 0.04$ $0.299_{-0.162}^{+0.254}$ HARPS 3, 5
59 4235732073427592704/4235731867269159680 100.27 $1.092\pm 0.239$ $0.861\pm 0.015$ $1.12\pm 0.03$ $0.339_{-0.113}^{+0.246}$ HARPS 3
60 6458951765971500672/6458952345790198144 44.30 $344\pm 0.175 0.0$$064\pm 0.012$ $0.45\pm 0.12$ $0.125_{-0.133}^{+0.255}$ HARPS 3
61 2201661297490051968/2201661091331626752 107.50 $347\pm 0.367 0.0$$018\pm 0.006$ $62\pm 0.07 0.0$$056_{-0.116}^{+0.244}$ MAROON-X 5
62 1871558941576158464/1871559697490418816 92.44 $426\pm 0.438 0.0$$024\pm 0.009$ $0.35\pm 0.07$ $0.243_{-0.114}^{+0.223}$ MAROON-X 1
63 1942384773344557184/1942384872124424832 52.00 $0.648\pm 0.312$ $0.266\pm 0.084$ $0.66\pm 0.15$ $0.114_{-0.050}^{+0.056}$ MAROON-X None
64 1762461893163118464/1762461309047562368 121.93 $0.535\pm 0.347$ $0.220\pm 0.004$ $55\pm 0.05 0.0$$012_{-0.127}^{+0.262}$ MAROON-X None
65 4230699363889120128/4230699329529382400 79.48 $0.018\pm 0.540$ $0.278\pm 0.022$ $79\pm 0.07 0.0$$055_{-0.102}^{+0.244}$ MAROON-X 1
66 3400292798990117888/3394298532176344960 14.54 $0.332\pm 0.186$ $0.253\pm 0.052$ $61\pm 0.11 0.0$$044_{-0.171}^{+0.265}$ Scarpa 6
67 3975129194660883328/3975223065466473216 39.41 $0.164\pm 0.242$ $0.141\pm 0.021$ $0.25\pm 0.06$ $0.361_{-0.130}^{+0.251}$ Scarpa None
68 6193279279612173952/6193280031230266752 29.90 $1.280\pm 0.193$ $1.014\pm 0.025$ $1.67\pm 0.06$ $0.519_{-0.103}^{+0.187}$ Scarpa None
69 1440518669436791296/1440425863783337856 45.59 $0.494\pm 0.183$ $0.442\pm 0.034$ $0.78\pm 0.07$ $0.120_{-0.116}^{+0.247}$ Scarpa 6
70 10584899657116672/10608573516849536 47.06 $0.676\pm 0.195$ $0.856\pm 0.027$ $1.27\pm 0.04$ $0.399_{-0.105}^{+0.213}$ Scarpa None
71 3359808231100381312/3359820016490648576 47.63 $0.995\pm 0.208$ $0.791\pm 0.017$ $1.11\pm 0.01$ $0.711_{-0.116}^{+0.131}$ Scarpa None
72 692119656035933568/692120029700390912 48.74 $0.404\pm 0.189$ $0.292\pm 0.015$ $0.69\pm 0.09$ $0.233_{-0.209}^{+0.320}$ Scarpa None
73 1019361632454363904/1019174509319377536 29.18 $1.115\pm 0.181$ $1.337\pm 0.016$ $1.52\pm 0.03$ $0.558_{-0.100}^{+0.175}$ Scarpa None
74 777967084390189696/777967702865481344 45.87 $1.273\pm 0.201$ $1.353\pm 0.015$ $1.39\pm 0.02$ $0.586_{-0.095}^{+0.175}$ Scarpa None
75 2882262637207289216/2882262529831237120 40.92 $0.668\pm 0.181$ $0.747\pm 0.021$ $0.92\pm 0.03$ $0.218_{-0.105}^{+0.223}$ Scarpa 6
Figure 5: The distribution of $v_{\rm obs}/v_{\rm escN}$ is shown for the basic-cut sample of 75 wide binaries that may include kinematically contaminated cases. Here $v_{\rm obs}$ is the magnitude of the observed relative 3D velocity between the pair while $v_{\rm escN}$ is the Newtonian escape velocity predicted by 3D elliptical orbit modeling with a fixed value of $G=G_{\rm N}(\Gamma=0)$ or $1.4G_{\rm N}(\Gamma\approx 0.07)$, which encompass the current range of the effective gravitational constant for Newtonian and MOND models. Both the distribution of the individual median values and that of random values from the Bayesian MCMC samples are shown. The numbers indicated are from the median distributions. Pseudo-Newtonian (MOND) approximate bound of $1.2$ is based on numerical MOND solutions (Chae and Milgrom, 2022) in the regime with acceleration $a_{0}$ for a test particle assuming one-particle equivalent description.
Although the basic cut is already strict to the point that many true pure binaries may have been excluded, the sample may still include contaminated or unreliable cases as Figure 5 indicates. For example, a star with a faint close companion can still have ruwe <1.25<1.25. In addition, even for a true pure binary, a reported value of $v_{r}$ with a small measurement error generally requires independent confirmation due to any possibilities of unknown systematic errors.
III.3 Step 3: Selection of the clean sample of wide binaries in the low-acceleration regime
The driving philosophy of our sample selection is to maximally control systematic uncertainties even if statistical power is significantly sacrificed. Therefore, we require that each binary satisfy at least one of the following additional criteria to remove potentially contaminated or unreliable systems:
  1. (1)
    Speckle interferometric observations of the fields around both stars were carried out and no objects were detected. In addition, the measured value of $v_{r}$ (whatever its source is) is consistent with the Gaia DR3 value within $2\sigma$.
  2. (2)
    Both stars were observed at two epochs with the LCO and two independent values of $v_{r}$ are consistent with each other within $2\sigma$.
  3. (3)
    $v_{r}$ values come from the HARPS RVs of the Saglia et al. (2025) sample.
  4. (4)
    $v_{r}$ values come from APOGEE, and value(s) of VSCATTER (scatter from multiple visits in an observation) is(are) available and its (maximum) value is less than $100$ m s-1.
  5. (5)
    Two independent values of $v_{r}$ are available from two different sources/telescopes and are consistent with each other (see Table 1).
  6. (6)
    Unusually precise Gaia DR3 values (with error <190<190 m s-1) are available and agree within $100$ m s-1 with the values from other sources.
Some systems satisfy multiple selection merits. In Table 2, we indicate the selection merit(s) for each binary.
We note that the targets of Speckle observations were selected from the samples recently defined and used by two of us (e.g. Hernandez et al. 2024b; Chae 2024b). These samples were selected with a number of indirect criteria to exclude apparent binaries hosting unresolved close companion star(s). The Speckle observations (Appendix E) show that only 5% of the target systems potentially have additional faint star(s), confirming that the employed criteria were efficient in selecting samples of pure binaries in a statistical sense. By comparison, in random samples of local stars of comparable mass, Speckle studies generally identify about 40% unresolved binaries.
Only 37 systems satisfy at least one of the requirements listed above. We manually check all objects found in the 3D space around each binary using the Gaia DR3 archive, and find that in one pair (Gaia DR3 1762461893163118464 & 1762461309047562368) there is a third star that may perturb the system. So we exclude this system and are left with 36 systems, which will define our clean sample that also passes additional tests as will be shown in the following two subsections. Aside from the membership inclusion criteria mentioned above, no system having Speckle observations identifying any extra component (some of which might just be background sources) was included in the final clean sample. The properties of the clean sample will be described in Section 3.4.
III.3.1 Additional check: Testing Gaia-Hipparcos proper motion consistency over 25 years
Figure 6: This figure compares sky-projected transverse velocities derived from the relative proper motions (PMs) between the pair reported in Gaia DR3 and Hipparcos for 41 wide binaries selected from the raw sample. The three systems with the $5\sigma$ discrepancies between the two observations show abnormal properties in the posterior PDFs of the parameters from Bayesian modeling (see the text).
We check whether the Gaia DR3 proper motions (PMs) of the stars in the clean sample are consistent with the Hipparcos PMs measured about 25 years earlier than Gaia. Although the Hipparcos precision is poorer by an order-of-magnitude and not all binaries in the Gaia DR3 were observed by Hipparcos, the comparison between the two can provide an additional test. We only consider the relative PM between the pair for this test because it is free from issues of absolute calibrations between the two observations (Brandt, 2018) and what is relevant for internal dynamics is the relative velocity. We find that 41 binaries from the raw sample (including 13 that satisfy the above selection merits) have the Hipparcos PMs for both stars.
Figure 6 shows the comparison between the two observations in terms of the sky-projected relative velocity components $v_{x^{\prime}}$ and $v_{y^{\prime}}$ derived from the PMs for 41 systems from the raw sample including 13 systems belonging to the clean sample. First of all, all 13 systems from the clean sample satisfy the Gaia-Hipparcos consistency at least within $3\sigma$, while there are cases that do not satisfy such a consistency.
Three binaries show $5\sigma$ discrepancy between Gaia DR3 and Hipparcos, strongly indicating that these binaries have been undergoing kinematic perturbations during the time baseline of $\approx 25$ years (see Figure 3) between the two observations. If this is true, the Bayesian modeling results for these systems are very likely to display abnormal properties in the posterior PDFs of the parameters beyond the ranges predicted by any viable gravity models including Newton and MOND. Indeed, we find that the posterior PDFs for the systems return abnormally large values of $\Gamma>0.5$ and $v_{\rm obs}/v_{\rm escN}>1.3$. Interestingly, all these systems are from the Scarpa sample (Gaia DR3 777967084390189696 & 777967702865481344, Gaia DR3 6193279279612173952 & 6193280031230266752, and Gaia DR3 1754191435419155456 & 1754191229260708736: the former two are included in the basic-cut sample of Table 2 while the latter is not), and none of them satisfy any of the selection merits for the clean sample and were therefore excluded. We note also that two of these systems fail the velocity threshold introduced by Chae (2024b) (see Appendix C).
While the $5\sigma$ discrepancy between Gaia DR3 and Hipparcos PMs appears to be a decisive indicator of kinematic contaminants, the interpretation of the less strong discrepancy $3\sigma$ (but not $5\sigma$) is unclear. Three systems exhibit such discrepancies but no abnormalities in the PDF of $\Gamma$ or the inferred value of $v_{\rm obs}/v_{\rm escN}$. This may indicate that a $3\sigma$-threshold in the Gaia-Hipparcos comparison is not effective or reliable. However, the detail on the threshold appears to be irrelevant for the present clean sample because none of the 13 systems from the clean sample show a $3\sigma$ discrepancy.
The remarkable agreement between the Gaia-Hipparcos $5\sigma$ discrepancy and the Bayesian-inferred large value of $v_{\rm obs}/v_{\rm escN}\gtrsim 1.3$ in identifying kinematically contaminated systems suggests that the PDF of $\Gamma$ (as was already discussed in Chae (2025c)) or the quantity $v_{\rm obs}/v_{\rm escN}$ may be used when the direct observational test such as the Gaia-Hipparcos PM test is not available. As will be shown below, all systems from the clean sample have $v_{\rm obs}/v_{\rm escN}\lesssim 1.2$ in concordance with the Gaia-Hipparcos test results.
III.3.2 Additional check: Comparing metallicities of the two components
Figure 7: The left panel shows APOGEE and HARPS measurements of iron abundances [Fe/H] for the stars of 25 wide binaries from the basic-cut sample including 13 from the clean sample. The brighter ($A$) and the fainter ($B$) components of binaries show a negligible difference: ${\rm[Fe/H]}_{B}-{\rm[Fe/H]}_{A}=-0.01\pm 0.07$. Two systems (marked by encircled points) gravitationally unbound by Newtonian gravity ($v_{\rm obs}>v_{\rm escN}$) also have nearly equal values of [Fe/H] for the two components. The middel panel shows Gaia DR3 metallicities [M/H] (Andrae et al., 2023) for 29 wide binaries from the clean sample. There is a good correlation between the two components of the binaries. The right panel shows the systematic difference between the two components. The systematic difference is largely due to the binaries marked by red five-pointed stars in which component A is brighter (more massive) than component B, which is consistent with the known systematic for Gaia’s [M/H] (Andrae et al., 2023). See the text for the details.
While the formation and long-term survival mechanisms of gravitationally-bound wide binaries are not well understood, wide binary stars are thought to be formed in a common interstellar gas cloud, through mechanisms such as dissolution of star clusters (e.g.,Kouwenhoven et al. 2010; Moeckel and Clarke 2011, dynamical unfolding of triples (Reipurth and Mikkola, 2012) ), formation from adjacent cores (Tokovinin, 2017), gravitational capture (Rozner and Perets, 2023)), and star formation in the turbulent interstellar medium (Xu et al., 2023). If the two main-sequence stars of a wide binary are born together from a common cloud, they will have identical or similar chemical compositions regardless of their ages or formation mechanisms. Observational studies generally find that metallicities of the two stars match well (e.g., Andrews et al. 2019; Hawkins et al. 2020; Nelson et al. 2021; Lim et al. 2024).
Thus, we seek to check whether our wide binaries satisfy the expected consistency of metallicities between the two components. In general, high-SNR and high-resolution spectra are needed to measure metallicities of stars, but we collect metallicities from the public databases/publications for as many stars of our sample as possible. Figure 7 shows the metallicities collected from APOGEE, [^4] https://www.sdss4.org/dr17/irspec/abundances/ HARPS (Saglia et al., 2025), and Gaia DR3 (Andrae et al., 2023). The former two provide accurate and precise results based on high-resolution spectra while the latter provides much less precise results.
The left panel of Figure 7 compares APOGEE or HARPS iron abundances [Fe/H] of the the brighter ($A$) and the fainter ($B$) components of 25 (17 APOGEE + 8 HARPS) wide binaries selected from the basic-cut sample that include 13 wide binaries of the clean sample. The metallicity difference ${\rm[Fe/H]}_{B}-{\rm[Fe/H]}_{A}$ is consistent with zero with an rms scatter <0.1<0.1. In particular, the 13 wide binaries of the clean sample have ${\rm[Fe/H]}_{B}-{\rm[Fe/H]}_{A}=-0.01\pm 0.07$. Moreover, the two systems with $v_{\rm obs}/v_{\rm escN}>1$ from the clean sample satisfy well the expected property ${\rm[Fe/H]}_{B}\simeq{\rm[Fe/H]}_{A}$ of true binaries.
The middle panel of Figure 7 shows Gaia DR3 metallicities [M/H] (where M represents all metal elements) for the 29 wide binaries of the clean sample for which this quantity is available for both components. The two metallicities ${\rm[M/H]}_{A}$ and ${\rm[M/H]}_{B}$ are well correlated, but there is a clear systematic shift for those with relatively large mass difference ($M_{B}/M_{A}<0.9$) between the two components while there is no tangible shift for those with similar masses or those having relatively small masses regardless of mass ratio. The selective systematic difference is well consistent with the known systematic of the Gaia DR3 general stellar parameterizer for photometric (GSP-Phot) [M/H] as a function of effective temperature $T_{\rm eff}$. Figure 11 of Andrae et al. (2023) shows that [M/H] for stars with $T_{\rm eff}\lesssim 4500$K have no systematic shift while [M/H] has a systematic shift $\approx-0.4$ - 0.3-0.3 dex for $T_{\rm eff}>4600$K, which corresponds to $M\approx 0.8M_{\odot}$ (Figure 7 of Eker et al. (2015)). Moreover, the same figure shows that the scatter around the median systematic shift is large $\gtrsim 0.3$ dex. Thus, the measured statistic of ${\rm[M/H]}_{B}-{\rm[M/H]}_{A}=0.36\pm 0.17$ dex (i.e., the more massive star has a lower value) is well consistent with random fluctuations.
The above results demonstrate that all available metallicties for our clean sample of wide binaries are consistent with the current understanding of wide binary stars that the two components have identical or similar metallicities.
III.4 Properties of the clean sample
Figure 8: Same as Figure 5 but for the clean sample. There are no cases with $v_{\rm obs}/v_{\rm escN}>1.22$ for this sample.
The clean sample was defined with various selection merits, and additional checks (Sections III.3.1 and III.3.2) using the comparison of Gaia and Hipparcos relative PMs and the comparison of metallicities of the two components have not found any noticeable issue in the clean sample. Thus, all 36 wide binaries from the clean sample will be used to infer gravity. Here we describe key statistical properties of the clean sample.
The distribution of $v_{\rm obs}/v_{\rm escN}$ for the clean sample can be found in Figure 8. Compared with Figure 5, all the systems with $v_{\rm obs}/v_{\rm escN}>1.2$ (or $1.22$ for $G=1.4G_{\rm N}$) and three (or one) systems with $1<v_{\rm obs}/v_{\rm escN}\leq 1.2$ are removed, but four systems with $1<v_{\rm obs}/v_{\rm escN}\lesssim 1.2$ remain. It is interesting to note that although the quality cuts introduced in going from Figure 5 to Figure 8 do not include the $v_{\rm obs}/v_{\rm escN}$ ratio in any way, the above requirements for the stability and reliability of radial velocities naturally remove all cases with $v_{\rm obs}/v_{\rm escN}>1.22$.
Figure 9: The distribution of the magnitude of relative velocity component is shown for the clean sample of wide binaries. The left panel shows the distribution for all wide binaries while the right panel shows the distribution excluding four systems that are not gravitationally bound by Newtonian gravity. In both panels, all three components are statistically consistent with one another, dots with error bars giving the mean velocity values for the 3 velocity components and their confidence intervals.
Figure 9 shows the distribution of the magnitude of three relative velocity components for the clean sample. For a randomly selected sample, the three components are expected to be statistically equivalent if they are measured with comparable precision. Indeed, the three components are consistent with one another, in terms of both the overall shape and the median value.
Figure 10: The relative radial velocity ($v_{r}$) measurements from different observations are compared for the wide binaries of the clean sample. The left panel compares the values collected in this work with the values from the Gaia DR3 database. Relatively precise DR3 values represented with magenta color match excellently with $v_{r}$. The middle panel compares two independent measurements within the LCO network as described in Appendix A.1. Those with $2\sigma$ discrepancy (marked by red dots) are precluded in this work. The right panel compares two independent measurements from different observations/sources separated by more than several years. Black dots are from Table 1 and magenta circles are from the left panel. Seven systems are represented by both black dots and magenta circles, meaning that this panel includes only 25 unique systems (see the text for the details).
In Figure 10, we compare various independently measured values of $v_{r}$. Because the major criterion used to define the clean sample is the availability of multiple values of $v_{r}$ measured at multiple (mostly two) epochs (although the time difference between epochs is quite diverse: see Figure 3), it is interesting to check/test how well independent values match each other. The left panel of Figure 10 compares the values ($v_{r}$) of the clean sample listed in Table 2 with the Gaia DR3 values ($v_{r,\rm{DR3}}$) that happen to be available for all the binaries of the sample. They are consistent with each other up to their measurement errors. The difference $v_{r,\rm{DR3}}-v_{r}$ has a mean of 0.034-0.034 km s-1 and an rms scatter of $0.385$ km s-1. We find that $v_{r,\rm{DR3}}$ has a much larger rms scatter of $0.475$ km s-1 (due to the larger measurement errors) compared with $0.312$ km s-1 for $v_{r}$. The latter is close to the intrinsic scatter. For the subsample with relative precise $v_{r,{\rm DR3}}$ (with error <0.3<0.3 km s-1), $v_{r,\rm{DR3}}$ has a much smaller scatter of $0.377$ km s-1 which is similar to $v_{r}$’s scatter of $0.354$ km s-1 for the same binaries. For this subsample, there is an excellent match between $v_{r}$ and $v_{r,{\rm DR3}}$ that were observed at different epochs separated by 2 - 9 years (Figure 3).
The middle panel of Figure 10 compares two independent values measured by the LCO at two epochs separated by a few months (see Appendix A.1). In 14 of the 18 systems, two values agree well within $2\sigma$. For the rest, one system agrees within $3\sigma$, but the other systems are discrepant by $3\sigma$. All such cases were excluded from the clean sample. In one system (Stars HD 101574 and BD-01 2557, Table 4) there is a large discrepancy between two measurements of $v_{r}$ with different telescopes (within the LCO) due to the discrepant RV values for HD 101574 (see Appendix A.1). It is unclear whether this kind of discrepancy is due to intrinsic variations from kinematic perturbations or issues in measurements. Whatever the case, this possibility of large measurement-to-measurement variation highlights the importance of reproduction and confirmation of $v_{r}$ values.
In the right panel of Figure 10, black dots compare two independent values of $v_{r}$ measured with different telescopes at two epochs separated by 3 - 11 years (except for one system which has a 0.3 year baseline) for 9 wide binaries satisfying the selection merit #(5). The values $v_{r,\rm{obs1}}$ and $v_{r,\rm{obs2}}$, star identifiers, and observation sources of the 9 systems can be found in Table 1 and Table 2. They match well each other within a few times the small nominal errors except for one system, which is not discarded as already mentioned in Section III.1. We note that nominal measurement errors for some systems shown in the panel are extremely small (<10<10m s-1), but realistic errors considering the effects of gravitational redshift and convective flow may be at least a few tens of m s-1 (Saglia et al., 2025). Magenta circles shown in the left panel are reproduced here to show all possible multi-epoch measurements that are reasonably precise. Because seven systems from the subsample represented by magenta circles are also included in the list of Table 1, there are only 25 systems with multi-epoch values. For 21 of them, time baselines are in the range of 3 - 11 years (see Figure 3).
Figure 11: The left panel shows the distribution of masses of the stars to be used in this work. The right panel compares the Gaia DR3 FLAME masses with our masses for those stars whose FLAME masses are available. The two masses agree well with a scatter of only 5%.
Regarding masses of the stars to be used as the observational input, individual estimates are available for wide binaries from the Saglia et al. (2025) sample and the LCO sample (Appendix A). For wide binaries for which individual estimates are not available from stellar modeling, we use the mass-magnitude relation derived by Chae (2023) to estimate the mass of a star. As explicitly shown in Chae (2025c) and Yoon et al. (2025), the Chae (2023) masses agree well with the Gaia DR3 FLAME masses for stars similar to those used in this work whenever the latter are available from the Gaia archive (note that not for all stars FLAME masses are available). Figure 11 shows the distribution of our masses and a comparison with the corresponding FLAME masses for those stars that have the FLAME masses. We note that our masses are statistically consistent with the FLAME masses with an rms scatter of 5% and our Bayesian methodology allows for the uncertainty of mass through the parameter $\log_{10}f_{M}$.
IV Results along with relevant discussions
We now present our inference of gravity at low acceleration ($10^{-9}$ m s-2) through the parameter $\Gamma$ (Equation (1)) by applying the Bayesian methodology (Section II) to the clean sample of wide binaries described in III. First of all, we note that the measured 3D velocity components suffer from minor geometric effects known as perspective effects (Shaya and Olling, 2011; Yoon et al., 2025), and thus we actually use velocities corrected from those described in Section III for perspective effects. The data file including the corrected velocities will be available on Zenodo. While the priors on inclination ($i$) and orbit true anomaly ($\Delta\phi$) described in Section II are fixed because they are statistical properties that a random sample of orbits must satisfy, we consider a wide range of possibilities on the prior probability distribution on eccentricity ($e$): $\alpha=0$ (flat), $\alpha=1$ (thermal), and $\alpha=1.3$ (superthermal for Gaia DR3 wide binaries of relevance: Hwang et al. 2022) in $f_{\rm pr}(e)=(1+\alpha)e^{\alpha}$. Our nominal choice will be the thermal distribution and it is implicitly assumed if not stated otherwise.
In Figure 12, we first examine the individual PDFs of $\Gamma$ for all 75 wide binaries with precise $v_{r}$ (uncertainty <100<100 m s-1) of the basic-cut sample listed in Table 2. This sample includes both the clean sample and the rest. Since the systems not included in the clean sample cannot be verified to be uncontaminated on the basis of the currently available observational information, this figure may include kinematically contaminated cases, although it probably contains mostly pure binaries. Indeed, there are clear outliers with the PDFs of $\Gamma$ covering only the range $\Gamma>0.2$. These abnormal PDFs do not overlap with one narrow PDF peaked at a negative value. In other words, the former and the latter are mutually exclusive, meaning that they cannot obey the same gravity law if their 3D velocities are not contaminated. For this reason, this basic-cut sample is not used to derive a consolidated value of $\Gamma$. However, it will be interesting to investigate those wide binaries with abnormal PDFs in the future, as one such individual system alone, if it is a pure binary with accurate data, can rule out certain gravity models.
IV.1 Main results
Figure 12: This figure shows the individual PDFs of $\Gamma$ for the entire sample of Table 2. Most PDFs overlap one another reasonably well. However, there are several individual cases peaked at large positive values of $\Gamma$ that does not overlap with one narrowly peaked at a negative value. It turns out that none of these mutually excluding systems are included in the clean sample.
The top left panel of Figure 13 shows the PDFs of $\Gamma$ for the clean sample of 36 wide binaries. In contrast to Figure 12, all individual PDFs behave well and overlap well with one another without exception. This is expected if all wide binaries obey the same gravity law and all velocities are uncontaminated and reliably measured. In other words, individual PDFs will represent only random scatters around the underlying gravity law. The properties of the individual PDFs are reassuring of the reliability of all the data involved and the process of selecting the clean sample. Then, the underlying gravity (parameterized by $\Gamma$ in the present study) may be recovered by the statistical consolidation described in Section 2.2 of Chae (2025c) based on Hill and Miller (2011).
The top left panel of Figure 13 shows the statistical consolidation for all wide binaries of the clean sample, while the other panels show various subsamples. The subsamples are considered to test how an increased precision of velocities or/and a designed control of unusually higher velocities (within the clean sample) can affect the inference of gravity. The consolidated PDF for the entire clean sample is well outside the Newtonian value $\Gamma=0$, in stark contrast with the samples with stronger internal acceleration ($10^{-9}$ m s-2) presented in Chae (2025c, b). As can be seen from the curve displayed on the logarithmic scale, the consolidated PDF is approximately Gaussian but not quite exactly. There is a mild left-right asymmetry, and the curve does not decline exactly like Gaussian. We find $\Gamma=0.102_{-0.021}^{+0.023}$, a $60_{-15}^{+17}$% boost to Newtonian gravity, where the quoted formal uncertainties represent the halves of the 95.4% bounds (rather than the 68.3% bounds) from the central median. Based on the Gaussian-like formal uncertainties or the actual PDF, Newton is ruled out at $\gtrsim 5\sigma$. In contrast, the predicted range of MOND gravity models is consistent with the consolidated PDF within $2\sigma$.
Figure 13: This figure shows results on the Bayesian inference of $\Gamma$ for the clean sample. In each panel thin blue curves represent individual PDFs and thick red curve is the consolidated PDF. The left panels show results including wide binaries with $1.0<v_{\rm obs}/v_{\rm escN}\leq 1.2$ while the right panels show results excluding them.
In the bottom left panel of Figure 13, we consider a subsample with highest-precision $v_{r}(=v_{z^{\prime}})$ with uncertainties <50<50 m s-1, to be fully comparable with the precision of the transverse velocity components $v_{x^{\prime}}$ and $v_{y^{\prime}}$. This subsample of 19 wide binaries gives essentially the same value of $\Gamma$ with somewhat larger statistical uncertainties due to the smaller sample size. Again, it is reassuring of our conclusions that increasing the precision cut on the radial velocities yields consistent results when compared to the full clean sample, showing convergence and robustness regarding this parameter.
In the right panels of Figure 13, we consider the subsamples obtained by excluding systems with unusually high 3D velocity $v_{\rm obs}(>v_{\rm escN})$ from the corresponding left-panel samples, where $v_{\rm escN}$ is the Newtonian escape velocity as would be obtained through modeling the system assuming Newtonian gravity with a fixed value of $\Gamma$ (see Figure 8). Based on the Newtonian modeling results, the excluded systems are likely to be individually incompatible with Newtonian gravity unless they are fly-bys with parabolic or hyperbolic orbits caught in instantaneous relative motion by chance or their current velocity data suffer from unknown systematic errors. While these systems can be compatible with MOND gravity with boosted effective gravitational constant, some or all of these systems may be regarded potentially suspicious from the viewpoint of standard gravity. Thus, it is of interest to check whether the MONDian nature of the consolidated PDF persists even if these systems are excluded.
The right panels of Figure 13 show that the consolidated value of $\Gamma$ is significantly reduced, as expected from numerical experiments. For example, modeling and consolidating simulated data (assuming Newtonian gravity) shown in Figures 4 and 8 of Chae (2025c) demonstrate that selectively removing individual PDFs covering only higher values of $\Gamma$ will result in a biased consolidated PDF that is shifted from the assumed gravity. However, the consolidated PDFs are still inconsistent with Newton at a $2\sigma$ level. In other words, the sample as a whole appears to have an anomaly or MOND(-like) signal that cannot be naturally removed. If one wants to interpret our results for the full clean sample within a Newtonian framework, one needs to assume that all $v_{\rm obs}>v_{\rm escN}$ cases, however unlikely they are (see below), must be unbound systems, the removal of which should yield agreement with Newtonian expectations, as all remaining binaries are individually consistent with Newtonian expectations. However, this is not the case. Even removing all $v_{\rm obs}>v_{\rm escN}$ binaries leaves us with a sample that collectively is still discrepant with the Newtonian hypothesis. Since the included binaries are individually consistent with Newton, it is clear that being inconsistent with Newton as a sample stems from requiring orbital parameter distributions (typically inclination and orbital phase) which are at odds with the isotropy and Keplerian assumptions of the method.
Interestingly, the consolidated PDFs excluding systems with $v_{\rm obs}/v_{\rm escN}>1$ agree well with the gravity boost range predicted by MOND gravity numerical simulations (Chae and Milgrom, 2022; Pflamm-Altenburg, 2025) although the distinction from Newton is less pronounced. Thus, it is tempting to exclude the systems with $v_{\rm obs}/v_{\rm escN}>1$ even from a MOND point of view, but we believe that such a choice is not warranted for several reasons.
First of all, the systems with $v_{\rm obs}/v_{\rm escN}>1$ have passed all predefined observational selection criteria for pure wide binaries as the other systems. There is no compelling reason to exclude them on the basis of the currently available observational information. Table 3 summarizes some properties of interest for these systems. All these systems have relatively precise values of $v_{r}$ from Gaia DR3 (Gaia uncertainties smaller than $\approx 0.3$ km s-1 are rare: see Chae 2025c) and they all agree excellently with the more precise independent values measured at different epochs. In all cases, the radial separation is consistent with zero up to the measurement errors. In three cases, the projected separation is less than 20 kau and the metallicity difference between the two stars is nearly zero or not significant considering the Gaia DR3 [M/H] realistic errors (see Section III.3.2). One system (ID #33) may have a moderate issue. The projected separation is relatively large with $s=30.34$ kau and the metallicity difference is the largest among the clean sample with $\Delta[M/H]\equiv[M/H]_{A}-[M/H]_{B}\approx-0.7$. Because the two stars of this system have effective temperatures >4600>4600K, the Gaia DR3 metallicity difference cannot be specifically explained by the known bias. However, the $1\sigma$ range of realistic errors of [M/H] can be as large as $\approx 0.7$ (see Figure 11 of Andrae et al. (2023)). Moreover, $v_{r}$ of this system was measured twice by the LCO and it is consistent with the relatively precise Gaia DR3 value from about 10 years earlier epoch. Thus, even this system needs not be excluded.
Here it is interesting to note that the presence of $v_{\rm obs}/v_{\rm escN}>1$ was first noticed in the Saglia et al. (2025) sample with the occurrence rate of $1/8$ among low-acceleration systems (Chae, 2025b). In our clean sample we have identified 3 more cases from the newly added 28 wide binaries. Thus, these newly added systems suggest that cases with $1<v_{\rm obs}/v_{\rm escN}\lesssim 1.2$ (mildly gravitationally-unbound from the Newtonian perspective) are common with an occurrence rate of about $\approx 0.1$, confirming that the system discovered by Saglia et al. (2025) is not a fluke.
Secondly, the (mildly) Newtonian-unbound systems are not particularly different from the Newtonian-bound systems with $0.9\lesssim v_{\rm obs}/v_{\rm escN}<1$. Table 3 summarizes the key properties of the data and the inferred values of $\Gamma$ for 11 wide binaries with $v_{\rm obs}/v_{\rm escN}\gtrsim 0.9$ that are most responsible for the amplitude of the gravitational anomaly in the clean sample. The values of $v_{p}$ (the scalar sky-plane relative velocity) and $|v_{r}|$ are statistically similar between the two subsamples with $v_{\rm obs}/v_{\rm escN}>1$ or $0.9\lesssim v_{\rm obs}/v_{\rm escN}<1$. The inferred PDFs of $\Gamma$ are all similar with similar median values in the range $0.22\leq\Gamma\leq 0.40$ and similar confidence widths. No individual PDF is abnormal compared with the other PDFs. This means that the gravitational anomaly is not driven by a few exceptional systems but by the collective properties of many systems.
[…]
Table 3: Some properties of wide binaries with $v_{\rm obs}/v_{\rm escN}\gtrsim 0.9$ from the clean sample: the first four are the Newtonian unbound systems with $v_{\rm obs}/v_{\rm escN}>1$.
[…]
[pc] [km s-1] [km s-1] [km s-1] [km s-1] [kau] [kau]
3 131.31 $0.369\pm 0.019$ $0.386\pm 0.051$ $0.527\pm 0.327$ $0.534\pm 0.039$ 19.32 $74.0\pm 94.4$ $0.400_{-0.142}^{+0.234}$ APOGEE
33 73.86 $0.243\pm 0.010$ $0.345\pm 0.060$ $0.181\pm 0.245$ $0.422\pm 0.049$ 30.34 $41.1\pm 37.1$ $0.313_{-0.130}^{+0.239}$ LCO
42 33.55 $0.228\pm 0.005$ $0.595\pm 0.082$ $0.692\pm 0.233$ $0.637\pm 0.076$ 7.870 $7.34\pm 7.35$ $0.277_{-0.131}^{+0.239}$ LCO
59 100.27 $0.383\pm 0.009$ $0.861\pm 0.015$ $1.092\pm 0.239$ $0.942\pm 0.014$ 5.779 $46.1\pm 52.0$ $0.339_{-0.113}^{+0.246}$ HARPS
25 93.79 $0.580\pm 0.009$ $0.337\pm 0.086$ $0.132\pm 0.257$ $0.670\pm 0.043$ 6.524 $13.3\pm 31.4$ $0.243_{-0.117}^{+0.231}$ LCO
32 105.03 $0.344\pm 0.008$ $0.327\pm 0.042$ $0.631\pm 0.254$ $0.474\pm 0.029$ 13.10 $5.01\pm 52.0$ $0.240_{-0.117}^{+0.248}$ LCO
36 82.84 $0.328\pm 0.006$ $0.561\pm 0.082$ $0.514\pm 0.236$ $0.650\pm 0.071$ 6.694 $22.7\pm 23.6$ $0.226_{-0.125}^{+0.223}$ LCO
39 59.47 $475\pm 0.006 0.0$$55\pm 0.046 0.0$$067\pm 0.187$ $0.478\pm 0.008$ 11.054 $42.9\pm 20.2$ $0.247_{-0.124}^{+0.224}$ LCO
53 108.61 $0.341\pm 0.013$ $0.468\pm 0.017$ $0.596\pm 0.286$ $0.579\pm 0.015$ 4.829 $98.8\pm 56.1$ $0.314_{-0.154}^{+0.254}$ HARPS
54 116.26 $0.660\pm 0.009$ $0.449\pm 0.004$ $0.354\pm 0.215$ $0.799\pm 0.008$ 4.935 $18.5\pm 50.6$ $0.222_{-0.108}^{+0.233}$ HARPS
75 40.92 $0.280\pm 0.004$ $0.747\pm 0.021$ $0.668\pm 0.181$ $0.798\pm 0.020$ 4.703 $3.0\pm 9.6$ $0.218_{-0.105}^{+0.223}$ Scarpa
Thirdly, these wide binaries selected from the solar neighborhood of $d<150$ pc are extremely unlikely to be random chance associations. With a local number density of stars $n$ estimated from the Gaia DR3 database and the well-known distribution of peculiar velocity components in the solar neighborhood, we estimate that a randomly selected star to have a fly-by with projected separation $s<s_{\rm max}$, radial separation $l<l_{\rm max}$, and scalar relative 3D velocity $v(\equiv\sqrt{v_{x^{\prime}}^{2}+v_{y^{\prime}}^{2}+v_{z^{\prime}}^{2}})<v_{\rm max}$ is
$p_{\rm chance}\approx$ $(1.4-3.4)\times 10^{-8}\frac{n}{0.15\text{ pc}^{-3}}$ (4)
$\times\left(\frac{s_{\rm max}}{30\text{ kau}}\right)^{2}\frac{l_{\rm max}}{100\text{ kau}}\left(\frac{v_{\rm max}}{1\text{ km s}^{-1}}\right)^{3},$
where the 1-dimensional velocity dispersion of peculiar velocities is assumed to be in the range 304030-40 km s-1. For the clean sample, the nominal choices of Equation (4) are conservative, and thus for a sample of 2 million stars within 150 pc, we expect $\lesssim 0.04$ cases, meaning that practically no fly-by cases meeting the thresholds should be observed. Thus, the observed 4 cases from the clean sample are extremely unlikely to be chance associations unrelated to gravity. Moreover, the 4 cases are a lower limit to the true number of binaries with $v_{\rm obs}>v_{\rm escN}$ because they were selected from samples of wide binaries that meet extremely strict observational constraints (e.g., many from the basic-cut sample were excluded for now, but some of them could be included in the future, once more high quality multi-epoch radial velocities are available). In other words, not all pairs within 150 pc have sufficient data to check whether they meet the selection criteria of the clean sample.
Lastly, existing MOND gravity models do predict [^5] We thank Cezary Migaszewski for sharing numerical simulations results prior to publication. the range of $v_{\rm obs}/v_{\rm escN}$ observed in the clean sample. Moreover, whatever MOND gravity models predict, they (i.e. nonrelativistic gravity models) are far from an established theory. Thus, even the predictions of the current MOND models need to be taken with a grain of salt. Above all, because this work is about a measurement of $\Gamma$, it should not be controlled by any existing gravity models.
Figure 14: This figure shows the inferred value of $\Gamma$ as a function of the internal acceleration $g_{\rm N}$ for the clean sample of 36 wide binaries shown in the top left panel of Figure 13.
Having noted in the above that the systems with $v_{\rm obs}/v_{\rm escN}>1$ need not be excluded from observational and theoretical points of view, we now examine whether $\Gamma$ shows any trend with the internal acceleration $g_{\rm N}$ within our considered limit $g_{\rm N}<10^{-9}$ m s-2. Figure 14 shows the individually inferred values of $\Gamma$ with respect to $\log_{10}g_{\rm N}$ for the clean sample. First of all, individual values provide insights why the consolidated value (which corresponds to the effective or representative value for the whole population) of $\Gamma$ cannot be zero. All individual values of $\Gamma<0$ are consistent with zero within about $1.5\sigma$. However, there are 10 individual values of $\Gamma>0$ that are $\gtrsim 2\sigma$ away from zero. These values shift the consolidated value by $\approx+0.1$ so that all values agree with the shifted consolidated value within $\approx 2\sigma$.
Now while it is not easy to find a well-defined trend of $\Gamma$ with $g_{\rm N}$ from this small sample, it is clear that a consolidated $\Gamma$ is clearly >0>0 in the lowest acceleration regime that is fully MONDian with $\lesssim 10^{-10}$ m s-2. Also, a subsample of the relatively higher acceleration range $10^{-9.5}<g_{\rm N}<10^{-9}$ m s-2 does not include any case with $v_{\rm obs}/v_{\rm escN}>1$ and has a relatively lower value for the consolidated $\Gamma$. This would be qualitatively in agreement with the general trend of MOND gravity. We refrain from further analysis on this point here, as our intent is to focus on a determination of the effective value of $\Gamma$, a well-defined quantity useful to asses the presence (or absence) of a gravitational anomaly with respect to Newtonian gravity.
IV.2 Additional and auxiliary results
Figure 15: Same as Figure 13 but with the flat prior on eccentricity.
In this subsection we present additional results obtained by varying the prior on eccentricity from the nominal choice of the thermal prior. We also check the posterior distributions of the orbit and orientation parameters with respect to the prior distributions required by the randomness of the sample.
Although our nominal choice is the thermal distribution, the most relevant distribution of eccentricities for wide binaries with $s>1$ kau would be a superthermal distribution of $\alpha\approx 1.3$ in $f_{\rm pr}(e)=(1+\alpha)e^{\alpha}$, according to the Hwang et al. (2022) Bayesian study of a large number of wide binaries with precise projected displacements and velocities. Thus, we consider a superthermal prior with $\alpha=1.3$. It turns out that the results with the superhermal prior are nearly indistinguishable from the results with the thermal prior. For the whole clean sample, we have $\Gamma=0.103_{-0.021}^{+0.027}$. This is not surprising because the moderate difference between the thermal distribution and the specific superthermal distribution can make only a tiny effect on the parameter inference when the precise 3D velocities along with the nearly exact two sky-projected displacements have some constraining power on the orbital parameters. We also consider the uninformative flat prior on eccentricity to investigate the role of the eccentricity prior in the gravity inference and at the same time to gauge the power of our sample in constraining the eccentricity distribution.
Figure 15 shows the results on $\Gamma$ with the flat prior on eccentricity. Compared with the nominal results, the inferred value of $\Gamma$ is slightly reduced to $\Gamma=0.092_{-0.017}^{+0.021}$ but the significance of the deviation from Newton is somewhat stronger (a $5.4\sigma$ deviation) because individual PDFs get somewhat narrower, and consequently the consolidated PDF is also narrower. Thus, perhaps contrary to expectation, the flat prior on eccentricity actually makes Newtonian gravity even more strongly discrepant with our wide binary sample.
Figure 16: Posterior distributions of the orbit, inclination, and mass parameters are compared with the priors represented by red lines/curves. The histogram of a parameter represents the composite of the individual posterior PDFs of the parameter for all 36 wide binaries in the clean sample. The prior distribution of $\Delta\phi(=\phi-\phi_{0})$ is the function (with a rescaled amplitude) given by Equation (3) convolved with $e$ shown in the leftmost column.
Figure 16 shows the posterior distributions of the orbit, inclination, and mass parameters for the two Bayesian results with the thermal or flat prior on eccentricity. Due to the degeneracy between the gravity parameter $\Gamma$ and the other parameters in fitting the data $(\mathbf{r},\mathbf{v})$, when $\Gamma$ is free while the other parameters are constrained by their priors, the posterior distributions of the latter are expected to collectively follow the prior distributions as a population. If the posterior distribution of a parameter is significantly different from the prior individually or collectively, that would strongly indicate that the kinematic data prefer a value or distribution different from the prior.
Figure 16 shows that the parameters overall follow the prior distributions. In particular, the posterior distribution of the orbit true anomaly parameter $\Delta\phi(=\phi-\phi_{0})$ is well consistent with the $e$-convolved distribution fulfilling the required self-consistency. This means that the inferred PDFs of $\Gamma$ are based on orbital inclination and occupancy distributions consistent with isotropy and random presentations.
For the thermal prior, the posterior distribution of $e$ matches well the prior distribution. However, for the flat prior, the posterior distribution of $e$ is tilted toward the upper limit indicating that the data prefer a non-flat distribution with $\alpha>0$. However, the relatively moderate tilt may indicate that the kinematic data alone (when $\Gamma$ is free) are not sufficient to constrain well the orbit parameters, especially due to the large uncertainty of the radial separation $z^{\prime}$.
Finally, we note that if gravity is fixed (e.g., Newtonian or a boosted gravity), the orbit and orientation parameters can be significantly constrained by the data $(\mathbf{r},\mathbf{v})$, and the posterior distributions of the parameters, in particular the orbit true anomaly parameter $\Delta\phi$, can be used to test the assumed gravity model, as was investigated in the pilot study by Chae (2025b). In such a test, priors should not be imposed on the orbit and orientation parameters because gravity is fixed (so the degeneracy between gravity and orbit parameters is not a concern) and we seek to test the distributions required by the assumed gravity without the influence of any priors. As already noted with a small sample of 8 or 9 wide binaries with $g_{\rm N}<10^{-9}$ m s-2 by Saglia et al. (2025) and Chae (2025b), Newtonian gravity requires a biased distribution of $\Delta\phi$ towards the periastron to compensate for the boosted gravity obtained for the correct prior distribution (see Figure 16), violating Kepler’s second law in a statistical sense for the population. With a 4 times larger low-acceleration sample than the Saglia et al. (2025) sample, the statistical significance is now clearly stronger. Thorough statistical tests of Newtonian gravity as well as control boosted gravity (toy) models will be presented in a separate work as those work involves statistical methods such as the Kolmogorov-Smirnov test, the Anderson-Darling test, and Bayesian information criterion (see Chae 2025b). We stress that this work is mainly devoted to the construction of the wide binary sample and the “measurement” of $\Gamma$, which is the most reliable and straightforward representation of the data (without requiring additional statistical methods) that can be readily carried out with the available code (Chae, 2025b).
IV.3 Exploring possible systematic errors: Can the gravitational anomaly be removed?
So far we have presented the main and additional results on the gravity inference along with relevant discussions based on the currently available observational information. Here we further consider extreme possibilities to check whether the low-acceleration gravitational anomaly from the clean sample can be removed. The gravity inference may be affected by variation in modeling inputs or sample selection criteria.
As for modeling inputs, we have already considered a sufficiently wide range of possibilities for the orbit eccentricity. It was also shown that our adopted stellar masses agree well with the Gaia FLAME masses (Figure 11) whenever the latter are available. However, given that gravity inference is directly and significantly affected by the input mass, here we consider the possibility that the masses of all our stars might be 10% larger than the values we have adopted. This is in part motivated from the study by Saglia et al. (2025) who quote a maximum of 10% systematic error for their masses based on the HARPS spectra. Our sample is a composite of several subsamples and masses were estimated through independent methods for different subsamples. So it is unlikely that all stellar masses are systematically biased in the same direction by 10%. Thus, the uniform increase of 10% for all stars is an extreme possibility. Figure 17 shows the result on $\Gamma$ with the increased mass. The inferred value is reduced to $\Gamma=0.086_{-0.020}^{+0.025}$, which is however still inconsistent with Newtonian at a more than $4\sigma$ level. This shows that any reasonable adjustment of the currently estimated stellar masses cannot remove the gravitational anomaly.
Figure 17: Same as the top left panel of Figure 13 but with all masses of the stars increased by 10%.
As for sample selection criteria, we have employed the philosophy of sacrificing statistical significance to maximize the purity of the sample, by taking only wide binaries that pass all the stringent selection criteria and have additional observational indicators for kinematically uncontaminated pure binaries. This is why only 17% of the raw sample with relatively precise $v_{r}$ (Figure 2) survived in the final clean sample. The purity of the clean sample may be reaffirmed by the fact that none of the posterior PDFs of $\Gamma$ are abnormal, and in particular, our recovered distributions of orbital phases and inclinations are consistent with isotropy and angular momentum conservation expectations, as shown in Figure 16. However, it is interesting to note that some component stars of our pure binaries are classified as unresolved binaries according to the Gaia DR3 multiple-star classifier (MSC; Fouesneau et al., 2023), making binaries triples or quadruples. In line with Saglia et al. (2025), we did not use the MSC because it is not reliable on an individual basis.
Figure 18: Same as the top left panel of Figure 13 but for a subsample excluding systems marked as unresolved binaries (potentially triples or quadruples) in the Gaia DR3 multiple-star classifier.
It turns out that the Saglia et al. (2025) sample and our clean sample include systems listed as triples or quadruples in the MSC-based Object Type given in the Gaia DR3 online archive. This classification is known to be significant only in a qualitative and statistical sense. Considering our general philosophy of this work, we consider removing systems marked as triples/quadruples by the MSC. Figure 18 shows the result on gravity after excluding the 12 systems (i.e. one third of our clean sample) that include MSC-based unresolved binary stars. The inferred value of $\Gamma=0.096_{-0.028}^{+0.031}$ for the remaining 24 binaries is nearly indistinguishable from that for the entire sample other than the slightly increased statistical errors consistent with the reduced sample size. This indicates that the excluded systems are not special but random. If they were triples/quadruples, the inferred $\Gamma$ for them would be abnormally large because of the systematically underestimated masses and the kinematic contamination due to multiplicity. In that hypothetical case, our inference would shift into accordance with Newtonian expectations once such cases were removed, which is not what happens. This exercise demonstrates that the gravitational anomaly found cannot be removed by invoking hidden unresolved close binaries as per the Gaia DR3 MSC in our clean sample.
Regarding the sample selection, it is of interest to consider only wide binaries whose $v_{r}$ are stable over more than several years from direct multi-epoch observations. Speckle observations show that our targets have a low chance of about 5% to have a faint Speckle resolved companion. Since only 10 wide binaries from the clean sample were directly verified to be free of resolved faint companion from our Speckle observations, it is in principle possible that the remaining 26 wide binaries may include one or two cases that may have a faint companion more than tens of kau apart from a component star. However, removing any one or two wide binaries from our clean sample has a negligible effect on the inferred value of $\Gamma$. Thus, the reaming concern is any unresolved companion within tens of kau from a component star.
Figure 19: Same as Figure 14 but with respect to distance to the binary system.
Companions of relatively close separation less than several tens au may be missed by Speckle observations with a probability increasing with distance because the same angular resolution limit means an increasing limit of physical separation. Our clean sample already excludes a relatively larger distance of >150>150 pc. However, if kinematic contamination in the intermediate separation (that cannot be flagged by Gaia’s ruwe values: see below) from $\approx 10$ au to several tens au is present, it is expected to cause a trend in the inferred $\Gamma$ with distance in our clean sample as wide binaries are distributed over a broad distance range of $10<d<150$ pc. Figure 19 exhibits the individual values of $\Gamma$ with respect to the distances of the binaries in the clean sample. There is not obvious trend of $\Gamma$ with $d$ and the four systems with $v_{\rm obs}/v_{\rm escN}>1$ are distributed over the entire distance range. This indicates that significant kinematic contamination in the intermediate range is unlikely.
Close companions ($\lesssim 10$ au) will have kinematic effects on either tangential velocities (preferentially when the orbit of the companion is close to face-on) or radial velocities (preferentially when the orbit of the companion is close to edge-on). The kinematic effects on tangential velocities can be flagged by Gaia’s ruwe (Belokurov et al., 2020) while those on radial velocities can be flagged through a monitoring observation covering a significant fraction of the orbital period. Because we have already required $\texttt{ruwe}<1.25$, requiring an observed stability of $v_{r}$ over more than several years can complete the observational requirement for pure binaries. Figure 20 shows the inference of $\Gamma$ for 21 wide binaries with time baselines of 3 - 11 years taken (excluding shorter baseline cases) from those shown in the right panel of Figure 10. Again, the inferred value of $\Gamma=0.121_{-0.026}^{+0.034}$ is strongly discrepant with Newton.
Figure 20: Same as the top left panel of Figure 13 but for a subsample with $v_{r}$ stable from two observations separated by 3 - 11 years (except for 4 systems from those shown in the right panel of Figure 10).
Therefore, it seems impossible to remove the gravitational anomaly seen in the clean sample within any conceivable variation of the currently available observational information. Yet, one might be still skeptical of the gravitational anomaly and want to find a way to remove it. As a cautionary tale of any fabricated result through an intentional or unintentional distortion of our sample (or any other samples present and future), we carry out an exercise of deliberately removing choice systems in order to produce a Newtonian result. For this we consider selectively removing wide binaries that are responsible for the gravitational anomaly. Figure 21 shows a contrived result obtained by removing 9 wide binaries (a quarter of the clean sample) that have highest-$v_{\rm obs}/v_{\rm escN}$ values. This is a false result that appears to agree with Newton remarkably well. The removal of the 9 systems with the smallest values of $v_{\rm obs}/v_{\rm escN}$ would correspondingly shift our results to artificially large values of $\Gamma$, while removing 9 random binaries on average leaves our inference unchanged, beyond the expected increase of the confidence intervals due to the reduced sample.
The Gaia limiting magnitude of 20.5 out to the limit of our sample of 150 pc implies that stellar companions at separations beyond the Gaia resolution of 0.3 arcsec can be excluded as contaminating our sample. At closer separations, the combination of only 5% close companions identified by our Speckle campaign and the clearing of a subsample of our final clean sample through this technique, exclude a further range of potential stellar companions. In order to bring our results into consistency with Newtonian expectations we would require, as detailed above, about 9 of our binaries to be contaminated, something which would require a probability of about $(0.05)^{9}$ if unresolved stellar companions were to blame. At even closer separation, where close binary periods become comparable to the Gaia DR3 temporal range of 34 months, the use of ruwe restrictions for all (e.g. see Castro-Ginard et al. (2024) where the presence of unresolved stellar companions of all types has been shown to be incompatible with stars in DR3 for values of $\texttt{ruwe}>1.15$, a limit only one of our binaries in the clean sample crosses) and consistent multi-epoch high quality radial velocity measurements, in practice eliminate stellar companion contamination as a concern.
The only remaining potential contaminants to reconcile our observations with Newtonian gravity would be brown dwarfs at separations above about 3 au, where the inner orbital periods would be larger than 34 months. Here we enter the ‘brown dwarf dessert’, where independent observations have consistently identified an absence of brown dwarf companions to stars of the type we are using, with masses a little below one solar mass (e.g., Kraus et al. 2011). Direct searches using eclipse and lensing surveys have shown this ‘dessert’ to extend out to a few tens of au. Dynamical models of brown dwarf formation have shown instabilities in fragmenting disks to preclude the formation of stable brown dwarfs about stars of the type used here out to about 200 pc (e.g., Vorobyov 2013). Beyond such distances the kinematic signal of such a hypothetical contaminant becomes comparable to the noise level in our study, and the possibility ceases to be relevant. Hence, invoking a hypothetical distribution of undetected brown dwarf companions of sufficient frequency to explain the gravitational anomaly we detect within a Newtonian framework would not only be a contrived and ad hoc proposal lacking any independent evidential support, but would also run counter to all observational and theoretical knowledge on the point.
Figure 21: Same as the top left panel of Figure 13 but for a subsample selectively excluding a quarter of the sample. The excluded 9 wide binaries are those having the highest values of $v_{\rm obs}/v_{\rm escN}$ (see Tables 2 and 3).
V Meanings and Implications of the Results
In this work, we have assembled a sample of pure binaries and then focused on testing Newtonian gravity at low internal acceleration $\lesssim 10^{-9}$ m s-2 by measuring $\Gamma$ assuming elliptical orbits. If observed wide binaries represent random phases of Newtonian elliptical orbits, the consolidated PDF of $\Gamma$ must be consistent with zero as demonstrated for wide binaries with relatively small separation $\lesssim 1$ kau or relatively strong internal acceleration $\gtrsim 10^{-8}$ m s-2 (Chae, 2025c, b). If not, it simply means that Newtonian gravity is broken at low acceleration (whatever the implication for gravity may be) because the result derived assuming the Newtonian framework does not agree with the Newtonian prediction.
The inferred value of $\Gamma$ is a pure measurement/quantification of the degree by which the data deviate from Newton under the specific assumption of elliptical orbits and angular momentum conservation. Thus, while it provides a direct test of Newtonian gravity, it is not straightforward to test specific nonstandard gravity such as MOND models (Bekenstein and Milgrom, 1984; Milgrom, 2010) using $\Gamma$. The orbits of wide binaries in a nonstandard model can only be obtained by numerically solving the nonlinear field equation (Pflamm-Altenburg, 2025). Mock data $(\mathbf{r},\mathbf{v})$ can be obtained from the mock orbits, and the mock data can be modeled in the same manner as the real data. Then, the inferred values of $\Gamma$ for the mock data can be compared with those for the real data obtained here to test the assumed gravity model.
Simply obtaining a boost factor for the gravitational parameter through $\gamma=G/G_{\rm N}=10^{2\Gamma}$ and testing a MOND model with it should be regarded as only a first-order approach. In concordance with the results mostly based on the 2D sky-plane velocities $v_{p}$ from recent works led independently by two of us (Chae, 2023, 2024b, 2024a, 2025c, 2025b; Hernandez, 2023; Hernandez et al., 2024b; Hernandez and Kroupa, 2025; Yoon et al., 2025), the first-order test shows that the MOND predicted gravitational boost is supported by our clean sample.
In the literature, there have been efforts (e.g., see Hernandez et al. (2024a) for a review) to discriminate between Newton and MOND (or MOND-like gravity) in the low-acceleration regime using general samples of wide binaries that include triples and higher-order multiples (and even gravitationally unbound fly-bys). In principle, such an approach can work and should be concordant with the direct measurement of $\Gamma$. For that approach to work, two conditions must be met. First, the prediction of MOND models must be accurately correct. As Pflamm-Altenburg (2025) showed, analytical or numerical calculations under simplifying assumptions (e.g., one-particle equivalent description) can be misleading for wide binary orbital motions under the strong ($\approx 1.8a_{0}$) external field of the Milky Way. None of the existing studies used correct numerical solutions to date. Second, the degeneracy among gravity, the fraction of triples and higher-order multiples ($f_{\rm multi}$) in a sample of apparent binaries, and the fraction of fly-bys should be well under control. The fraction $f_{\rm multi}$ can be calibrated using a control subsample of small-separation binaries whose individual stars are required to satisfy the same observational criteria (e.g., $S/N$ of velocities, ruwe, etc) as those of the main subsample of wide binaries of interest. Fly-bys can be removed by requiring a threshold on the scalar relative velocity between the pair. Unfortunately, the studies in question failed to do either calibrate $f_{\rm multi}$ or remove fly-bys. Our clean sample of wide binaries (and future extended samples) have no fly-bys and satisfy $f_{\rm multi}=0$, so it may be used to discriminate between Newton and MOND in the future.
As identified by recent independent studies (e.g. Pittordis et al., 2025), the most serious systematic up to our present study was the potential presence of undetected multiple systems. Here we have very substantially reduced that possibility through a series of independent approaches such as the use of multi-epoch radial velocity observations, the consistency of Hipparcos and Gaia proper motion observations, the presence of consistent metallicities for the two components of a binary, and the use of Speckle imaging to rule out extra components. Note that Manchanda et al. (2023) recently explored a range of follow-up observational techniques towards solving this issue, and identify Speckle interferometry as the leading technique towards guaranteeing pure binary samples.
VI Summary, Conclusion, and outlook
We have collected a purest possible statistical sample of 36 solar neighborhood wide binaries having low internal accelerations ($\lesssim 10^{-9}$ m s-2) based on an unprecedented combination of various observational data including our crucial new observations of Speckle interferometric imaging and spectroscopic measurements of radial velocities. We have then carried out the most through and complete study of the internal dynamics of wide binaries to date, through the Bayesian 3D modeling methodology.
We started from assembling a new collection of 306 Gaia DR3 wide binaries with relatively precise radial velocities. Speckle interferometric imaging of 414 wide binaries of similar qualities shows that about 5% can have resolvable faint companions. From the initial collection, we carefully selected a high-quality sample of 75 isolated wide binaries, based on limits such as distance <150<150 pc from the Sun, Gaia’s $\texttt{ruwe}<1.25$, a narrow locus in the CM diagram, and relative radial velocity ($v_{r}$) error <100<100 m s-1.
We then included in the final statistical clean sample only those wide binaries for which multi-epoch measurements of $v_{r}$ showed no detectable variability, in most cases over more than several years, and/or Speckle observations have not detected any resolved faint companion. Wide binaries in the sample have highest-quality 3D relative velocities: the scalar sky-plane relative velocity $v_{p}$ (between the pair) has an error smaller than 20 m s-1 with a median of 7.5 m s-1, and the relative radial velocity $v_{r}$ has an error smaller than 100 m s-1 with a median of 47 m s-1. The clean sample is further verified by a consistency check between Hipparcos and Gaia relative proper motions over the $\approx 25$-year period covered by the two observations, and the metallicity consistency between the component stars of the binaries.
The Bayesian inference method of gravity (Chae, 2025c, b) using 3D relative velocities is applied to recover the effective value of the gravitational constant $G(=\gamma G_{N})$ in the low acceleration regime, under the assumption of elliptical orbits and angular momentum conservation. Despite the relatively small sample of 36 wide binaries, the 3D velocity modeling algorithm allows strong constraints on gravity. Our findings can be summarized as follows:
[…]
  • We find that all binaries satisfy $v_{\rm obs}/v_{\rm escN}\lesssim 1.2$ where $v_{\rm obs}$ is the magnitude of the observed 3D relative velocity and $v_{\rm escN}$ is the theoretical escape velocity assuming Newtonian gravity. The upper limit of $\approx 1.2$ is consistent with numerical orbits from numerical solutions of MOND nonlinear field equations. However, it is unclear whether this is a coincidence due to the small sample size or systems with $v_{\rm obs}/v_{\rm escN}>1.2$ will be found in future samples.
  • We find 4 Newtonian-unbound systems with $1<v_{\rm obs}/v_{\rm escN}\lesssim 1.2$ from our clean sample, with an occurrence rate of 1/9. Because $v_{\rm obs}<1$ km s-1 (i.e., they have essentially the same 3D velocities within the Milky Way with a small velocity gradient consistent with a relative motion under mutual gravity) in these systems, the probability that any of these systems is a chance association (fly-by) in 3D space is negligibly small. It will be interesting to see whether this occurrence rate persists in larger future samples.
Our $2\sigma$ range of the boost factor $1.32<\gamma<1.94$ can be consistent with some nonstandard theories of gravity including MOND gravity models (Bekenstein and Milgrom, 1984; Milgrom, 2010). However, the current precision is not sufficient to distinguish between such models, and a proper test of such models require additional analyses based on mock data from numerical solutions of the nonlinear gravitational field equations (see Section V). Ongoing high quality radial velocity and Speckle interferometry observational campaigns will result in larger samples to further refine the results presented here.
We thank Cezary Migaszewski for sharing the distributions of $v_{\rm obs}/v_{\rm escN}$ obtained from numerical solutions of MOND gravity. We thank R. A. M. Cortés for the assistance with the Hipparcos archive data. We thank Arthur Kosowsky, Cezary Migaszewski, and Jan Pflamm-Altenburg for discussions. This work was supported by the National Research Foundation of Korea (NRF-2022R1A2C1092306). X.H. acknowledges financial assistance from SECIHTI SNII and UNAM DGAPA PAPIIT grant IN-102624. V.O. acknowledges financial assistance from UNAM DGAPA PAPIIT grant IN-114123. D.L. acknowledges support from Basic Science Research Program through the National Research Foundation of Korea funded by the Ministry of Education (RS-2025-25419201 and RS-2022-NR070872). Y.-W.L. acknowledges support from the NRF of Korea to the Center for Galaxy Evolution Research (RS-2022-NR070872, RS-2022-NR070525). This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. The LCO measurements of radial velocities of stars were carried out with telescopes at McDonald Observatory in the USA, South African Astronomical Observatory (SAAO) in South Africa, and WISE Observatory in Israel. The Speckle observations reported were acquired at the Observatorio Astronómico Nacional in the Sierra San Pedro Mártir (OAN-SPM), Baja California, México. We thank the daytime and night support staff at the OAN-SPM for facilitating and helping obtain our observations. The MAROON-X observations were supported by K-GMT Science Program (PID: GN-2024B-Q-122) of Korea Astronomy and Space Science Institute (KASI).
Appendix A Description of LCO observations and data reduction of radial velocities
A sample of 60 wide binaries were selected from the El-Badry et al. (2021) catalog with the following requirements: (1) Both stars of a binary are brighter than $G=11$ mag and are in the absolute magnitude range $3\lesssim M_{G}<8$, (2) the sky-plane separation $s>3.5$ kau, and (3) the relative Gaia DR3 RV satisfies
$|v_{r}|\equiv|{\rm RV}_{A}-{\rm RV}_{B}|<\sqrt{9(\sigma_{A}^{2}+\sigma_{B}^{2})+(\Delta V)^{2}}$ (A1)
where $\Delta V=0.9419\sqrt{M_{\rm tot}/s}\times 1.3\times 1.2$ km s-1 for the binary total mass $M_{\rm tot}$ given in $M_{\odot}$ and the sky-plane separation $s$ given in kau. This threshold is slightly relaxed from that given by Equations (2) and (3) of Chae (2024b). We note that we impose only a requirement on the relative RV because the El-Badry et al. (2021) catalog was obtained with a requirement on the relative PM.
[…]
The data extracted from the LCO archive were bias and flat-field corrected images processed with the BANZAI pipeline (McCully et al., 2018). BANZAI-NRES is designed to handle all data from the NRES of the LCO network. Stellar RVs in LCO/NRES are estimated by comparing the BLAZ-extracted spectrum with a corresponding ZERO file, derived from PHOENIX stellar models (Husser et al., 2013). Each target star is assigned a specific ZERO file to ensure consistency. The pipeline first determines an initial redshift by cross-correlating the BLAZ and ZERO spectra. It then interpolates the ZERO spectrum to this redshift and divides each spectral order into wavelength segments. The residual redshift for each segment is then estimated along with formal errors. Finally, multiple estimates of the mean redshift are computed using different averaging methods.
The BANZAI-NRES pipeline provides extracted and wavelength-corrected spectra. If the target is a star, the pipeline also delivers RV measurements and stellar classification parameters, such as effective temperature and surface gravity. The RV precision of the BANZAI-NRES pipeline has been demonstrated to be as good as 10 m s-1 for bright (V $\approx$ 6) standard stars. For this study, we used the cross-correlation function (CCF) implemented in the pipeline, with approximate errors of 71 m s-1 for the entire sample. A summary of the observations can be found in Table 4.
[…]
Table 4: Summary of Wide Binaries Observed with LCO/NRES
[…]
[kau] $M_{\odot}$] [km s-1] [km s-1] [km s-1]
1A BD+65 14 528186492427483392 5.246 1.18 F6V $7.967\pm 0.261$ $8.116\pm 0.064$ $0.652\pm 0.087$ (1)
1B BD+65 12 528187248343141376 1 G1V $8.442\pm 0.342$ $8.768\pm 0.059$ (1)
2A HD 236377 428811394564909568 5.063 1.2 F8 $6.402\pm 0.234 66$$6.074\pm 0.037$ $0.647\pm 0.149$ (1)
2B BD+59 32B 428811428924644352 0.85 G0 $65.447\pm 0.245$ $65.427\pm 0.144$ (1)
3A BD-09 71 2426718786382002304 9.547 0.97 $14.245\pm 0.570$ $13.311\pm 0.084$ $0.772\pm 0.142$ (1)
3B TYC 5262-512-1 2426718614583009280 0.99 $13.280\pm 0.487$ $12.539\pm 0.115$ (2)
4A HD 4552 2776055105362407680 4.839 1.2 F8 $14.913\pm 0.236$ $14.957\pm 0.140$ $0.605\pm 0.150$ (3)
4B BD+12 90 2776054899203977728 1.0 G0 $15.509\pm 0.161$ $15.562\pm 0.054$ (3)
5A BD-03 107 2529661799483080192 5.373 1.13 F6V $12.528\pm 0.274$ $12.456\pm 0.100$ $0.560\pm 0.109$ (3)
5B TYC 4674-49-1 2529474126591876096 0.83 G6V $2.306\pm 0.303 11$$1.896\pm 0.044$ (1)
6A HD 8745 322414398818886656 13.951 1.16 F8 $8.390\pm 0.130$ $7.901\pm 0.039$ $0.423\pm 0.062$ (1),(1)
6B BD+36 251 322413681559182720 0.86 K0 $8.050\pm 0.173$ $7.478\pm 0.048$ (1),(1)
7A HD 9769 5038962632187714432 10.152 1.33 F5V $31.107\pm 0.144$ $31.413\pm 0.054$ $0.212\pm 0.088$ (2)
7B CD-24 668 5038962219872444160 1.09 $30.965\pm 0.220$ $31.201\pm 0.069$ (2)
8A HD 11584 4940794866807373952 4.933 1.18 F6/8V $2.731\pm 0.145 22$$2.773\pm 0.038$ $0.387\pm 0.069$ (2),(2)
8B CD-50 524 4940794488850252928 1.15 $2.377\pm 0.159 22$$2.386\pm 0.058$ (3),(3)
9A HD 18014 5185524920830592128 4.044 0.74 K4V $38.489\pm 0.141$ $38.950\pm 0.051$ $0.276\pm 0.086$ (2)
9B BD-04 488 5185536774940328448 0.7 K5V $38.386\pm 0.267$ $38.674\pm 0.069$ (1)
10A BD-12 743 5114547700047886976 3.670 0.91 $86.111\pm 0.204$ $86.371\pm 0.046$ $0.392\pm 0.112$ (3)
10B TYC 5307-1283-1 5114544745110388352 0.82 $86.370\pm 0.229$ $86.763\pm 0.102$ (2)
11A HD 24820 3193531802050857472 10.787 1.24 F6V $1.067\pm 0.211$ $1.821\pm 0.222$ $0.715\pm 0.228$ (2)
11B BD-10 790 3193508059471646464 1.05 $0.737\pm 0.195$ $1.106\pm 0.052$ (2)
12A HD 25384 5090701698022451200 5.449 1.31 F5V $5.975\pm 0.173$ $5.915\pm 0.097$ $0.072\pm 0.161$ (2)
12B TYC 5888-565-1 5090701698022451072 1.14 $5.824\pm 0.238$ $5.843\pm 0.129$ (2)
13A HD 26440 3203601851092155776 22.891 1.07 G3V $24.222\pm 0.175$ $24.604\pm 0.088$ $0.361\pm 0.200$ (3)
13B TYC 4729-523-1 3203683695989446400 0.95 $24.626\pm 0.225$ $24.965\pm 0.180$ (3)
14A BD+05 650 3285755951169814656 26.874 1.12 G0 $23.149\pm 0.225$ $23.765\pm 0.105$ $0.050\pm 0.120$ (3)
14B BD+05 653 3285744612456149888 1.05 G2 $23.260\pm 0.307$ $23.715\pm 0.058$ (3)
15A HD 29356 3230677565443833088 4.972 1.16 F7/G0 $38.397\pm 0.135$ $38.989\pm 0.042$ $0.076\pm 0.057$ (3),(3)
15B HD 29355 3230677874682668672 1.15 F3/6 $38.724\pm 0.133$ $39.065\pm 0.038$ (3),(3)
16A zeta Dor 4763906879239461632 3.760 1.09 F9VFe-0.5 $1.450\pm 0.125$ $0.976\pm 0.102$ $0.581\pm 0.123$ (2),(2)
16B CD-57 1079 4763897739549071744 0.57 K7Vk $1.135\pm 0.130$ $0.395\pm 0.068$ (3),(3)
17A HD 35376 4763357363943383808 5.022 0.95 G3/5V $17.604\pm 0.152$ $17.985\pm 0.018$ $0.283\pm 0.092$ (2)
17B CD-57 1160 4763357260864168320 0.82 $17.798\pm 0.231$ $18.268\pm 0.090$ (3)
18A HD 247101 3445112363273415808 25.589 1.3 $1.655\pm 0.277$ $1.822\pm 0.060$ $0.147\pm 0.077$ (1)
18B HD 247123 3445112122755273472 1.14 $2.167\pm 0.195$ $1.675\pm 0.049$ (1)
19A CD-22 2824 2913801715936708352 22.722 1.23 F5 $14.846\pm 0.191$ $14.501\pm 0.210$ $0.232\pm 0.221$ (2)
19B TYC 5945-1833-1 2913814119802247296 0.99 $14.241\pm 0.254$ $14.269\pm 0.070$ (3)
20A HD 48540 2925423072806388608 22.211 1.3 F3V $1.764\pm 0.141$ $1.768\pm 0.121$ $0.226\pm 0.132$ (2)
20B TYC 6521-2073-1 2925423931801365760 1.09 $1.558\pm 0.234$ $1.994\pm 0.054$ (2)
21A HD 49496 1003223614961194752 4.710 1.01 G0V $23.958\pm 0.226$ $23.738\pm 0.042$ $0.430\pm 0.077$ (1)
21B TYC 3778-1205-1 1003223584897948160 0.87 K2V $23.903\pm 0.220$ $23.308\pm 0.065$ (1)
22A TYC 752-1389-1 3158926322836178816 10.968 0.77 $17.510\pm 0.429$ $16.958\pm 0.058$ $0.023\pm 0.066$ (2)
22B TYC 752-1649-1 3158878734598549376 0.89 NaN $17.575\pm 0.427$ $16.981\pm 0.032$ (2)
23A HD 51943 2932313231147240960 9.840 1.28 F5/6V $17.921\pm 0.157$ $17.739\pm 0.077$ $0.073\pm 0.085$ (2)
23B CPD-19 1595 2932313196787509376 1.11 F6V $17.943\pm 0.169$ $17.666\pm 0.037$ (2)
24A HD 53566 3116331104937277952 30.340 1.26 F3/5V $15.010\pm 0.140$ $14.859\pm 0.046$ $0.345\pm 0.060$ (2),(2)
24B TYC 170-2913-1 3116324881524878208 0.84 F9 $14.830\pm 0.201$ $14.514\pm 0.039$ (2),(2)
25A TYC 5389-255-1 3046204180298475264 6.127 0.80 G7V $59.201\pm 0.356$ $59.393\pm 0.075$ $0.413\pm 0.131$ (2)
25B TYC 5389-869-1 3046204150242468480 0.80 G9V $59.243\pm 0.375$ $59.806\pm 0.108$ (2)
26A BD+82 192 1142787168495168000 3.772 0.80 G5 $1.756\pm 0.218 11$$1.980\pm 0.054$ $0.317\pm 0.075$ (1)
26B BD+82 193 1142786996696476288 0.74 G5 $11.858\pm 0.272$ $12.297\pm 0.052$ (1)
27A TYC 1364-1623-1 3170300942420466176 9.723 1.04 $27.738\pm 0.243$ $26.993\pm 0.038$ $0.055\pm 0.058$ (1),(2)
27B TYC 1364-1760-1 3170394607068638336 1.01 $26.989\pm 0.207$ $26.938\pm 0.044$ (2),(2)
28A HD 71050 914241517609344128 19.612 1.16 G0 $3.634\pm 0.339$ $3.251\pm 0.107$ $0.133\pm 0.123$ (1)
28B HD 70986 914244399532441472 1.11 G5 $636\pm 0.322 3.3$$384\pm 0.061$ (1)
29A HD 72584 5755402720924702464 7.910 1.11 F6V $12.566\pm 0.259$ $12.674\pm 0.049$ $0.196\pm 0.127$ (1),(3)
29B TYC 4874-1253-1 5755402484702384128 0.79 G3 $12.516\pm 0.206$ $12.869\pm 0.117$ (3),(3)
30A HD 78796 5651775953326498688 7.518 0.96 G5V $741\pm 0.150 3.3$$314\pm 0.042$ $0.119\pm 0.081$ (2),(3)
30B CD-23 8096 5651752515687994368 0.8 K0 $3.872\pm 0.194$ $3.195\pm 0.069$ (3),(1)
31A HD 81268 5741345125461459200 13.101 1.04 G0V $14.433\pm 0.174$ $14.853\pm 0.024$ $0.327\pm 0.042$ (1),(3)
31B BD-08 2665 5741344919302959360 0.97 G0 $15.064\pm 0.185$ $15.180\pm 0.034$ (1),(2)
32A HD 85137 5305981470567619456 6.524 1.14 F7V $3.841\pm 0.167$ $3.917\pm 0.048$ $0.337\pm 0.086$ (2),(2)
32B CD-57 2838 5305981745445719680 0.84 $3.974\pm 0.195$ $4.254\pm 0.071$ (3),(3)
33A HD 88418 3861210890850859392 10.090 1.24 F8 $43.494\pm 0.205$ $42.783\pm 0.085$ $0.339\pm 0.111$ (2),(2)
33B TYC 251-458-1 3861210925210595072 0.88 $42.704\pm 0.238$ $42.444\pm 0.071$ (3),(1)
34A TYC 6068-1392-1 5668676409117097984 5.728 1.13 $14.247\pm 0.340$ $14.649\pm 0.074$ $0.254\pm 0.139$ (3)
34B TYC 6068-1401-1 5668675653202854528 1.06 $14.349\pm 0.333$ $14.903\pm 0.118$ (2)
35A HD 92677 3749791158495959552 4.181 0.92 G6/8(IV) $1.042\pm 0.172$ $1.408\pm 0.045$ $0.383\pm 0.073$ (2),(2)
35B HD 92652 3749791055416743552 0.9 G8/K1 $0.642\pm 0.177$ $1.025\pm 0.057$ (3),(3)
36A HD 93528 3550081879381593728 7.870 0.82 K0V $23.759\pm 0.148$ $24.129\pm 0.044$ $0.595\pm 0.082$ (2)
36B BD-21 3153 3550084490721711872 0.7 K4.5Vk $24.451\pm 0.180$ $24.724\pm 0.069$ (2)
37A HD 95532 3556353734223925504 7.777 1.27 F7/8V $27.258\pm 0.134$ $26.931\pm 0.040$ $0.252\pm 0.080$ (3)
37B HD 95531 3556356070686135040 1.05 G2/3(V) $27.093\pm 0.191$ $26.679\pm 0.069$ (3)
38A HD 101574 3793107930900527616 16.371 1.22 F5V $0.382\pm 0.166$ $0.826\pm 0.061$ $0.260\pm 0.094$ (2),(1)
38B BD-01 2557 3793106419072038272 1.03 $0.367\pm 0.208$ $0.566\pm 0.072$ (3),(3)
39A HD 103231 3487243037508315648 10.721 1.12 G0V $8.786\pm 0.135$ $8.343\pm 0.052$ $0.170\pm 0.060$ (2)
39B HD 103206 3487237024554098560 0.98 G6V $8.823\pm 0.167$ $8.513\pm 0.030$ (2)
40A HD 233884 787833551986183168 4.547 1.07 K0 $5.034\pm 0.199$ $4.746\pm 0.079$ $0.386\pm 0.088$ (1)
40B TYC 3454-370-1 787833483266126080 0.86 $5.199\pm 0.244$ $5.132\pm 0.038$ (1)
41A HD 105350 6147940161728145536 22.167 0.98 G5V $28.584\pm 0.142$ $29.068\pm 0.061$ $0.250\pm 0.109$ (3),(3)
41B TYC 7763-590-1 6147937516029209728 0.74 $28.704\pm 0.205$ $29.318\pm 0.090$ (2),(2)
42A HD 107434 6151377578674752896 11.180 1.15 F6V $10.499\pm 0.129$ $10.067\pm 0.080$ $0.569\pm 0.108$ (2),(2)
42B CD-37 7822 6151049133934677248 0.8 $10.061\pm 0.192$ $9.499\pm 0.073$ (3),(3)
43A BD+82 377 1719835231806217472 8.476 1.07 G0 $3.369\pm 0.277 33$$3.532\pm 0.038$ $0.297\pm 0.060$ (1)
43B BD+82 376 1719835407900844544 1.05 G5 $3.177\pm 0.232 33$$3.235\pm 0.047$ (1)
44A HD 123033 1258410612976538368 4.273 1.2 F6V $20.046\pm 0.121$ $19.847\pm 0.066$ $0.632\pm 0.082$ (3),(1)
44B BD+26 2522 1258410750415492864 0.87 K0 $19.854\pm 0.146$ $19.215\pm 0.048$ (1),(3)
45A HD 124711 1482432155767129728 12.657 1.21 $2.591\pm 0.141$ $2.090\pm 0.083$ $0.068\pm 0.101$ (1)
45B BD+36 2455 1479430076707007488 0.93 $2.355\pm 0.195$ $2.022\pm 0.058$ (1)
46A HD 127058 5899243585161684864 6.692 1.03 F8/G0V $3.428\pm 0.148$ $3.955\pm 0.032$ $0.566\pm 0.056$ (2)
46B CD-49 8796 5899244375435693952 0.92 G5 $4.332\pm 0.159$ $4.521\pm 0.046$ (3)
47A BD+08 2889 1172915990414659328 24.684 0.85 G5 $18.422\pm 0.248$ $17.964\pm 0.029$ $0.027\pm 0.077$ (2),(3)
47B BD+08 2887 1172920487244742912 0.8 K2 $18.415\pm 0.261$ $17.937\pm 0.071$ (3),(3)
48A HD 129171 1282815063829295360 16.772 1.03 G0 $1.925\pm 0.121 11$$1.406\pm 0.026$ $0.283\pm 0.041$ (1),(1)
48B HD 129209 1282817022334383232 1 G2IV $1.598\pm 0.124 11$$1.123\pm 0.032$ (1),(1)
49A TYC 364-256-1 4430185068482324864 4.426 0.96 $34.756\pm 0.198$ $34.665\pm 0.032$ $0.025\pm 0.072$ (3)
49B TYC 364-158-1 4430185034123000960 0.99 $34.649\pm 0.192$ $34.640\pm 0.064$ (2)
50A 41 Her 4435689739087756800 7.159 1.12 G8 $6.595\pm 0.119$ $6.124\pm 0.019$ $0.637\pm 0.097$ (3)
50B Ross 643 4435683451255623808 0.71 K3 $7.031\pm 0.165$ $6.761\pm 0.095$ (2)
51A BD+27 2769 4574764697142222592 14.269 1.19 F5 $2.903\pm 0.184 33$$3.382\pm 0.250$ $1.078\pm 0.266$ (1)
51B BD+27 2768 4574670895056352896 1.07 F8 $32.563\pm 0.162$ $32.304\pm 0.092$ (1)
52A HD 158226 4599984642025088128 4.731 0.99 G1V $73.239\pm 0.132$ $72.809\pm 0.028$ $0.086\pm 0.062$ (1)
52B BD+31 3025 4599984504586131456 0.88 G8 $73.024\pm 0.234$ $72.895\pm 0.055$ (1)
53A HD 352384 1815165535636339072 11.565 1.14 G0 $51.028\pm 0.171$ $51.307\pm 0.026$ $0.853\pm 0.047$ (1)
53B HD 352383 1815165883534980992 0.97 G0 $51.103\pm 0.213$ $52.160\pm 0.039$ (1)
54A HD 201706 1760471948915107200 4.263 1.02 G0 $14.790\pm 0.139$ $15.896\pm 0.027$ $0.947\pm 0.047$ (1)
54B BD+14 4553 1760477618271932672 0.81 G5 $14.779\pm 0.160$ $14.949\pm 0.038$ (1)
55A BD+41 4134 1967283042361454848 6.695 1.04 G0 $24.763\pm 0.163$ $24.385\pm 0.038$ $0.561\pm 0.082$ (1)
55B BD+41 4133 1967282939282261120 0.84 $24.249\pm 0.171$ $23.824\pm 0.073$ (1)
56A HD 207397 6840365718216434816 4.515 0.97 G3V $18.847\pm 0.195$ $19.153\pm 0.180$ $0.005\pm 0.185$ (1)
56B BD-14 6130 6840365615137220096 0.81 $19.154\pm 0.264$ $19.148\pm 0.044$ (1)
57A HD 209032 6570796871887419648 3.952 0.96 K1(IVp) $18.897\pm 0.157$ $19.334\pm 0.018$ $0.105\pm 0.069$ (2)
57B TYC 8001-622-1 6570797524722448896 0.92 $19.156\pm 0.146$ $19.229\pm 0.067$ (3)
58A TYC 4264-485-1 2201661297490051968 8.370 1.03 $12.734\pm 0.251$ $12.514\pm 0.077$ $0.002\pm 0.110$ (1)
58B TYC 4264-252-1 2201661091331626752 1.04 $12.387\pm 0.268$ $12.516\pm 0.079$ (1)
59A HD 220721 1938247517245907456 11.055 1.05 G0 $15.506\pm 0.125$ $15.045\pm 0.029$ $0.055\pm 0.046$ (1)
59B HD 220748 1938247654684985216 0.97 G5 $15.439\pm 0.139$ $14.990\pm 0.036$ (1)
60A HD 224478 2739496687336985984 7.122 1.13 F6V $14.919\pm 0.340$ $16.068\pm 0.300$ $0.716\pm 0.344$ (1)
60B AG+02 2992 2739498199165473920 1.01 G $15.254\pm 0.324$ $15.352\pm 0.168$ (1)
A.1 Systems that were observed more than once with LCO
Eighteen of the 60 binary systems were observed more than once. Their respective properties are listed below.
6A (HD 8745) was observed twice at McDonald Observatory. RVs of $7.955\pm 0.040$ km s-1 and $7.846\pm 0.037$ km s-1 were measured on 2024 December 14 (BJD 2460659.703) and 2025 February 9 (BJD 2460716.58759), respectively. The mean RV is $7.901\pm 0.039$ km s-1. 6B (BD+36 251) was observed twice at McDonald Observatory. RVs of $7.545\pm 0.031$ km s-1 and $7.411\pm 0.056$ km s-1 were measured on 2024 December 14 (BJD 2460659.764) and 2025 February 9 (BJD 2460716.6159), respectively. The mean RV is $7.478\pm 0.048$ km s-1.
8A (HD 11584) was observed twice at the South African Astronomical Observatory (SAAO). RVs of $22.761\pm 0.030$ km s-1 and $22.784\pm 0.044$ km s-1 were measured on 2024 December 14 (BJD 2460659.437) and 2025 February 11 (BJD 2460718.29242), respectively. The mean RV is $22.773\pm 0.038$ km s-1. 8B (CD–50 524) was observed twice at the WISE Observatory. RVs of $22.398\pm 0.056$ km s-1 and $22.374\pm 0.060$ km s-1 were measured on 2024 December 18 (BJD 2460663.566) and 2025 February 17 (BJD 2460724.5223), respectively. The mean RV is $22.386\pm 0.058$ km s-1.
15A (HD 29356) was observed twice at the WISE Observatory. RVs of $38.938\pm 0.050$ km s-1 and $39.039\pm 0.030$ km s-1 were measured on 2025 January 8 (BJD 2460684.54458) and 2025 February 9 (BJD 2460716.5881), respectively. The mean RV is $38.989\pm 0.042$ km s-1. 15B (HD 29355) was observed twice at the WISE Observatory. RVs of $39.036\pm 0.033$ km s-1 and $39.093\pm 0.043$ km s-1 were measured on 2025 January 7 (BJD 2460683.71539) and 2025 February 7 (BJD 2460714.62182), respectively. The mean RV is $39.065\pm 0.038$ km s-1.
16A ($\zeta$ Dor) was observed twice at SAAO. RVs of $0.966\pm 0.103$ km s-1 and $0.985\pm 0.100$ km s-1 were measured on 2024 December 14 (BJD 2460659.584) and 2025 February 28 (BJD 2460735.37616), respectively. The mean RV is $0.976\pm 0.102$ km s-1. 16B (CD-57 1079) was observed twice at the WISE Observatory. RVs of $0.185\pm 0.074$ km s-1 and $0.605\pm 0.061$ km s-1 were measured on 2024 December 18 (BJD 2460663.73559) and 2025 March 4 (BJD 2460739.55929), respectively. The mean RV is $0.395\pm 0.068$ km s-1.
24A (HD 53566) was observed twice at SAAO. RVs of $14.822\pm 0.045$ km s-1 and $14.896\pm 0.047$ km s-1 were measured on 2025 January 10 (BJD 2460686.47896) and 2025 February 12 (BJD 2460719.37888), respectively. The mean RV is $14.859\pm 0.046$ km s-1. 24B (TYC 170-2913-1) was observed twice at SAAO. RVs of $14.510\pm 0.024$ km s-1 and $14.519\pm 0.049$ km s-1 were measured on 2025 January 7 (BJD 2460683.4967) and 2025 February 14 (BJD 2460721.37678), respectively. The mean RV is $14.515\pm 0.039$ km s-1.
27A (TYC 1364-1623-1) was observed twice, at SAAO and McDonald Observatory, respectively. RVs of $27.039\pm 0.026$ km s-1 and $26.946\pm 0.047$ km s-1 were measured on 2024 December 19 (BJD 2460664.47497) and 2025 February 16 (BJD 2460723.81325), respectively. The mean RV is $26.993\pm 0.038$ km s-1. 27B (TYC 1364-1760-1) was observed twice at SAAO. RVs of $26.961\pm 0.039$ km s-1 and $26.916\pm 0.049$ km s-1 were measured on 2024 December 19 (BJD 2460664.47497) and 2025 February 14 (BJD 2460721.35215), respectively. The mean RV is $26.939\pm 0.044$ km s-1.
29A (HD 72584) was observed twice, at McDonald Observatory and WISE Observatory, respectively. RVs of $12.749\pm 0.043$ km s-1 and $12.599\pm 0.055$ km s-1 were measured on 2024 December 14 (BJD 2460659.941) and 2025 March 3 (BJD 2460738.70737), respectively. The mean RV is $12.674\pm 0.049$ km s-1. 29B (TYC 4874-1253-1) was observed twice at WISE Observatory. RVs of $12.854\pm 0.077$ km s-1 and $12.885\pm 0.146$ km s-1 were measured on 2024 December 15 (BJD 2460660.700) and 2025 March 2 (BJD 2460737.69814), respectively. The mean RV is $12.870\pm 0.117$ km s-1.
30A (HD 78796) was observed twice at WISE Observatory and SAAO, respectively. RVs of $3.313\pm 0.047$ km s-1 and $3.316\pm 0.037$ km s-1 were measured on 2024 December 14 (BJD 2460659.772) and 2025 February 11 (BJD 2460718.34031), respectively. The mean RV is $3.315\pm 0.042$ km s-1. 30B (CD-23 8096) was observed twice at McDonald Observatory and WISE Observatory, respectively. RVs of $3.195\pm 0.057$ km s-1 and $3.195\pm 0.079$ km s-1 were measured on 2024 December 18 (BJD 2460663.93401) and 2025 February 11 (BJD 2460718.71122), respectively. The mean RV is $3.195\pm 0.069$ km s-1.
31A (HD 81268) was observed twice at McDonald Observatory and WISE Observatory, respectively. RVs of $14.892\pm 0.026$ km s-1 and $14.815\pm 0.021$ km s-1 were measured on 2024 December 16 (BJD 2460661.891) and 2025 February 13 (BJD 2460720.64292), respectively. The mean RV is $14.854\pm 0.024$ km s-1. 31B (BD-08 2665) was observed twice at McDonald Observatory and SAAO, respectively. RVs of $15.214\pm 0.036$ km s-1 and $15.147\pm 0.031$ km s-1 were measured on 2024 December 16 (BJD 2460661.908) and 2025 February 15 (BJD 2460722.53986), respectively. The mean RV is $15.181\pm 0.034$ km s-1.
32A (HD 85137) was observed twice at SAAO. RVs of $3.960\pm 0.048$ km s-1 and $3.874\pm 0.048$ km s-1 were measured on 2024 December 13 (BJD 2460658.486) and 2025 February 9 (BJD 2460716.33343), respectively. The mean RV is $3.917\pm 0.048$ km s-1. 32B (CD-57 2838) was observed twice at WISE Observatory. RVs of $4.353\pm 0.081$ km s-1 and $4.155\pm 0.060$ km s-1 were measured on 2024 December 13 (BJD 2460658.715) and 2025 February 9 (BJD 2460716.70867), respectively. The mean RV is $4.254\pm 0.071$ km s-1.
33A (HD 88418) was observed twice at SAAO. RVs of $42.705\pm 0.096$ km s-1 and $42.861\pm 0.072$ km s-1 were measured on 2025 February 15 (BJD 2460722.37206) and 2025 March 4 (BJD 2460739.43796), respectively. The mean RV is $42.783\pm 0.085$ km s-1. 33B (TYC 251-458-1) was observed twice at McDonald Observatory and WISE Observatory, respectively. RVs of $42.430\pm 0.081$ km s-1 and $42.458\pm 0.059$ km s-1 were measured on 2025 February 15 (BJD 2460722.84552) and 2025 March 3 (BJD 2460738.78113), respectively. The mean RV is $42.444\pm 0.071$ km s-1.
35A (HD 92677) was observed twice at SAAO. RVs of $1.398\pm 0.030$ km s-1 and $1.418\pm 0.056$ km s-1 were measured on 2024 December 13 (BJD 2460658.572) and 2025 February 10 (BJD 2460717.46585), respectively. The mean RV is $1.408\pm 0.045$ km s-1. 35B (HD 92652) was observed twice at WISE Observatory. RVs of $1.030\pm 0.047$ km s-1 and $1.019\pm 0.066$ km s-1 were measured on 2024 December 13 (BJD 2460658.842) and 2025 February 9 (BJD 2460716.72902), respectively. The mean RV is $1.025\pm 0.057$ km s-1.
38A (HD 101574) was observed twice at McDonald Observatory and SAAO, respectively. RVs of $1.665\pm 0.045$ km s-1 and $0.013\pm 0.073$ km s-1 were measured on 2024 December 15 (BJD 2460734.79018) and 2025 February 10 (BJD 2460717.43468), respectively. The mean RV is $0.826\pm 0.061$ km s-1. 38B (BD-01 2557) was observed twice at WISE Observatory. RVs of $0.573\pm 0.047$ km s-1 and $0.559\pm 0.090$ km s-1 were measured on 2024 December 15 (BJD 2460660.840) and 2025 February 10 (BJD 2460717.67514), respectively. The mean RV is $0.566\pm 0.072$ km s-1.
41A (HD 105350) was observed twice at WISE Observatory. RVs of $29.135\pm 0.074$ km s-1 and $29.000\pm 0.045$ km s-1 were measured on 2024 December 14 (BJD 2460659.794) and 2025 March 3 (BJD 2460738.86556), respectively. The mean RV is $29.068\pm 0.061$ km s-1. 41B (TYC 7763-590-1) was observed twice at SAAO. RVs of $29.453\pm 0.107$ km s-1 and $29.182\pm 0.069$ km s-1 were measured on 2024 December 24 (BJD 2460669.51094) and 2025 March 4 (BJD 2460739.56682), respectively. The mean RV is $29.318\pm 0.090$ km s-1.
42A (HD 107434) was observed twice at SAAO. RVs of $9.966\pm 0.084$ km s-1 and $10.169\pm 0.076$ km s-1 were measured on 2024 December 13 (BJD 2460658.555) and 2025 February 28 (BJD 2460735.54704), respectively. The mean RV is $10.067\pm 0.080$ km s-1. 42B (CD-37 7822) was observed twice at WISE Observatory. RVs of $9.649\pm 0.068$ km s-1 and $9.348\pm 0.077$ km s-1 were measured on 2024 December 12 (BJD 2460657.841) and 2025 February 28 (BJD 2460735.65839), respectively. The mean RV is $9.499\pm 0.073$ km s-1.
44A (HD 123033) was observed twice at McDonald Observatory and WISE Observatory. RVs of $19.677\pm 0.069$ km s-1 and $20.018\pm 0.064$ km s-1 were measured on 2025 February 17 (BJD 2460724.83569) and 2025 March 3 (BJD 2460738.8245), respectively. The mean RV is $19.847\pm 0.066$ km s-1. 44B (BD+26 2522) was observed twice at McDonald Observatory and WISE Observatory. RVs of $19.314\pm 0.042$ km s-1 and $19.117\pm 0.054$ km s-1 were measured on 2025 February 8 (BJD 2460715.83734) and 2025 March 5 (BJD 2460740.82101), respectively. The mean RV is $19.215\pm 0.048$ km s-1.
47A (BD+08 2889) was observed twice at SAAO and WISE Observatory. RVs of $17.980\pm 0.025$ km s-1 and $17.948\pm 0.033$ km s-1 were measured on 2025 February 11 (BJD 2460718.58422) and 2025 February 28 (BJD 2460735.77632), respectively. The mean RV is $17.964\pm 0.029$ km s-1. 47B (BD+08 2887) was observed twice at WISE Observatory. RVs of $17.954\pm 0.085$ km s-1 and $17.921\pm 0.055$ km s-1 were measured on 2025 February 11 (BJD 2460718.83524) and 2025 February 26 (BJD 2460733.78058), respectively. The mean RV is $17.937\pm 0.071$ km s-1.
48A (HD 129171) was observed twice at McDonald Observatory. RVs of $11.419\pm 0.022$ km s-1 and $11.393\pm 0.030$ km s-1 were measured on 2024 December 23 (BJD 2460668.97809) and 2025 February 5 (BJD 2460712.89457), respectively. The mean RV is $11.406\pm 0.026$ km s-1. 48B (HD 129209) was observed twice at McDonald Observatory. RVs of $11.129\pm 0.032$ km s-1 and $11.118\pm 0.032$ km s-1 were measured on 2024 December 14 (BJD 2460660.018) and 2025 February 5 (BJD 2460712.99954), respectively. The mean RV is $11.123\pm 0.032$ km s-1.
Appendix B Description of MAROON-X observations and data reduction of radial velocities
[…]
Table 5: MAROON-X measurements of relative radial velocities in 6 wide binaries
[…]
Name Gaia DR3 identifier Sepaa2D separation: Sky-plane physical separation from El-Badry et al. (2021). MassbbStellar mass based on the mass-magnitude relation derived by Chae (2023). ruwe Spec Type RV(DR3) $v_{r}$(MAROON-X)ccRelative radial velocity $v_{r}(\equiv{\rm RV}_{A}-{\rm RV}_{B})$.
[kau] $M_{\odot}$] [km s-1] [km s-1]
TYC 4264-485-1 2201661297490051968 8.370 1.02 0.963 $12.734\pm 0.251$ $0.0182\pm 0.006$
TYC 4264-252-1 2201661091331626752 0.999 0.860 $12.387\pm 0.268$
TYC 2700-274-1 1871558941576158464 7.750 0.791 0.877 $22.921\pm 0.291$ $0.0239\pm 0.009$
TYC 2700-210-1 1871559697490418816 0.771 0.992 $23.347\pm 0.326$
TYC 2743-1157-1 1902676117063910016 6.271 0.899 1.282 $3.745\pm 0.365$ $0.9863\pm 0.005$
LSPM J2231+3454 1902679033343158528 0.879 1.271 $3.330\pm 0.390$
TYC 3640-273-1 1942384773344557184 9.209 0.84 0.956 $8.452\pm 0.202$ $0.2663\pm 0.084$
PM J23201+4819 1942384872124424832 0.565 1.116 $9.100\pm 0.238$
TYC 1101-86-1 1762461893163118464 8.828 0.941 0.931 $75.799\pm 0.252$ $0.2199\pm 0.004$
TYC 1101-87-1 1762461309047562368 0.939 1.156 $76.334\pm 0.238$
BD+00 4500 4230699363889120128 5.247 1.091 1.124 G0 $38.226\pm 0.165$ $0.2781\pm 0.022$
BD+00 4497 4230699329529382400 0.878 0.949 K8 $38.244\pm 0.514$
Appendix C Selection of wide binaries from the Scarpa et al. sample
Here we describe our selection of wide binary candidates from the Scarpa et al. (2017) sample of 58 pairs that have precise RVs measured in January and December 2013 with the fiber-fed FIES Echelle spectrograph at the 2.5m Nordic Optical Telescope, the Roque de Los Muchachos observatory in the Canary islands. Although this sample is far from a well-defined sample of wide binaries and contains many unbound pairs, we consider it because it contains true wide binaries with very precise RVs.
Since we eventually need only uncontaminated pure binaries, we remove obviously unbound, contaminated, or problematic cases by individually examining the pairs. Specifically, we exclude any pair in which (1) Gaia’s ruwe has unacceptably large value(s) (e.g., >1.5>1.5), (2) RV(s) show(s) unacceptably large variation(s) in time or with respect to Gaia DR3, (3) the scalar relative RV ($|v_{r}|$) is too large (e.g., $|v_{r}|>2$ km s-1) indicating an obviously gravitationally-unbound system, or (4) the sky-plane 2D physical separation is too large >0.22>0.22 pc or >45>45 kau, because any systems beyond the limit are more likely to be gravitationally unbound and require much more precise velocities. We are left with only 24 systems, which are listed in Table 6. All 24 systems are included in the raw sample of 306 systems shown in Figure 2.
Although the majority of the Scarpa et al. (2017) sample has already been excluded, it turns out that the remaining sample is relatively more contaminated than our other samples such as LCO and MAROON-X described in Appendix A and B, which are selected from previous statistical samples used by one of us (e.g., Chae 2024b). This may be because the above exclusion criterion for $|v_{r}|$ is not tight enough, so that boosted velocities due to unseen perturbers may be present. Indeed, as shown in Section III.3.1, the Hipparcos-Gaia relative PM comparison identified three cases of unacceptably large variation between the two observations, and they all are from the Scarpa et al. (2017) sample listed in Table 6.
The relatively high fraction of contaminated cases can be seen by the scalar 3D velocity $v_{\rm obs}$ listed in the penultimate column of Table 6. There are 9 systems with $v_{\rm obs}>1$ km s-1 including 3 systems with $v_{\rm obs}>1.5$ km s-1. Three of them are the above contaminated cases identified by the Hipparcos-Gaia test: HIP54692/HIP54681 ($v_{\rm obs}=1.690$), HIP65602/HIP65574 (1.053), and HIP101916/HIP101932 (1.632). In contrast, the clean statistical sample does not contain any case with $v_{\rm obs}>1$ km s-1 (see Table 3 for systems with relatively large values of $v_{\rm obs}$ among the clean sample).
The systems in Table 6 can be further tested with the relative velocity threshold of Equation (A1) for $v_{r}$ and $v_{p}$. The velocity $v_{p}$ is also considered because the relative PMs were not used in selecting the systems. Three systems fail this test, as indicated in the last column of Table 6. These systems are excluded from gravity tests in this work. These systems exactly match the three systems with $v_{\rm obs}>1.5$ km s-1, and two of them are the contaminated cases identified by the Hipparcos-Gaia test. Thus, 4 out of the 9 systems with $v_{\rm obs}>1$ km s-1 are excluded by the velocity threshold and the Hipparcos-Gaia test. It is unclear whether the remaining 5 cases with $1<v_{\rm obs}<1.5$ km s-1 from the Scarpa et al. (2017) sample are also contaminated. It will be interesting to see what future data will reveal about these systems.
Despite the relatively high fraction of contaminated cases, the Scarpa et al. (2017) sample is useful in this study because it provides independent measurements at different epochs for true pure binaries. Indeed, for 7 systems from the clean sample the Scarpa et al. (2017) values confirm the stability of $v_{r}$ in conjunction with other measurements over more than several years (see Table 1). Also, as demonstrated above, the sample inadvertently provided useful negative examples of contaminated cases through the Hipparcos-Gaia test and the velocity threshold. The Scarpa et al. (2017) sample seems to provide a useful testbed for kinematically contaminated cases.
[…]
Table 6: Wide binaries selected from Scarpa et al. (2017)
[…]
Name Gaia DR3 identifier Sepaa2D separation: Sky-plane physical separation from El-Badry et al. (2021). MassbbStellar mass based on the mass-magnitude relation derived by Chae (2023). RV(DR3)ccRadial velocity from Gaia DR3. RV(Scarpa)ddRadial velocity from Scarpa et al. (2017). $v_{r}$eeRelative radial velocity $v_{r}(\equiv{\rm RV}_{A}-{\rm RV}_{B})$ from Scarpa et al. (2017). $v_{p}$ffScalar sky-plane 2D velocity $v_{p}(\equiv\sqrt{v_{x^{\prime}}^{2}+v_{y^{\prime}}^{2}}$) from Gaia DR3. $v_{\rm obs}$ggScalar 3D velocity $v_{\rm obs}(\equiv\sqrt{v_{p}^{2}+v_{r}^{2}}$). TesthhTest with $v_{r}$ and $v_{p}$ based on the threshold given by Equation (A1).
[kau] $M_{\odot}$] [km s-1] [km s-1] [km s-1] [km s-1] [km s-1]
HIP11137 76300510625993344 2.090 0.998 $27.138\pm 0.175$ $27.129\pm 0.014$ $0.157\pm 0.021$ $0.535\pm 0.007$ $0.558\pm 0.009$ Pass
HIP11134 76300476266255488 0.925 $27.010\pm 0.139$ $27.286\pm 0.016$
HIP15304 10584899657116672 7.340 1.162 $31.392\pm 0.133$ $31.860\pm 0.019$ $0.856\pm 0.027$ $0.547\pm 0.009$ $1.016\pm 0.023$ Pass
HIP15310 10608573516849536 1.092 $32.068\pm 0.143$ $32.716\pm 0.019$
HIP15527 5060104351007433472 9.082 0.960 $39.849\pm 0.118$ $40.287\pm 0.019$ $0.377\pm 0.024$ $0.225\pm 0.004$ $0.439\pm 0.020$ Pass
HIP15526 5060105897197110144 0.880 $40.270\pm 0.120$ $40.664\pm 0.014$
HIP19859 3285218186904332288 1.417 1.070 $7.277\pm 0.137$ $6.858\pm 0.012$ $0.572\pm 0.020$ $0.966\pm 0.005$ $1.122\pm 0.011$ Pass
HIP19855 3285218255623808640 0.960 $8.010\pm 0.130$ $7.430\pm 0.016$
HIP21537 3230677565443833088 4.972 1.160 $38.397\pm 0.135$ $38.947\pm 0.016$ $0.135\pm 0.022$ $0.265\pm 0.013$ $0.298\pm 0.015$ Pass
HIP21534 3230677874682668672 1.150 $38.724\pm 0.133$ $39.082\pm 0.015$
HIP22611 4873223829966552192 5.959 1.503 $45.755\pm 0.120$ $46.203\pm 0.015$ $0.241\pm 0.022$ $0.661\pm 0.006$ $0.704\pm 0.009$ Pass
HIP22604 4873226853623529856 0.997 $45.859\pm 0.165$ $46.444\pm 0.016$
HIP25278 3400292798990117888 10.311 1.135 $37.701\pm 0.138$ $38.350\pm 0.047$ $0.253\pm 0.052$ $0.106\pm 0.006$ $0.274\pm 0.048$ Pass
HIP25220 3394298532176344960 0.738 $38.032\pm 0.124$ $38.603\pm 0.023$
HIP33705 5607190344506642432 12.331 1.150 $16.441\pm 0.125$ $16.700\pm 0.027$ $0.441\pm 0.032$ $0.100\pm 0.005$ $0.452\pm 0.031$ Pass
HIP33691 5607189485513198208 0.880 $16.726\pm 0.141$ $17.141\pm 0.017$
HIP34426 3359808231100381312 8.189 1.106 $1.840\pm 0.141 11$$1.544\pm 0.013$ $0.791\pm 0.017$ $1.612\pm 0.007$ $1.796\pm 0.010$ Fail
HIP34407 3359820016490648576 1.086 $12.835\pm 0.153$ $12.335\pm 0.011$
HIP39457 5595858262287843840 4.421 1.200 $26.293\pm 0.139$ $26.807\pm 0.015$ $0.741\pm 0.022$ $0.791\pm 0.008$ $1.084\pm 0.016$ Pass
HIP39452 5595858159201264768 1.016 N/A $27.548\pm 0.016$
HIP44858 692119656035933568 2.524 1.014 $29.976\pm 0.139$ $30.473\pm 0.010$ $0.292\pm 0.015$ $0.565\pm 0.007$ $0.636\pm 0.009$ Pass
HIP44864 692120029700390912 1.013 $30.380\pm 0.129$ $30.765\pm 0.011$
HIP45836 1019361632454363904 6.732 1.206 $895\pm 0.128 7.7$$761\pm 0.012$ $1.337\pm 0.016$ $0.673\pm 0.004$ $1.497\pm 0.014$ Pass
HIP45859 1019174509319377536 0.919 $6.780\pm 0.127$ $6.424\pm 0.010$
HIP52787 3550081879381593728 7.870 0.820 $23.759\pm 0.148$ $24.212\pm 0.019$ $0.410\pm 0.029$ $0.228\pm 0.005$ $0.469\pm 0.026$ Pass
HIP52776 3550084490721711872 0.700 $24.451\pm 0.180$ $24.622\pm 0.022$
HIP54692 777967084390189696 6.185 1.194 $1.222\pm 0.132 11$$1.696\pm 0.011$ $1.353\pm 0.015$ $1.012\pm 0.008$ $1.690\pm 0.013$ Fail
HIP54681 777967702865481344 0.996 $9.949\pm 0.152$ $10.343\pm 0.010$
HIP58067 3975129194660883328 2.893 0.956 $5.965\pm 0.198$ $6.461\pm 0.015$ $0.141\pm 0.021$ $0.202\pm 0.006$ $0.246\pm 0.013$ Pass
HIP58073 3975223065466473216 0.927 $5.801\pm 0.140$ $6.320\pm 0.014$
HIP64057 3945118265299248128 1.458 0.948 $1.565\pm 0.130$ $1.247\pm 0.014$ $0.017\pm 0.019$ $0.361\pm 0.009$ $0.362\pm 0.009$ Pass
HIP64059 3945118643256370688 0.896 $1.673\pm 0.147$ $1.264\pm 0.013$
HIP65602 6193279279612173952 9.539 0.816 $13.492\pm 0.144$ $12.884\pm 0.018$ $1.014\pm 0.025$ $0.282\pm 0.004$ $1.053\pm 0.025$ Pass
HIP65574 6193280031230266752 0.811 $2.211\pm 0.128 11$$1.870\pm 0.018$
HIP71726 1282815063829295360 16.772 1.030 $1.925\pm 0.121 11$$1.468\pm 0.015$ $0.170\pm 0.021$ $0.142\pm 0.008$ $0.221\pm 0.017$ Pass
HIP71737 1282817022334383232 1.000 $1.598\pm 0.124 11$$1.298\pm 0.015$
HIP74442 1274568245587206016 2.293 1.152 $60.315\pm 0.141$ $60.207\pm 0.016$ $0.962\pm 0.023$ $0.330\pm 0.007$ $1.017\pm 0.022$ Pass
HIP74439 1274562370068531712 0.975 $59.383\pm 0.174$ $59.245\pm 0.016$
HIP85620 1440518669436791296 8.824 1.094 $4.155\pm 0.127 33$$3.736\pm 0.027$ $0.442\pm 0.034$ $0.139\pm 0.006$ $0.463\pm 0.033$ Pass
HIP85575 1440425863783337856 0.970 $3.661\pm 0.132 33$$3.294\pm 0.021$
HIP99729 4249652990144051840 2.736 1.060 $464\pm 0.129 0.0$$11\pm 0.018 0.0$$061\pm 0.026$ $0.834\pm 0.010$ $0.836\pm 0.010$ Pass
HIP99727 4249652783985617920 1.050 $388\pm 0.138 0.0$$072\pm 0.019$
HIP101082 2298101352139398144 13.722 1.949 $14.467\pm 0.118$ $14.069\pm 0.014$ $0.322\pm 0.020$ $0.127\pm 0.010$ $0.346\pm 0.019$ Pass
HIP101166 2298101901895214720 1.042 $14.117\pm 0.149$ $13.747\pm 0.014$
HIP101916 1754191435419155456 6.401 1.566 $54.097\pm 0.124$ $53.457\pm 0.015$ $0.339\pm 0.022$ $1.596\pm 0.010$ $1.632\pm 0.011$ Fail
HIP101932 1754191229260708736 0.841 $53.510\pm 0.136$ $53.118\pm 0.016$
HIP118254 2882262637207289216 4.703 1.041 $29.940\pm 0.128$ $30.368\pm 0.016$ $0.747\pm 0.021$ $0.280\pm 0.004$ $0.798\pm 0.020$ Pass
HIP118251 2882262529831237120 0.973 $29.271\pm 0.128$ $29.621\pm 0.014$
Appendix D Selection of wide binaries with radial velocities from SDSS4 APOGEE
The SDSS4 APOGEE survey measured RVs for a large number of stars using high-resolution, multi-object, near-infrared spectroscopy. APOGEE RVs are available on the SDSS DR17 website https://www.sdss4.org/dr17/irspec/use-radial-velocities/. The precision reported for an individual RV is often better than 100m s-1 and in some cases better than 50m s-1. Thus, APOGEE RVs potentially provide precise relative RVs for a number of wide binaries. The file allStar-dr17-synspec-rev1.fits downloadable from the website provides RVs for 733901 stars, 717925 of which have Gaia DR3 identifications.
We search for wide binaries from the 717925 stars with Gaia DR3 identifications. Wide binary candidates are selected from the El-Badry et al. (2021) catalog. Once a candidate binary with APOGEE RVs for both stars is found, we check whether the radial velocity difference between the pair can be consistent with a gravitationally bound system. We use the threshold given by Equation (A1). We have found 208 wide binaries within 300 pc from the Sun, which are listed in Table 7. For 15 of them, RVs for both components were measured multiple times at multiple epochs. For these systems, Table 7 gives only the error-weighted mean and its error, while Table 8 gives all the measured values.
Each measurement of RV for a star is based on NVISITS ($\geq 1$) “visit(s)” of the star. The given value of RV refers to a signal-to-noise ratio (SNR)-weighted average if NVISITS >1>1 stored in the APOGEE parameter VHELIO_AVG. The given nominal error (Err) refers to the SNR-weighted uncertainty stored in the APOGEE parameter VERR. When multiple observations (each of which may have multiple visits) were made at multiple epochs, the given value of RV and its error refer to the error-weighted mean of RVs and its uncertainty. For each value of RV, two additional errors are given in the table: ‘Disp’ and ‘Scatt’. Disp simply refers to the standard deviation of RVs from multiple observations at multiple epochs. It is set to zero when only a single-epoch observation (regardless of NVISITS) was made. Scatt refers to the scatter in multiple visits in an observation stored in the APOGEE parameter VSCATTER. When multiple observations were made, the given value of Scatt refers to the maximum value. Scatt is zero when there was only a single observation with a single visit. Disp is provided as a qualitative check of the variability of RV when multiple observations are available while Scatt is used to select reliable RVs. Because Scatt is not available for all RVs, the nominal error ‘Err’ is used in estimating the nominal uncertainty of the relative RV $v_{r}$ between the two stars. Note that Err is either smaller or larger than Scatt as can be seen in Table 7. The majority of wide binaries have $\sigma_{v_{r}}<0.1$ km s-1 and 195 wide binaries with $\sigma_{v_{r}}<0.35$ km s-1 are included in the raw sample of the main part shown in Figure 4.
[…]
Table 7: Summary of wide binaries with radial velocities from the SDSS DR17 APOGEE database
[…]
identifier ssaaSky-plane physical separation from El-Badry et al. (2021). MassbbStellar mass based on the mass-magnitude relation derived by Chae (2023). RVccRadial velocity stored in the APOGEE parameter VHELIO_AVG (the barycentric velocity) which represents a signal-to-noise weighted average if multiple visits took place (i.e., the APOGEE parameter NVISITS >1>1): When multiple values from multiple observations (note here that each observation may have multiple visits) are reported, the given value refers to the error-weighted mean $\mu$ calculated by $\mu=\sum({\rm RV}_{i}/\sigma_{i}^{2})/\sum(1/\sigma_{i}^{2})$. ErrddReported error: When multiple values from multiple observations are reported, the given value refers to the error of $\mu$ calculated by $\sigma_{\mu}=1/\sqrt{\sum(1/\sigma_{i}^{2})}$. DispeeThis value refers to the standard deviation of multiple values from multiple observations. It is set to zero when a single observation was made. ScattffThe APOGEE parameter VSCATTER which represents the scatter from multiple visits of an observation. It is set to zero when an observation consists of a single visit. When multiple values of VSCATTER from observations are available, the maximum value is given. $v_{r}$ggRelative radial velocity $v_{r}(\equiv{\rm RV}_{A}-{\rm RV}_{B})$ where ${\rm RV}_{A}$ and ${\rm RV}_{B}$ refer respectively to the first and second values. $\sigma_{v_{r}}$ multi-epochhhWhether multiple observations were made at multiple epochs.
[kau] $M_{\odot}$] [km s-1] [km s-1] [km s-1] [km s-1] [km s-1] [km s-1]
Gaia DR3 5961072491289480960 4.449 0.537 -1.252 0.054 0.000 0.000 -0.067 0.081
Gaia DR3 5961072456929737216 0.524 -1.185 0.060 0.000 0.014
Gaia DR3 466294295706341760 6.201 1.025 -1.956 0.034 0.000 0.000 -0.660 0.065
Gaia DR3 466291787445425408 0.392 -1.296 0.055 0.000 0.240
Gaia DR3 750184379766219648 10.229 0.790 19.668 0.026 0.000 0.053 -0.188 0.125
Gaia DR3 750184311046743168 0.688 19.856 0.122 0.000 0.089
Gaia DR3 6619469571588636288 3.274 0.626 -8.789 0.060 0.000 0.077 0.134 0.102
Gaia DR3 6619469571588635904 0.497 -8.923 0.082 0.000 0.107
Gaia DR3 2344436971156155904 16.197 1.032 23.537 0.024 0.000 0.037 -0.398 0.069
Gaia DR3 2344437246034051968 0.507 23.935 0.065 0.000 0.112
Gaia DR3 2342666791794987264 11.690 0.917 22.908 0.025 0.000 0.036 -0.380 0.112
Gaia DR3 2342666138959958400 0.387 23.288 0.109 0.000 0.252
Gaia DR3 607218804112579968 15.567 0.821 27.141 0.031 0.000 0.000 -0.341 0.050
Gaia DR3 607218765457083904 0.610 27.482 0.039 0.000 0.000
Gaia DR3 605713022937691136 2.744 1.551 26.605 0.167 0.000 0.000 0.174 0.188
Gaia DR3 605712812485668352 0.834 26.431 0.086 0.000 0.000
Gaia DR3 2642922251741817216 4.581 0.548 23.539 0.028 0.059 0.044 -0.156 0.040 Yes
Gaia DR3 2642922286101533952 0.528 23.695 0.029 0.042 0.062
Gaia DR3 5279010553386994048 13.361 0.486 44.107 0.036 0.000 0.000 -0.021 0.107
Gaia DR3 5278997561113174016 0.401 44.128 0.101 0.000 0.000
Gaia DR3 3053911383155592832 12.824 0.528 56.530 0.035 0.000 0.050 -0.225 0.054
Gaia DR3 3053907221325174144 0.498 56.755 0.041 0.000 0.068
Gaia DR3 2119132916774654080 2.522 1.242 -21.516 0.023 0.000 0.044 0.536 0.043
Gaia DR3 2119132916774654208 0.863 -22.052 0.036 0.000 0.030
Gaia DR3 2119404186908641792 1.828 0.637 -14.241 0.048 0.000 0.089 0.487 0.061
Gaia DR3 2119404186908642048 0.623 -14.728 0.037 0.000 0.000
Gaia DR3 2029433521248546304 2.849 1.041 -44.770 0.033 0.000 0.000 0.224 0.152
Gaia DR3 2029432043779954432 0.196 -44.994 0.148 0.000 0.154
Gaia DR3 5183999279727496448 18.627 1.275 -5.200 0.030 0.000 0.078 0.056 0.042
Gaia DR3 5183999309791637504 0.924 -5.256 0.030 0.000 0.018
Gaia DR3 5174322615330339456 4.858 0.714 18.562 0.028 0.000 0.030 -0.668 0.111
Gaia DR3 5174322855848505728 0.262 19.230 0.107 0.000 0.146
Gaia DR3 3840226230398314368 48.488 0.523 -4.704 0.056 0.000 0.113 -0.140 0.085
Gaia DR3 3840219358450646656 0.494 -4.564 0.064 0.000 0.414
Gaia DR3 652232737138721280 23.846 0.590 97.005 0.050 0.114 0.000 -0.288 0.132
Gaia DR3 652239441584849280 0.586 97.293 0.122 0.000 0.000
Gaia DR3 48167959442012672 1.854 0.853 10.825 0.028 0.000 0.000 -0.848 0.046
Gaia DR3 48167955146062720 0.669 11.673 0.037 0.000 0.000
Gaia DR3 2501323948860951296 19.938 1.027 -26.327 0.027 0.031 0.066 -0.014 0.040 Yes
Gaia DR3 2501325525113734912 0.749 -26.313 0.030 0.064 0.115
Gaia DR3 2501911740905243392 3.483 0.817 53.014 0.027 0.000 0.000 -0.684 0.042
Gaia DR3 2501911740905243136 0.776 53.698 0.032 0.000 0.044
Gaia DR3 1356763886586884864 5.370 0.765 -36.526 0.021 0.039 0.029 0.177 0.043
Gaia DR3 1356763852227145344 0.704 -36.703 0.037 0.000 0.003
Gaia DR3 1352381027080312832 24.561 1.236 -25.485 0.020 0.008 0.010 -0.151 0.039
Gaia DR3 1352427996842670464 0.698 -25.334 0.033 0.000 0.023
Gaia DR3 1357040379401720448 1.506 0.929 -21.890 0.028 0.000 0.010 -0.137 0.040
Gaia DR3 1357040345041982464 0.701 -21.753 0.028 0.605 0.087
Gaia DR3 2125816160770555904 2.111 0.800 11.132 0.042 0.000 0.045 0.659 0.379
Gaia DR3 2125816160770558592 0.431 10.473 0.377 0.000 2.102
Gaia DR3 3861880665230888960 19.322 1.301 13.512 0.043 0.000 0.017 0.386 0.051
Gaia DR3 3858878375716469120 0.922 13.126 0.027 0.000 0.051
Gaia DR3 3861953095559329536 29.682 0.740 14.441 0.039 0.000 0.010 -0.301 0.058
Gaia DR3 3861952369709297536 0.720 14.742 0.043 0.000 0.007
Gaia DR3 217107577453306112 3.417 0.835 22.261 0.020 0.013 0.144 0.433 0.030 Yes
Gaia DR3 217107607515576704 0.660 21.828 0.022 0.045 0.056
Gaia DR3 665747663485734016 4.126 0.748 -0.190 0.036 0.000 0.065 -0.491 0.062
Gaia DR3 665747663485733632 0.627 0.301 0.050 0.000 0.145
Gaia DR3 663509878149021824 33.873 0.854 16.977 0.017 0.040 0.000 0.070 0.043
Gaia DR3 663515861040979584 0.607 16.907 0.039 0.000 0.000
Gaia DR3 666293742807553024 2.991 0.862 21.017 0.037 0.000 0.013 -0.578 0.279
Gaia DR3 666293777167291008 0.411 21.595 0.277 0.000 0.000
Gaia DR3 664055515090902144 1.352 0.363 26.148 0.056 0.000 0.000 0.120 0.096
Gaia DR3 664055412011687424 0.354 26.028 0.078 0.000 0.000
Gaia DR3 664315652670220032 2.429 0.438 -2.690 0.123 0.000 0.848 -0.392 0.193
Gaia DR3 664315652670219776 0.424 -2.298 0.149 0.000 0.525
Gaia DR3 2099859268216511488 0.337 0.428 -17.723 0.064 0.000 0.063 0.512 0.099
Gaia DR3 2099859268216512000 0.385 -18.235 0.075 0.000 0.000
Gaia DR3 2100011516217168640 0.810 0.566 -42.882 0.051 0.000 0.118 0.510 0.110
Gaia DR3 2100011516217169920 0.553 -43.392 0.098 0.000 0.000
Gaia DR3 2100451630105041152 2.030 1.203 -7.960 0.043 0.000 0.049 -0.757 0.055
Gaia DR3 2100451630105040256 0.790 -7.203 0.035 0.000 0.107
Gaia DR3 2100501760964021504 17.569 1.060 10.209 0.030 0.000 0.000 0.184 0.040
Gaia DR3 2100502001482188032 1.002 10.025 0.027 0.000 0.000
Gaia DR3 3811688440459523200 2.030 0.946 7.522 0.045 0.000 0.068 -0.907 0.082
Gaia DR3 3811688440459523456 0.593 8.429 0.068 0.000 0.000
Gaia DR3 638830003928974336 0.365 0.540 14.947 0.131 0.000 5.215 2.290 0.181
Gaia DR3 638829999633204864 0.426 12.657 0.125 0.000 4.541
Gaia DR3 2532890137420583168 20.057 0.533 -22.431 0.393 0.000 0.000 2.332 1.010
Gaia DR3 2532890858975088512 0.470 -24.763 0.930 0.000 0.000
Gaia DR3 2532980366093464832 28.332 0.968 9.318 0.059 11.282 0.000 0.828 0.279
Gaia DR3 2532980499236841984 0.404 8.490 0.273 0.000 0.000
Gaia DR3 2534384064484982016 1.217 0.293 6.042 0.155 0.000 0.079 -0.265 0.217
Gaia DR3 2534384064484982400 0.279 6.307 0.152 0.000 0.152
Gaia DR3 2534623758019863296 12.474 0.447 0.291 0.057 0.000 0.000 -0.035 0.092
Gaia DR3 2534624685732797824 0.386 0.326 0.072 0.000 0.025
Gaia DR3 2507987436001873408 6.302 1.081 -3.735 0.022 0.039 0.087 0.576 0.043
Gaia DR3 2507987431706333184 0.967 -4.311 0.037 0.000 0.056
Gaia DR3 2435485982863717120 1.885 0.999 -28.884 0.025 0.000 0.000 0.296 0.044
Gaia DR3 2435485982863716992 0.640 -29.180 0.036 0.000 0.057
Gaia DR3 2435551609963991936 27.394 1.537 -3.344 0.021 0.044 0.131 -0.916 0.330
Gaia DR3 2435554599261229824 0.296 -2.428 0.329 0.000 0.581
Gaia DR3 2128304183783972992 0.504 0.960 -21.532 0.032 0.000 0.085 -1.257 0.049
Gaia DR3 2128304179486008704 0.816 -20.275 0.037 0.000 0.072
Gaia DR3 2126777988630768000 1.792 1.059 -15.875 0.061 0.000 0.082 -0.297 0.080
Gaia DR3 2126777992924832128 0.912 -15.578 0.051 0.000 0.052
Gaia DR3 4088301956580684416 13.816 0.562 -11.303 0.082 0.000 0.000 0.524 0.243
Gaia DR3 4088301823454290176 0.354 -11.827 0.229 0.000 0.000
Gaia DR3 2783518418493208832 0.165 0.755 -2.501 0.052 0.000 0.433 0.908 0.078
Gaia DR3 2783518212334778880 0.700 -3.409 0.058 0.000 0.000
Gaia DR3 315289980082496000 3.346 0.730 12.902 0.026 0.000 0.043 -0.869 0.061
Gaia DR3 315289980082496256 0.389 13.771 0.055 0.000 0.262
Gaia DR3 3641315024925921024 2.553 1.003 -15.066 0.025 0.000 0.000 0.192 0.033
Gaia DR3 3641315024927514880 0.703 -15.258 0.022 0.071 0.094
Gaia DR3 3975013093104939904 1.935 0.400 28.683 0.074 0.000 0.174 0.005 0.344
Gaia DR3 3975013093104939648 0.210 28.678 0.336 0.000 1.055
Gaia DR3 3961746072270936320 27.208 0.964 -45.842 0.013 0.078 0.154 -0.091 0.021 Yes
Gaia DR3 3961745217573223296 0.720 -45.751 0.017 0.071 0.129
Gaia DR3 3888309828090932608 2.493 0.479 -9.621 0.084 0.000 0.686 0.041 0.113
Gaia DR3 3888309832385988992 0.432 -9.662 0.075 0.000 0.406
Gaia DR3 3888520938618115456 1.114 0.386 64.304 0.225 0.000 0.453 0.320 0.536
Gaia DR3 3888520938618496768 0.285 63.984 0.487 0.000 2.111
Gaia DR3 2734917221406795648 1.600 1.017 -2.107 0.027 0.000 0.028 -0.691 0.059
Gaia DR3 2734917221406795520 0.482 -1.416 0.052 0.075 0.480
Gaia DR3 6823356823090355712 9.032 1.281 -20.233 0.041 0.000 0.066 -0.033 0.063
Gaia DR3 6823356715715859840 1.253 -20.200 0.048 0.000 0.072
Gaia DR3 6827446284791133184 3.608 0.875 -46.987 0.025 0.000 0.021 0.776 0.069
Gaia DR3 6827446280496122112 0.601 -47.763 0.064 0.000 0.000
Gaia DR3 2778254433563650048 1.213 0.794 2.065 0.026 0.000 0.006 -0.196 0.049
Gaia DR3 2778254437855626240 0.566 2.261 0.041 0.000 0.004
Gaia DR3 3016902642094840704 0.697 0.844 19.935 0.040 0.145 0.538 0.553 0.057
Gaia DR3 3016902646391459328 0.755 19.382 0.041 0.000 0.000
Gaia DR3 1261431280655790208 17.706 0.699 -31.270 0.102 0.000 0.674 -0.021 0.236
Gaia DR3 1261431589893436288 0.556 -31.249 0.213 0.000 1.113
Gaia DR3 1262819482805829760 2.507 0.341 -50.351 0.429 0.000 0.000 0.427 0.451
Gaia DR3 1262819551525306624 0.258 -50.778 0.139 0.000 0.235
Gaia DR3 2687603311918406528 11.324 0.985 12.188 0.034 0.000 0.021 0.090 0.053
Gaia DR3 2687603208839189376 0.626 12.098 0.041 0.023 0.111
Gaia DR3 2242465891977167872 1.233 1.330 -2.749 0.189 0.000 0.194 0.239 0.192
Gaia DR3 2242465891977204992 0.916 -2.988 0.031 0.000 0.019
Gaia DR3 1291638988239949056 7.655 0.681 -41.205 0.039 0.000 0.007 -0.025 0.146
Gaia DR3 1291638988239948800 0.353 -41.180 0.141 0.000 0.217
Gaia DR3 1292334704222207104 1.923 0.578 -5.693 0.040 0.000 0.118 -0.343 0.237
Gaia DR3 1292334704222207232 0.248 -5.350 0.234 0.000 1.037
Gaia DR3 2603390513756129152 22.756 1.150 -10.734 0.038 0.000 0.000 -0.208 0.073
Gaia DR3 2603390410676914816 0.729 -10.526 0.062 0.000 0.000
Gaia DR3 2615313579192094080 0.983 0.928 -16.558 0.028 0.000 0.018 0.703 0.360
Gaia DR3 2615313579193124992 0.222 -17.261 0.359 0.000 0.276
Gaia DR3 2601074839188320384 3.130 0.925 -1.307 0.033 0.000 0.053 -0.356 0.046
Gaia DR3 2601074903612842880 0.908 -0.951 0.032 0.000 0.013
Gaia DR3 2615509842018541184 1.940 0.919 -25.133 0.032 0.000 0.182 -0.512 0.085
Gaia DR3 2615509842020990336 0.906 -24.621 0.079 0.000 0.000
Gaia DR3 2597900304305485952 22.582 0.938 6.717 0.045 0.000 0.299 -0.066 0.068
Gaia DR3 2597899930643579520 0.768 6.783 0.051 0.000 0.187
Gaia DR3 2613542407694482816 1.031 0.995 15.920 0.023 0.000 0.067 -1.237 0.045
Gaia DR3 2613542403398991360 0.952 17.157 0.039 0.000 0.000
Gaia DR3 2627346157705716352 2.975 0.873 22.895 0.035 0.000 0.092 0.145 0.057
Gaia DR3 2627346535662838272 0.654 22.750 0.045 0.000 0.041
Gaia DR3 2623865584928390784 21.862 1.219 -31.354 0.028 0.000 0.000 -0.052 0.040
Gaia DR3 2624616237837226368 1.051 -31.302 0.028 0.000 0.000
Gaia DR3 2543548528262965248 5.541 1.294 -21.561 0.033 0.000 0.025 -0.702 0.049
Gaia DR3 2543548528262965504 0.749 -20.859 0.036 0.000 0.048
Gaia DR3 2565313665876247168 6.931 0.963 -13.701 0.025 0.000 0.000 -0.308 0.044
Gaia DR3 2565313567092405888 0.784 -13.393 0.036 0.000 0.000
Gaia DR3 2581807955200716800 19.842 0.764 -2.898 0.029 0.000 0.000 -0.011 0.072
Gaia DR3 2581806619466249728 0.450 -2.887 0.066 0.000 0.000
Gaia DR3 2581883413481515008 17.044 1.241 -14.699 0.023 0.055 0.000 -0.310 0.039 Yes
Gaia DR3 2581859911420471808 0.741 -14.389 0.032 0.104 0.000
Gaia DR3 2582614416915244416 7.599 0.699 6.468 0.036 0.313 0.000 0.597 0.074
Gaia DR3 2582614412620059520 0.467 5.871 0.065 0.000 0.000
Gaia DR3 2588048478257634304 2.242 0.847 -5.600 0.039 0.000 0.077 0.347 0.049
Gaia DR3 2588048473962161280 0.738 -5.947 0.029 0.027 0.091
Gaia DR3 2588052051670468224 2.136 1.259 14.045 0.025 0.000 0.068 0.604 0.030
Gaia DR3 2588051948591253376 0.808 13.441 0.017 0.007 0.079
Gaia DR3 2594556109625136128 12.607 0.451 -8.800 0.050 0.000 0.000 0.128 0.092
Gaia DR3 2594555491149842432 0.373 -8.928 0.077 0.000 0.000
Gaia DR3 3696682204254353920 34.216 0.950 10.107 0.046 0.000 0.102 -0.224 0.064
Gaia DR3 3696681860656968064 0.899 10.331 0.044 0.000 0.017
Gaia DR3 3698627961878292864 7.590 1.007 2.092 0.025 0.000 0.075 0.087 0.036
Gaia DR3 3698627927518342144 0.861 2.005 0.026 0.245 0.108
Gaia DR3 2157622725756418816 12.259 0.992 -37.939 0.023 0.000 0.000 0.472 0.035
Gaia DR3 2157623451608162688 0.706 -38.411 0.027 0.000 0.000
Gaia DR3 765428249492127104 13.781 0.485 5.792 0.069 0.000 0.367 0.201 0.104
Gaia DR3 765417666692750080 0.332 5.591 0.078 0.000 0.155
Gaia DR3 785716030026546688 18.845 0.810 -15.570 0.030 0.026 0.086 -0.089 0.051 Yes
Gaia DR3 785716682861545344 0.655 -15.481 0.041 0.059 0.166
Gaia DR3 792284031455893376 5.031 0.973 -10.027 0.028 0.000 0.023 -0.274 0.046
Gaia DR3 792284031455893504 0.779 -9.753 0.037 0.000 0.006
Gaia DR3 782803458087840000 2.502 0.814 -27.082 0.022 0.000 0.000 0.641 0.040
Gaia DR3 782803389372882944 0.586 -27.723 0.033 0.000 0.097
Gaia DR3 783007245705158400 13.815 0.649 -15.864 0.026 0.198 0.039 -0.437 0.051
Gaia DR3 783007280064896128 0.591 -15.427 0.044 0.000 0.119
Gaia DR3 822171987312735232 9.197 0.594 11.361 0.025 0.017 0.035 -0.189 0.253
Gaia DR3 822172227830881408 0.311 11.550 0.252 0.000 0.908
Gaia DR3 4418851263967008000 4.959 0.900 -6.898 0.022 0.000 0.051 -0.056 0.298
Gaia DR3 4418850890305439872 0.225 -6.842 0.297 0.000 1.656
Gaia DR3 1500748301498613248 3.249 1.626 -15.431 0.104 0.000 2.872 1.014 0.145
Gaia DR3 1500748198419397760 0.519 -16.445 0.101 0.000 0.000
Gaia DR3 1500827225817637504 3.446 0.177 -9.750 0.211 0.000 0.866 -0.447 0.353
Gaia DR3 1500826882220275712 0.139 -9.303 0.283 0.000 0.412
Gaia DR3 1508340498008710784 8.226 1.277 -36.941 0.019 0.055 0.000 -0.502 0.102
Gaia DR3 1508340532367957376 0.528 -36.439 0.100 0.000 0.049
Gaia DR3 1466043419558823296 15.233 0.639 -2.255 0.037 0.000 0.131 -0.204 0.079
Gaia DR3 1466040395901846528 0.446 -2.051 0.070 0.000 0.183
Gaia DR3 1460823041429644160 1.987 0.939 -8.137 0.010 0.069 0.078 0.414 0.056
Gaia DR3 1460823041429644032 0.620 -8.551 0.055 0.000 0.132
Gaia DR3 1460984429120711808 1.519 0.858 -54.308 0.028 0.080 0.077 0.082 0.079
Gaia DR3 1460984429120711552 0.717 -54.390 0.074 0.000 0.159
Gaia DR3 1461182444292609536 3.719 1.105 -13.677 0.024 0.007 0.073 -0.489 0.030 Yes
Gaia DR3 1461182684811100288 1.084 -13.188 0.018 0.038 0.048
Gaia DR3 1469596938060496768 3.735 0.617 -48.401 0.156 0.000 0.043 -0.352 0.160
Gaia DR3 1469596938060497024 0.508 -48.049 0.036 0.089 0.349
Gaia DR3 1469606455707961088 19.502 0.997 -12.462 0.019 0.010 0.058 -0.185 0.026 Yes
Gaia DR3 1469606283909297152 0.905 -12.277 0.018 0.080 0.045
Gaia DR3 1475867727751525632 5.102 0.483 2.719 0.063 0.000 0.559 -0.236 0.248
Gaia DR3 1475867727751525504 0.279 2.955 0.240 0.000 1.262
Gaia DR3 1459718891236328192 21.967 0.664 -20.889 0.012 0.076 0.145 -0.019 0.065
Gaia DR3 1459715764500136192 0.449 -20.870 0.064 0.000 0.186
Gaia DR3 1476646250703083264 4.150 0.545 -10.535 0.056 0.000 0.000 -0.172 0.092
Gaia DR3 1476646285062821888 0.445 -10.363 0.073 0.000 0.337
Gaia DR3 1476762627137479552 15.898 0.540 -1.850 0.058 0.000 0.000 -0.512 0.152
Gaia DR3 1476768433933264128 0.320 -1.338 0.141 0.000 0.211
Gaia DR3 1593535881608710784 8.875 0.366 -0.661 0.065 0.000 0.000 0.027 0.105
Gaia DR3 1593537084199554560 0.357 -0.688 0.082 0.226 0.820
Gaia DR3 1604175580752706432 2.686 1.015 -14.773 0.028 0.000 0.027 1.616 0.461
Gaia DR3 1604175477673086592 0.224 -16.389 0.460 0.000 0.395
Gaia DR3 1590641417247939584 9.153 0.882 -6.554 0.042 0.000 0.050 0.252 0.065
Gaia DR3 1590641417247939328 0.766 -6.806 0.050 0.000 0.049
Gaia DR3 4008273560363584768 10.198 0.371 5.286 0.105 0.000 0.558 1.848 0.756
Gaia DR3 4008273903960972416 0.160 3.438 0.749 0.000 1.791
Gaia DR3 4008339427981993088 3.022 1.360 -12.339 0.036 0.000 0.087 0.238 0.080
Gaia DR3 4008339359262516224 0.679 -12.577 0.072 0.000 0.106
Gaia DR3 4008743086188165248 11.678 0.606 -9.820 0.061 0.000 0.033 0.804 0.343
Gaia DR3 4008743154907642496 0.272 -10.624 0.338 0.000 0.596
Gaia DR3 63739517993652864 4.950 0.684 -5.527 0.043 0.000 0.149 0.173 0.061
Gaia DR3 63739311835222784 0.676 -5.700 0.043 0.000 0.062
Gaia DR3 6246952489482132736 13.867 0.962 9.765 0.027 0.000 0.028 -0.017 0.038
Gaia DR3 6246952592561351040 0.755 9.782 0.027 0.020 0.013
Gaia DR3 1534089651581519488 18.119 0.903 3.515 0.026 0.000 0.016 -0.433 0.081
Gaia DR3 1534086421765263360 0.450 3.948 0.077 0.000 0.126
Gaia DR3 1535360373488627200 17.981 1.314 -11.257 0.019 0.055 0.053 -0.270 0.031 Yes
Gaia DR3 1535360343424477440 1.092 -10.987 0.024 0.026 0.038
Gaia DR3 1538594140265624064 1.287 0.977 -33.814 0.027 0.000 0.003 -0.886 0.054
Gaia DR3 1538594144560846848 0.639 -32.928 0.047 0.000 0.008
Gaia DR3 1533327415144066432 5.666 0.872 -2.323 0.043 0.000 0.022 -0.172 0.068
Gaia DR3 1533327483863543296 0.700 -2.151 0.053 0.000 0.054
Gaia DR3 1549029953017127808 1.932 0.929 9.866 0.037 0.000 0.023 1.070 0.054
Gaia DR3 1549030056098162176 0.860 8.796 0.040 0.000 0.075
Gaia DR3 5477438450481258240 2.497 0.878 8.484 0.030 0.000 0.000 -0.185 0.049
Gaia DR3 5477438454778666624 0.759 8.669 0.039 0.000 0.113
Gaia DR3 1305035575352515840 3.031 0.484 7.595 0.031 0.000 0.011 -0.049 0.084
Gaia DR3 1305036228187622912 0.277 7.644 0.078 0.000 0.099
Gaia DR3 1302543326089734144 14.656 0.998 -63.058 0.036 0.000 0.025 -0.049 0.052
Gaia DR3 1302449489644570496 0.848 -63.009 0.038 0.000 0.006
Gaia DR3 1320379809873587968 1.716 1.007 -129.817 0.039 0.000 0.004 0.040 0.053
Gaia DR3 1320379741154111104 0.807 -129.857 0.036 0.000 0.005
Gaia DR3 1321929846390816640 36.773 0.793 -9.730 0.022 0.016 0.033 0.250 0.053
Gaia DR3 1321930842823231616 0.603 -9.980 0.048 0.000 0.014
Gaia DR3 1555391177542038400 8.101 1.089 -4.412 0.017 0.045 0.029 -0.119 0.079
Gaia DR3 1555388222603962752 0.398 -4.293 0.077 0.000 0.199
Gaia DR3 1580778076392231424 10.792 0.641 -26.414 0.056 0.000 0.145 -0.142 0.085
Gaia DR3 1580778213830765824 0.527 -26.272 0.064 0.000 0.154
Gaia DR3 76974648692629376 1.397 0.483 45.236 0.032 0.025 0.019 -0.100 0.064
Gaia DR3 76974751771844096 0.422 45.336 0.055 0.000 0.011
Gaia DR3 1381408782592655616 2.363 0.812 -4.565 0.040 0.000 0.091 0.442 0.066
Gaia DR3 1381455722290898176 0.635 -5.007 0.053 0.000 0.045
Gaia DR3 1391176092275606656 1.832 1.079 -16.754 0.027 0.000 0.000 -0.566 0.092
Gaia DR3 1391176092275606528 0.663 -16.188 0.088 0.000 0.137
Gaia DR3 1331462302965590400 3.889 1.283 -23.271 0.052 0.528 0.434 0.986 0.082
Gaia DR3 1331462302965590144 0.472 -24.257 0.064 0.000 0.120
Gaia DR3 1332438183958298496 5.681 0.968 -22.793 0.040 0.000 0.317 -0.688 0.045
Gaia DR3 1332438188254551808 0.933 -22.105 0.021 0.014 0.075
Gaia DR3 1332473922382951424 0.653 0.634 -12.902 0.039 0.165 0.593 -0.310 0.409
Gaia DR3 1332473922382506624 0.529 -12.592 0.407 0.000 0.667
Gaia DR3 4572807566444752128 0.782 0.804 27.156 0.022 0.000 0.000 0.108 0.093
Gaia DR3 4572807566444751616 0.317 27.048 0.090 0.000 0.119
Gaia DR3 1397692244857993728 2.797 0.713 -34.692 0.033 0.000 0.000 0.838 0.042
Gaia DR3 1397692279217731456 0.693 -35.530 0.026 0.994 0.327
Gaia DR3 1403853152105526272 4.334 0.375 10.955 0.056 0.000 0.000 0.015 0.087
Gaia DR3 1403852430551017984 0.297 10.940 0.066 0.797 6.370
Gaia DR3 1407596954838724864 1.159 0.726 -9.612 0.042 0.000 0.043 -0.019 0.086
Gaia DR3 1407596954838646912 0.701 -9.593 0.075 0.000 0.044
Gaia DR3 1407786689313958784 1.351 0.636 -54.605 0.105 0.000 0.000 -0.749 0.149
Gaia DR3 1407786654954220160 0.472 -53.856 0.106 0.000 0.125
Gaia DR3 2076871091425586944 3.106 0.849 -25.476 0.026 0.000 0.000 0.198 0.033
Gaia DR3 2076871091425583232 0.785 -25.674 0.021 0.004 0.083
Gaia DR3 576837717289109248 1.839 1.060 40.557 0.027 0.000 0.036 -0.290 0.038
Gaia DR3 576837717289109504 0.886 40.847 0.027 0.000 0.031
Gaia DR3 4819204205221996160 2.454 0.692 44.578 0.062 0.000 0.079 -0.347 0.098
Gaia DR3 4819204200922632064 0.626 44.925 0.076 0.000 0.105
Gaia DR3 3221172596658836736 36.462 0.699 75.807 0.034 0.000 0.047 -0.059 0.059
Gaia DR3 3221172729803195904 0.552 75.866 0.048 0.000 0.044
Gaia DR3 1247610934890765312 0.720 0.791 6.407 0.048 0.000 0.099 -0.794 0.059
Gaia DR3 1247610934890765184 0.721 7.201 0.034 0.000 0.100
Gaia DR3 1240752559313435520 3.125 0.777 -18.826 0.026 0.000 0.019 0.004 0.072
Gaia DR3 1240752559313476224 0.379 -18.830 0.067 0.000 0.072
Gaia DR3 2116748247853607168 1.402 1.145 -8.060 0.070 0.000 0.000 -0.562 0.075
Gaia DR3 2116748243555689728 0.889 -7.498 0.028 0.000 0.000
Gaia DR3 2116892318236994816 1.981 1.119 7.002 0.038 0.000 0.000 0.056 0.050
Gaia DR3 2116892313938771584 0.738 6.946 0.033 0.000 0.000
Gaia DR3 2078886290079138560 1.333 1.004 -10.477 0.037 0.000 0.071 -0.246 0.069
Gaia DR3 2078886285778453632 0.917 -10.231 0.058 0.000 0.000
Gaia DR3 1022456139210632064 0.108 0.590 11.460 0.026 0.589 0.000 0.559 0.043
Gaia DR3 1022456104850892928 0.580 10.901 0.034 0.000 0.000
Gaia DR3 1029034448560364928 1.641 0.735 -24.567 0.118 0.000 0.000 -0.243 0.123
Gaia DR3 1029034448560364800 0.605 -24.324 0.035 0.000 0.000
Gaia DR3 145203159127518336 43.714 0.808 17.197 0.058 0.052 0.075 0.081 0.081
Gaia DR3 145203811962545152 0.608 17.116 0.056 0.000 0.219
Gaia DR3 152109054223716480 7.241 0.561 16.995 0.067 0.000 0.535 0.570 0.258
Gaia DR3 152108882425024128 0.258 16.425 0.249 0.000 0.413
Gaia DR3 157039397504919168 0.623 1.555 30.297 0.064 0.000 0.305 -0.160 0.068
Gaia DR3 157039401801742976 0.988 30.457 0.024 0.000 0.000
Gaia DR3 880933187236449792 45.245 0.990 -11.592 0.030 0.000 0.023 0.233 0.051
Gaia DR3 880933702632515840 0.669 -11.825 0.041 0.000 0.076
Gaia DR3 5262754514488149632 15.417 1.406 1.239 0.027 0.063 0.081 -0.049 0.048 Yes
Gaia DR3 5262754273969986688 1.371 1.288 0.040 0.053 0.106
Gaia DR3 5268812686118172928 10.585 1.149 29.892 0.099 0.000 0.000 -0.180 0.107
Gaia DR3 5268812853620242048 1.112 30.072 0.041 0.000 0.000
Gaia DR3 3605478986040028544 14.727 1.776 -31.166 0.029 0.049 0.163 -1.201 0.467
Gaia DR3 3605479020399765888 0.210 -29.965 0.466 0.000 2.162
Gaia DR3 425040000962559616 0.079 1.054 9.130 0.042 0.000 0.301 0.244 0.059
Gaia DR3 425040000962497792 0.598 8.886 0.041 0.121 0.184
Gaia DR3 4676751272565519872 1.718 0.443 -19.618 0.118 0.000 0.000 0.999 0.578
Gaia DR3 4676751341284996480 0.390 -20.617 0.566 0.000 0.522
Gaia DR3 4669281259285732480 21.759 1.119 35.245 0.031 0.000 0.000 -0.113 0.045
Gaia DR3 4669281121846284160 1.018 35.358 0.033 0.000 0.000
Gaia DR3 3265508650701713664 2.612 1.030 29.067 0.050 0.000 0.107 -0.211 0.058
Gaia DR3 3265508685061451648 1.018 29.278 0.029 0.000 0.019
Gaia DR3 3266980170921153920 6.036 0.879 88.594 0.234 0.000 0.000 -0.287 0.240
Gaia DR3 3266980243936341248 0.417 88.881 0.055 0.000 0.054
Gaia DR3 4757024829820200704 5.504 0.974 -6.794 0.056 0.000 0.000 0.633 0.084
Gaia DR3 4757026135488693760 0.676 -7.427 0.062 0.000 0.000
Gaia DR3 3609320267350909696 10.223 0.787 -56.802 0.045 0.000 0.002 0.353 0.055
Gaia DR3 3609320503573212416 0.743 -57.155 0.032 0.000 0.086
Gaia DR3 1652245748082610304 1.342 0.961 54.463 0.032 0.000 0.041 0.348 0.059
Gaia DR3 1652245748082125440 0.557 54.115 0.049 0.000 0.000
Gaia DR3 2104999343337767424 19.187 0.953 -18.034 0.025 0.000 0.000 0.164 0.048
Gaia DR3 2104998900960077056 0.738 -18.198 0.041 0.000 0.134
Gaia DR3 3922083599776523392 1.955 0.772 -11.732 0.038 0.000 0.035 -0.581 0.215
Gaia DR3 3922083599776523520 0.429 -11.151 0.212 0.000 0.632
Gaia DR3 3928893253244519168 6.337 1.197 -2.068 0.026 0.076 0.092 -0.109 0.182
Gaia DR3 3928893253244519552 0.408 -1.959 0.180 0.000 0.578
Gaia DR3 3929154078017979264 2.895 1.166 -36.255 0.036 0.000 0.093 -0.365 0.100
Gaia DR3 3929154078017979008 0.642 -35.890 0.093 0.000 0.313
Gaia DR3 3936935283153115008 38.482 1.309 2.580 0.043 0.000 0.131 -0.120 0.078
Gaia DR3 3936947278997113216 0.580 2.700 0.065 0.000 0.365
Gaia DR3 3958567766407880576 0.909 0.573 -14.002 0.014 0.110 0.048 0.682 0.024 Yes
Gaia DR3 3958567766407880704 0.444 -14.684 0.019 0.110 0.044
Gaia DR3 690460424271103360 22.780 0.920 -32.999 0.031 0.000 0.054 -0.126 0.062
Gaia DR3 690272545221668096 0.606 -32.873 0.054 13.924 0.925
Gaia DR3 694284491350980864 3.288 0.636 -2.491 0.031 0.000 0.062 -0.276 0.100
Gaia DR3 694284628789933568 0.318 -2.215 0.095 0.000 0.092
Gaia DR3 2164378370045365376 1.115 0.413 6.758 0.053 0.000 0.147 0.098 0.070
Gaia DR3 2164331434639561088 0.377 6.660 0.046 0.028 0.200
Gaia DR3 678053947384823936 1.619 0.876 -2.451 0.019 0.014 0.019 -1.144 0.144
Gaia DR3 678053878665346816 0.421 -1.307 0.143 0.000 1.342
Gaia DR3 678095763181573632 1.901 0.876 -8.700 0.019 0.018 0.203 -0.298 0.029 Yes
Gaia DR3 678095763182104704 0.749 -8.402 0.022 0.051 0.027
Gaia DR3 676003255114571008 1.101 0.730 46.643 0.027 0.000 0.000 -0.324 0.038
Gaia DR3 676003255114571264 0.594 46.967 0.027 0.006 0.085
Gaia DR3 712074211531234816 23.757 0.397 23.960 0.061 0.000 0.023 -0.278 0.101
Gaia DR3 712024802229009920 0.326 24.238 0.081 0.000 0.059
Gaia DR3 705170274943504000 3.125 1.424 7.595 0.028 0.000 0.073 0.040 0.036
Gaia DR3 705170309303241984 1.103 7.555 0.023 0.000 0.016
Gaia DR3 705997794881180544 2.148 1.024 10.026 0.030 0.000 0.042 0.905 0.057
Gaia DR3 705997794881180672 0.892 9.121 0.049 0.000 0.016
Gaia DR3 2462426800883134336 0.996 1.233 -8.747 0.026 0.000 0.000 0.072 0.035
Gaia DR3 2462426800883156480 0.599 -8.819 0.024 0.000 0.000
Gaia DR3 2475292457022466048 4.118 0.946 18.337 0.111 0.000 0.319 0.389 0.402
Gaia DR3 2475292530037116288 0.652 17.948 0.386 0.000 0.335
Gaia DR3 923398868922049792 1.639 0.783 16.981 0.038 0.000 0.026 -0.402 0.066
Gaia DR3 923398864626108416 0.644 17.383 0.054 0.000 0.000
Gaia DR3 926109886638467840 0.697 0.558 93.745 0.047 0.000 0.135 0.436 0.083
Gaia DR3 926109886638467968 0.483 93.309 0.069 0.000 0.173
Gaia DR3 3686261307923490688 4.502 0.469 41.577 0.038 0.000 0.060 -0.177 0.084
Gaia DR3 3686259727375524864 0.301 41.754 0.075 0.000 0.045
Gaia DR3 1434064639260540288 1.857 1.299 -2.610 0.070 0.000 0.000 0.187 0.076
Gaia DR3 1434064669325509760 0.862 -2.797 0.030 0.000 0.000
Gaia DR3 1427740214018187392 2.646 0.618 -31.220 0.058 0.000 0.125 -0.691 0.172
Gaia DR3 1427740214018187648 0.372 -30.529 0.162 0.000 0.230
Gaia DR3 4763619318293079168 6.382 0.937 31.561 0.021 0.000 0.000 -0.180 0.075
Gaia DR3 4763618910272182400 0.348 31.741 0.072 0.000 0.000
Gaia DR3 2086713404118555648 0.442 0.974 -29.198 0.032 0.000 0.000 0.382 0.043
Gaia DR3 2086713404118555520 0.863 -29.580 0.029 0.000 0.000
Gaia DR3 5496774878580968832 21.476 0.563 23.446 0.047 0.000 0.000 -0.046 0.097
Gaia DR3 5496779998182670976 0.405 23.492 0.085 0.000 0.000
Gaia DR3 162758236656524416 13.291 0.696 17.969 0.047 0.309 1.291 -0.456 0.078 Yes
Gaia DR3 162757545164429696 0.683 18.425 0.062 1.186 1.353
Gaia DR3 1444305829863259648 6.749 0.791 -5.889 0.019 0.043 0.106 0.349 0.038
Gaia DR3 1444305937237949184 0.612 -6.238 0.033 0.000 0.000
Gaia DR3 1448490437975655552 35.425 1.323 -36.479 0.028 0.031 0.163 0.194 0.046
Gaia DR3 1448490369254832896 1.099 -36.673 0.036 0.000 0.070
Gaia DR3 1456403004684961536 0.600 0.398 -16.261 0.065 0.000 0.012 0.436 0.089
Gaia DR3 1456403004684961408 0.394 -16.697 0.061 0.000 0.062
Gaia DR3 2129540099271049216 0.786 0.919 -6.230 0.038 0.000 0.070 1.331 0.050
Gaia DR3 2129540103571086592 0.864 -7.561 0.032 0.000 0.362
Gaia DR3 2130903223113004032 4.305 1.185 -5.509 0.035 0.000 0.000 0.210 0.056
Gaia DR3 2130903047014230528 0.957 -5.719 0.044 0.000 0.031
Gaia DR3 832823020875083776 2.767 0.601 -2.311 0.029 0.095 0.102 0.320 0.101
Gaia DR3 832823025170612736 0.361 -2.631 0.097 0.000 0.272
Gaia DR3 839281109796627968 17.651 0.766 3.133 0.020 0.043 0.035 0.043 0.030 Yes
Gaia DR3 839281796991751168 0.737 3.090 0.022 0.065 0.064
Gaia DR3 839567326416583552 2.362 0.733 -20.752 0.041 0.000 0.073 0.347 0.070
Gaia DR3 839567326416583808 0.336 -21.099 0.057 0.085 0.118
Gaia DR3 6236104295448198272 3.850 0.561 -2.831 0.060 0.000 0.000 -0.679 0.195
Gaia DR3 6236103917491077120 0.418 -2.152 0.186 0.000 0.842
Gaia DR3 6041787193156436992 0.739 0.390 -38.218 0.086 0.000 0.000 0.157 0.585
Gaia DR3 6041787193156437248 0.305 -38.375 0.579 0.000 0.000
Gaia DR3 6044225119671978752 11.803 0.596 -2.069 0.061 0.000 0.000 -0.104 0.085
Gaia DR3 6044230960827500032 0.526 -1.965 0.059 0.000 0.000
Gaia DR3 972821779152541824 19.865 0.601 64.726 0.031 0.000 0.050 -0.160 0.048
Gaia DR3 972815560039899264 0.546 64.886 0.037 0.000 0.037
Gaia DR3 988478859091454976 4.261 0.698 41.484 0.044 0.000 0.059 -0.661 0.141
Gaia DR3 988478957873964544 0.417 42.145 0.134 0.000 0.118
Gaia DR3 134816244579883008 3.828 0.752 -3.744 0.041 0.000 0.074 0.050 0.062
Gaia DR3 134816244579883776 0.674 -3.794 0.046 0.000 0.072
Gaia DR3 1159502498311670144 3.022 0.821 23.549 0.022 0.000 0.027 0.126 0.042
Gaia DR3 1159502154714284544 0.489 23.423 0.036 0.000 0.032
Gaia DR3 4665709113509264768 5.923 0.783 22.638 0.018 0.058 0.000 0.173 0.046
Gaia DR3 4665710629633988736 0.560 22.465 0.042 0.000 0.000
Gaia DR3 1172915990414659328 24.684 0.885 -17.931 0.023 0.000 0.044 0.043 0.033
Gaia DR3 1172920487244742912 0.852 -17.974 0.023 0.000 0.051
Gaia DR3 935566717429616000 8.615 0.980 38.446 0.016 0.128 0.049 -0.617 0.054 Yes
Gaia DR3 935566511271187072 0.452 39.063 0.052 0.935 52.862
Gaia DR3 933708229245319552 2.249 0.321 4.559 0.074 0.000 0.037 0.106 0.113
Gaia DR3 933708164821511424 0.270 4.453 0.086 0.000 0.012
[…]
Table 8: Summary of wide binaries with multiple radial velocities of both components from multi-epoch APOGEE observationsaafootnotetext: Reported error.
[…]
Star A RVaaReported radial velocity. Errbbfootnotemark: ScattccReported value of VSCATTER. Star B RVaaReported radial velocity. Errbbfootnotemark: ScattccReported value of VSCATTER.
[km s-1] [km s-1]
Gaia DR3 2642922251741817216 23.611124 0.04403066 0.04110121 2642922286101533952 23.745401 0.04636222 0.06235123
23.493046 0.03530691 0.04366402 23.661783 0.03730566 0.02735676
Gaia DR3 2501323948860951296 -26.34998 0.03360082 0.00961134 2501325525113734912 -26.389238 0.04771343 0.11454166
-26.28739 0.04346129 0.06616001 -26.260921 0.03940159 0.04780705
Gaia DR3 217107577453306112 22.248499 0.02819021 0.07399079 217107607515576704 21.878296 0.03318193 0.05599995
22.275375 0.02912341 0.144448 21.788877 0.02936134 0.
Gaia DR3 3961746072270936320 -45.91673 0.03888413 0. 3961745217573223296 -45.767345 0.05168254 0.
-45.911835 0.05633869 0. -45.684467 0.05476467 0.03869113
-45.86499 0.04045845 0.06876723 -45.90399 0.04622763 0.01500438
-45.8177 0.03833721 0.05585103 -45.69478 0.08768474 0.12935062
-45.823917 0.06022181 0.15365449 -45.790802 0.06783935 0.05524765
-45.71305 0.05185273 0.06499686 -45.664474 0.04946134 0.
-45.744324 0.0405033 0. -45.800285 0.05546138 0.01956754
-45.77147 0.04337924 0.01045496 -45.666336 0.05314275 0.02631607
-45.743183 0.04535097 0.06701745 -45.732433 0.03973199 0.
-45.965115 0.03278913 0. -45.759113 0.05460643 0.06531424
-45.867214 0.04078658 0.00449147
Gaia DR3 2581883413481515008 -14.654112 0.02970796 0. 2581859911420471808 -14.30785 0.04154328 0.
-14.7650795 0.03625005 0. -14.515479 0.05171923 0.
Gaia DR3 785716030026546688 -15.540678 0.04524035 0.03031975 785716682861545344 -15.415195 0.06123991 0.1664266
-15.592761 0.04024156 0.08631156 -15.533735 0.05448572 0.01872353
Gaia DR3 1461182444292609536 -13.681904 0.03039815 0.01888097 1461182684811100288 -13.209225 0.02909379 0.04755092
-13.66857 0.04013753 0.07308564 -13.213476 0.02948592 0.01699376
-13.130007 0.03311201 0.02252192
Gaia DR3 1469606455707961088 -12.468412 0.03486331 0.0247436 1469606283909297152 -12.182982 0.02783493 0.0102103
-12.4709635 0.03095368 0.05807732 -12.376394 0.03245912 0.04527421
-12.447786 0.03108544 0.00479869 -12.307321 0.03310763 0.00351568
Gaia DR3 1535360373488627200 -11.203867 0.02588832 0.01143761 1535360343424477440 -10.9632845 0.03258488 0.01365597
-11.3142395 0.02703807 0.0534603 -11.014411 0.03550287 0.03829292
Gaia DR3 5262754514488149632 1.2070447 0.03156854 0. 5262754273969986688 1.3154067 0.04626257 0.
1.3321234 0.05412297 0.08124872 1.2100439 0.07827925 0.10565944
Gaia DR3 3958567766407880576 -14.007257 0.03145121 0.0253697 3958567766407880704 -14.654441 0.05386632 0.
-13.967669 0.03088553 0.0154797 -14.469742 0.05273892 0.0115321
-13.915713 0.0328168 0.034364 -14.670951 0.05127405 0.02754166
-13.89397 0.03500261 0.04774026 -14.732671 0.05338896 0.03030096
-14.2029 0.03196205 0. -14.631961 0.05492502 0.03866151
-14.683698 0.0498655 0.
-14.773074 0.05134532 0.00627914
-14.874403 0.05639074 0.04404995
Gaia DR3 678095763181573632 -8.719339 0.02869016 0.20294154 678095763182104704 -8.34763 0.03163528 0.01059777
-8.683017 0.02654715 0. -8.449814 0.02956442 0.02662994
Gaia DR3 162758236656524416 18.270649 0.06570818 1.2907226 162757545164429696 19.58776 0.08644723 1.3527985
17.651651 0.06742151 0.29228374 17.215908 0.08812028 1.2935578
Gaia DR3 839281109796627968 3.1859643 0.03458128 0.0264379 839281796991751168 3.1243968 0.03509628 0.0466863
3.0797095 0.03401986 0.00547901 3.007658 0.03551762 0.06423731
3.1357276 0.0380827 0.03508588 3.1594172 0.04301288 0.01149944
Gaia DR3 935566717429616000 38.319893 0.02177464 0.02721165 935566511271187072 38.15186 0.07279947 52.861546
38.576614 0.0222188 0.04888935 40.022377 0.07470552 40.95351
Appendix E Description of Speckle OAN-SPM observations and data reduction
E.1 Speckle observations at 2.1m telescope of the Observatorio Astronómico Nacional at Sierra de San Pedro Mártir (SPM), México.
[…]
Table 9:
Run Cubes Cubes Cubes
$256\times 256$ $512\times 512$ $1024\times 1024$
Feb-24 35 8
Sep-24 131 156
Nov-24 70 139 1
Mar-25 149
Jun-25 267 63
Total 652 358 9
E.2 Data Processing
[…]
The first step of the data processing is the dark field correction of detected images $i^{\prime}_{n}(\vec{x})$:
$i_{n}(\vec{x})=i^{\prime}_{n}(\vec{x})-Dark(\vec{x}),$ (E1)
where $i_{n}(\vec{x})$ is the intensity at point $(\vec{x})$ in the corrected image, $Dark(\vec{x})$ is the average dark image captured with closed shutter. Then for each specklegram we calculate the following criteria:
$Sh1_{n}=\left[\sum i_{n}(\vec{x})\right]^{2}/\sum i_{n}^{2}(\vec{x}),$ (E2)
where the summations extend over the complete n-th specklegram (McCarthy and Cobb, 1986). Speckle images are then ranked by their Sh1, where the ”best” frames are those with the lowest Sh1. Two images in Figure 22 show the best (a) and worst (b) speckle images in a series of 2000 specklegrams.
Figure 22: Data processing. Left to right: Best frame, worst speckle images in a series of 2000 specklegrams
E.3 High-Resolution
The intensity distribution $i_{n}(\vec{x})$ of the $n$’th short exposure image (specklegram) can be described by:
$i_{n}(\vec{x})=o(\vec{x})\otimes PSF_{n}(\vec{x}),$ (E3)
where $\vec{x}$ is a 2D spatial coordinate, $o(\vec{x})$ is the object intensity distribution, $PSF_{n}(\vec{x})$ is the Point Spread Function for the $n$’th specklegram and $\otimes$ denotes convolution.
Then, applying the Fourier transform to equation E3
$I_{n}(\vec{\omega})=O(\vec{\omega})\times OTF_{n}(\vec{\omega}),$ (E4)
where $I_{n}(\vec{\omega})$ and $O(\vec{\omega})$ are the Fourier transform of the image and object intensity distributions respectively, and $OTF_{n}(\vec{\omega})$ is the Optical Transfer Function (OTF), which is the equivalent of the $PSF$ in Fourier space. The $OTF_{n}(\vec{\omega})$ is the same for all objects within the isoplanatic region. Equation E4 is the base for all high-resolution techniques. We applied various image processing techniques to reconstruct the power spectrum and obtain a high-resolution image of the star. This involves analyzing the speckle patterns in each short-exposure frame and combining them.
E.4 Power Spectrum
The next step is to calculate the averaged power spectrum (PS) for each star:
$PS(\vec{\omega})=\left\langle\left|I_{n}(\vec{\omega})\right|^{2}\right\rangle,$ (E5)
[…]
In case of low light images, the averaged power spectrum can be expressed as (Kerp et al., 1992):
$PS(\vec{\omega})=P(\vec{\omega})\cdot\left|G(\vec{\omega})\right|^{2}+q\left|G(\vec{\omega})\right|^{2},$ (E6)
where $P(\vec{\omega})$ is the unshifted estimation of the power spectrum, $q$ is some constant, $\left|G(\vec{\omega})\right|^{2}$ is the power spectrum of the photon event shape function also known as photon bias. The photon bias $\left|G(\vec{\omega})\right|^{2}$ can be determined as the normalized power spectrum of the night sky. The $\left|G(\vec{\omega})\right|^{2}$ is constant in Y direction for this camera. Thus, it can be determined directly from $PS(\vec{\omega})$ (Figure 23, (a)) by analysis of it’s part beyond the cut-off frequency of telescope. The unshifted power spectrum of specklegrams $P({\bf f})$ shown in Figure 23 (b). Therefore, it can be presented as:
$P(\vec{\omega})=\left|O(\vec{\omega})\right|^{2}\left\langle\left|S_{n}(\vec{\omega})\right|^{2}\right\rangle,$ (E7)
[…]
where $\left|O(\vec{\omega})\right|^{2}$ power spectrum of the object, $\left\langle\left|S_{n}(\vec{\omega})\right|^{2}\right\rangle$ is the speckle interferometric transfer function. The speckle interferometric transfer function can be obtained by observing of a reference star or one can construct universal synthetic speckle interferometric transfer function (Tokovinin et al., 2010). Since we are concerned only with the presence or absence of a close stellar component, this information can be obtained directly from $P(\vec{\omega})$ without requiring the speckle interferometric transfer function. In particular, we calculated the high resolution autocorrelation function in polar coordinates $ACF_{p}$:
$ACF_{p}(\rho,\theta)=const\int_{0}^{\infty}\int_{0}^{2\pi}\cos(2\pi r\rho\cos(\theta-\phi))P({r,\phi})W(r,\phi)rd{r}d{\phi},$ (E8)
where $W(r,\phi)$ is the window which excludes part of $P(r,\phi)$ beyond the cut-off frequency of telescope $f_{T}$ and for frequencies lower atmospheric cutoff $f_{A}$. Also taking in to account central symmetry of $P(r,\phi)$ equation E8 can be rewritten as:
$ACF_{p}(\rho,\theta)=const\int_{f_{A}}^{f_{T}}\int_{0}^{\pi}\cos(2\pi r\rho\cos(\theta-\phi))P({r,\phi})rd{r}d{\phi}.$ (E9)
A star has a component if the high-resolution $ACF_{p}$ has a pronounced maximum. $ACF_{p}$ allows finding the component even when the power spectrum is distorted by vibrations and strong aberrations of the telescope (Orlov, 2021). One example of $ACF_{p}$ is shown in figure 23 (c).
E.5 Speckle Holography
Figure 24: The typical wide binary system, composed of Gaia DR3 1172915990414659328 (G = 9.2) and Gaia DR3 1172920487244742912 (G = 9.4). (a) High-resolution autocorrelation of Gaia DR3 1172915990414659328. (b) High-resolution speckle holography image of Gaia DR3 1172915990414659328. (c) High-resolution autocorrelation of Gaia DR3 1172920487244742912. (d) High-resolution speckle holography image of Gaia DR3 1172920487244742912.
For most wide binaries we found no hidden stellar components (Figure 24). However, during data processing we identified several interesting systems — for example, Gaia DR3 4493544185516358656, which turned out to be a triple star system (Fig. 25). To investigate such cases in more detail, we applied speckle holography to the entire dataset. The method can be summarized as follows:
If one has a point source $O^{r}(\vec{x})=\delta(\vec{x}=\vec{x^{r}})$ inside this area, the equation E4 for this star becomes:
$I_{n}^{r}(\vec{\omega})=OTF_{n}(\vec{\omega}).$ (E10)
Substitution of equation E10 into equation E4 gives:
$I_{n}(\vec{\omega})=O(\vec{\omega})\times I_{n}^{r}(\vec{\omega}).$ (E11)
The Fourier transform of the object intensity distributions can be estimated by :
$O(\vec{\omega})=\frac{\langle I_{n}(\vec{\omega})I_{n}^{r*}(\vec{\omega})\rangle}{\langle|I_{n}^{r}(\vec{\omega})|^{2}\rangle},$ (E12)
where $\langle$$\rangle$ means averaging over all images and is the complex conjugation. Finally, the object intensity distributions can be obtained by the inverse Fourier transform.
Figure 25: The triple system ABC, composed of Gaia DR3 4493544185516358656 (G = 12.86), Gaia DR3 4493544082437142400 (G = 13.26), and Gaia DR3 4493544082437142784 (G = 14.13). (a) Power spectrum of subsystem BC. (b) High-resolution autocorrelation of subsystem BC. (c) High-resolution speckle holography image. It was derived from Speckle Interferometry (Labeyrie, 1970), being the main difference that speckle holography works to remove atmospheric aberrations from each short exposure image and the phase information is kept. The separation between components B and C is 0.1″(8 pc).
E.6 detection limit
The detection limit describes how faint a companion star could be and still be identifiable in the data. It is defined by the minimum signal-to-noise ratio needed to separate a real signal from random noise. In practice, astronomers usually adopt a threshold of $5\sigma$, which makes the chance of mistaking noise for a real star well below one percent. Sometimes a lower threshold of $3\sigma$ is used, but this is risky. At this level, it is rare, but possible, to detect spurious companions caused by noise (Tokovinin et al., 2010).
At large separations from the primary star (more than about three times the seeing disk) the noise reaches a steady level, and the detection limit becomes roughly constant. At smaller separations, however, it strongly depends on angular distance from the primary star. This behavior is summarized with a contrast curve (Figure 26). The contrast curve plots the faintest detectable companion (in terms of brightness difference, $\Delta m$) as a function of separation from the primary star. In other words, it tells you: If there had been a companion this bright at this distance, we would have seen it.
Figure 26: The $5\sigma$ Contrast curves. (a) The curve is for a star with a magnitude of 12, while (b) is the curve for a star with a magnitude of 10.
Since no companions were detected in the vast majority of cases, we present for each star only two representative points of the contrast curve, corresponding $\Delta m=3$ and $\Delta m=4$.
In total, we obtained 1019 measurements for 391 selected wide binaries. Many targets were observed two or more times, allowing us to retain only the highest-quality measurements for each system. The table 10 lists the results for 782 stars.
[…]
Table 10: Summary of Speckle Observations
[…]
Comment Identifier Julian Year Dm 4 Dm 3
Gaia DR3 1003223584897948160 2024.8650 0.28 0.20
Gaia DR3 1003223614961194752 2024.8650 0.38 0.27
Gaia DR3 1020666649676726784 2025.2123 0.26 0.20
Gaia DR3 1020667714828594304 2025.2123 0.37 0.26
Gaia DR3 1023887978228642176 2025.2176 0.31 0.22
Gaia DR3 1023888012588380032 2024.8680 0.25 0.18
Gaia DR3 1028432805245868544 2024.8680 0.24 0.17
Gaia DR3 1028432805250731264 2024.8680 0.26 0.19
Gaia DR3 104998928046427264 2024.6869 0.32 0.22
Gaia DR3 105004842216905344 2024.6869 0.32 0.22
Gaia DR3 1159502154714284544 2025.4422 1.00 0.61
Gaia DR3 1159502498311670144 2025.4422 0.19 0.14
Gaia DR3 1172915990414659328 2025.4393 0.18 0.12
Gaia DR3 1172920487244742912 2025.4393 0.19 0.13
Gaia DR3 1174143182830505856 2025.4448 0.17 0.12
Gaia DR3 1174143182830505984 2025.4448 0.11 0.10
Gaia DR3 1178121422057906688 2025.4393 0.19 0.13
Gaia DR3 1178121422057907200 2025.4393 0.27 0.19
Gaia DR3 1182257922240686720 2025.4339 1.33 1.01
Gaia DR3 1182257922240687232 2025.4339 1.27 0.90
Gaia DR3 1214182826468995584 2025.4367 0.45 0.30
Gaia DR3 1214182826468995840 2025.4367 0.55 0.34
Gaia DR3 121690893241571968 2024.8618 0.92 0.49
Gaia DR3 121691266902954240 2024.8618 0.64 0.37
Gaia DR3 1224242120913963776 2025.4340 0.40 0.27
Gaia DR3 1224242120913963904 2025.4340 0.39 0.25
Gaia DR3 1227416514060783232 2025.4448 0.45 0.31
Gaia DR3 1227417304334766208 2025.4366 0.27 0.21
Gaia DR3 1247610934890765184 2025.4420 0.53 0.36
Gaia DR3 1247610934890765312 2025.4420 0.39 0.26
Gaia DR3 1258410612976538368 2025.4393 0.10 0.09
Gaia DR3 1258410750415492864 2025.4393 0.13 0.11
Gaia DR3 1271824002003198464 2025.4367 0.98 0.59
Gaia DR3 1271824002003198976 2025.4367 0.98 0.58
Gaia DR3 1274562370068531712 2025.4367 0.19 0.13
Gaia DR3 1274568245587206144 2025.4367 0.13 0.11
Gaia DR3 1282815063829295360 2025.4393 0.10 0.09
Gaia DR3 1282817022334383104 2025.4394 0.11 0.10
Gaia DR3 1282817022334383232 2025.2182 0.18 0.13
Gaia DR3 1297419700686304896 2025.4341 0.84 0.45
Gaia DR3 1297419808063095936 2025.4342 0.71 0.40
Gaia DR3 1299746168507031424 2025.4450 0.17 0.12
Gaia DR3 1299746168508720384 2025.4450 0.17 0.12
Gaia DR3 1302449489644570496 2025.4395 0.00 0.00
Gaia DR3 1302543326089734144 2025.4395 1.07 0.65
Gaia DR3 1305035575352515840 2025.4396 0.50 0.32
Gaia DR3 1305036228187622912 2025.4396 0.00 0.00
Gaia DR3 1306220264772422016 2025.4450 0.29 0.21
Gaia DR3 1306220642729543936 2025.4450 0.28 0.20
Gaia DR3 1308215328620272256 2025.4450 0.28 0.20
Gaia DR3 1308215328620272384 2025.4450 0.26 0.19
Gaia DR3 1320009686772023296 2025.4367 0.34 0.23
Gaia DR3 1320009721131760640 2025.4367 0.27 0.19
Gaia DR3 1320379809873587968 2025.4396 0.38 0.27
Gaia DR3 1321677920789034752 2025.4341 0.45 0.30
Gaia DR3 1321677920789214080 2025.4341 0.69 0.42
Gaia DR3 1327463825851820800 2025.4450 0.28 0.19
Gaia DR3 1327463825851820928 2025.4450 0.26 0.18
Gaia DR3 1331462302965590144 2025.4396 0.00 0.00
Gaia DR3 1331462302965590400 2025.4396 0.19 0.13
Gaia DR3 1332438183958298496 2025.4423 0.40 0.28
Gaia DR3 1332438188254551808 2025.4423 0.49 0.33
Gaia DR3 1332473922382506624 2025.4341 1.32 1.01
Gaia DR3 1332473922382951424 2025.4341 1.02 0.57
Gaia DR3 1346694112423491712 2025.4370 0.91 0.53
Gaia DR3 1346694146783637504 2025.4370 0.30 0.22
Gaia DR3 1348285896022947584 2025.4370 0.13 0.11
Gaia DR3 1348286651937191040 2025.4370 0.23 0.16
Gaia DR3 1352381027080312832 2025.4423 0.19 0.14
Gaia DR3 1352427996842670464 2025.4423 0.58 0.38
Gaia DR3 1356763852227145344 2025.4424 1.03 0.73
Gaia DR3 1356763886586884864 2025.4424 0.92 0.59
Gaia DR3 1357040345041982464 2025.4424 0.90 0.58
Gaia DR3 1357040379401720448 2025.4424 0.44 0.31
Gaia DR3 1371903199628510464 2025.4394 0.15 0.12
Gaia DR3 1371903199628531200 2025.4394 0.21 0.13
Gaia DR3 1379737421839743104 2025.4341 0.30 0.21
Gaia DR3 1379737426135359360 2025.4341 0.29 0.20
Gaia DR3 1381408782592655616 2025.4424 0.45 0.31
Gaia DR3 1381455722290898176 2025.4424 1.13 0.80
Gaia DR3 1391176092275606528 2025.4394 0.00 0.00
Gaia DR3 1391176092275606656 2025.4394 0.29 0.20
Gaia DR3 1396388773823564416 2025.4340 0.34 0.24
Gaia DR3 1396388773823907584 2025.4340 0.31 0.22
Gaia DR3 1410982832177435008 2025.4341 1.13 0.72
Gaia DR3 1410982862240353664 2025.4341 0.40 0.27
Gaia DR3 1419999819661235840 2025.4344 0.19 0.13
Gaia DR3 1420000579871414528 2025.4344 0.17 0.12
Gaia DR3 1434064639260540288 2025.4397 0.37 0.27
Gaia DR3 1434064669325509760 2025.4397 1.15 0.78
Gaia DR3 1444305829863259648 2025.4420 0.54 0.36
Gaia DR3 1444305937237949184 2025.4420 0.92 0.60
Gaia DR3 1467776078085413888 2025.4392 0.31 0.22
Gaia DR3 1467776078085414400 2025.4393 0.32 0.23
Gaia DR3 1478999686625786368 2025.4366 0.28 0.21
Gaia DR3 1478999755345263744 2025.4366 0.39 0.28
Gaia DR3 1479430076707007488 2025.4394 0.25 0.17
Gaia DR3 1482432155767129728 2025.4394 0.13 0.11
Gaia DR3 1508470171660833536 2025.4366 0.53 0.33
Gaia DR3 1508470206020209664 2025.4366 0.53 0.34
Gaia DR3 1509241238549271808 2025.4365 0.34 0.25
Gaia DR3 1509241238549292800 2025.4365 0.62 0.39
Gaia DR3 1509783778818258176 2025.4365 0.23 0.18
Gaia DR3 1509783778818258304 2025.4365 0.33 0.25
Gaia DR3 1527328376624427776 2025.4393 0.33 0.24
Gaia DR3 1527328376624971520 2025.4393 0.35 0.26
Gaia DR3 153741691551129216 2024.1277 0.40 0.28
Gaia DR3 153741760270606464 2024.1277 0.40 0.28
Gaia DR3 1543083553620064768 2024.1286 0.00 0.00
Gaia DR3 1543083691059006592 2024.1286 0.38 0.28
Gaia DR3 1549520949383005568 2025.4365 0.28 0.21
Gaia DR3 1549521125478552064 2025.4365 0.18 0.13
Gaia DR3 1585800782946765312 2025.4448 0.46 0.32
Gaia DR3 1585800851666242688 2025.4448 0.25 0.19
Gaia DR3 1586977737129182848 2025.4448 0.20 0.14
Gaia DR3 1586977844504488576 2025.4448 0.13 0.11
Gaia DR3 1590641417247939328 2025.4394 0.00 0.00
Gaia DR3 1590641417247939584 2025.4394 0.82 0.52
Gaia DR3 1592707502675396992 2025.4339 0.38 0.26
Gaia DR3 1592707502675731456 2025.4339 0.37 0.26
Gaia DR3 1595968929041676544 2025.4340 0.48 0.32
Gaia DR3 1595968929041676672 2025.4340 0.44 0.30
Gaia DR3 1605833880509338752 2025.4448 0.29 0.21
Gaia DR3 1605833884804824704 2025.4448 0.40 0.28
Gaia DR3 1610186095424889088 2025.4448 0.21 0.16
Gaia DR3 1610191970940150272 2025.4448 0.27 0.20
Gaia DR3 1610236260642857344 2025.4448 0.32 0.24
Gaia DR3 1610236260642871552 2025.4448 0.32 0.24
Gaia DR3 162781051522771072 2024.8621 0.38 0.26
Gaia DR3 162781051522771200 2024.8621 0.41 0.29
Gaia DR3 167549908330274176 2024.8618 0.64 0.36
Gaia DR3 167549942690010368 2024.8618 0.94 0.54
Gaia DR3 1730273101846974080 2024.6838 0.54 0.34
Gaia DR3 1730273106143055488 2024.6838 0.44 0.31
Gaia DR3 1738269854212128512 2024.6810 0.35 0.24
Gaia DR3 1738269854213155712 2024.6810 0.37 0.26
Gaia DR3 1749987585851186816 2024.6783 0.57 0.34
Gaia DR3 1749987585854235776 2024.6783 1.84 1.33
Gaia DR3 1760471948915107200 2025.4455 0.18 0.13
Gaia DR3 1760477618271932672 2025.4455 0.26 0.18
Gaia DR3 1762461309047562240 2024.6782 0.45 0.30
Gaia DR3 1762461309047562368 2025.4400 0.35 0.24
Gaia DR3 1762461893163118464 2025.4455 0.33 0.23
Gaia DR3 1762461893163118592 2024.6782 0.67 0.37
Gaia DR3 1778929480673414784 2024.6838 0.23 0.16
Gaia DR3 1778930240883745536 2024.6839 0.00 0.00
Gaia DR3 1781495706454162944 2024.6866 0.27 0.18
Gaia DR3 1781495710748864256 2024.6866 0.23 0.16
Gaia DR3 1789024749766697088 2024.6810 1.16 0.85
Gaia DR3 1789024857138962048 2024.6810 0.20 0.14
The third component is DR3 181318611409243136, Gaia DR3 181312735893067008 2024.8650 0.20 0.15
no parallax data available.
The third component is DR3 181318611409243136, Gaia DR3 181318611408327552 2024.8650 0.17 0.12
no parallax data available.
The third component, DR3 1815165879238575232, Gaia DR3 1815165535636339072 2025.4400 0.16 0.12
is too far away, DM=4.3
Gaia DR3 1861752775327557504 2024.6781 0.37 0.25
Gaia DR3 1861752775327558656 2024.6781 0.45 0.30
Gaia DR3 1868257722327168768 2024.6811 0.72 0.38
Gaia DR3 1868257726634854784 2024.6811 0.67 0.36
Gaia DR3 1868985324150840192 2024.6864 0.22 0.14
Gaia DR3 1868985328454953856 2024.6864 0.24 0.15
Gaia DR3 1871558941576158464 2024.6810 0.38 0.25
Gaia DR3 1871559697490418816 2025.4455 0.46 0.30
Gaia DR3 1894003238061739648 2024.6864 1.01 0.61
Gaia DR3 1894003238061739776 2024.6864 1.19 0.74
Gaia DR3 1895940302674258304 2024.6865 1.26 0.82
Gaia DR3 1895940302674258688 2024.6865 0.28 0.19
Gaia DR3 1899383114098895616 2024.6864 0.28 0.19
Gaia DR3 1899383114098895872 2024.6864 0.33 0.22
Gaia DR3 1902676117063909888 2024.6865 0.43 0.29
Gaia DR3 1902679033343158528 2024.6865 0.43 0.28
Gaia DR3 1904247284819944448 2024.6865 0.35 0.25
Gaia DR3 1904247284820463104 2024.6865 0.25 0.17
Gaia DR3 1909810023381344768 2024.6866 0.40 0.27
Gaia DR3 1909810126460559616 2024.6866 0.43 0.29
Gaia DR3 1910071432270904832 2024.6866 0.37 0.25
Gaia DR3 1910071535350120192 2024.6866 0.42 0.28
Gaia DR3 1921150828692050304 2024.6866 0.51 0.34
Gaia DR3 1921150832986231680 2024.6866 1.19 0.79
Gaia DR3 1923116072583279232 2024.6865 0.24 0.16
Gaia DR3 1923116072583279488 2024.6759 0.60 0.39
Gaia DR3 1931673464205703936 2024.6865 0.55 0.34
Gaia DR3 1931673532926948864 2024.6865 0.96 0.52
Gaia DR3 1932297578791924992 2024.6865 0.22 0.14
Gaia DR3 1932297578791925760 2024.6865 0.22 0.14
Gaia DR3 1938247517245907456 2024.6866 0.17 0.12
Gaia DR3 1938247654684985344 2024.6866 0.15 0.12
Gaia DR3 1942018494236897792 2024.6866 0.62 0.36
Gaia DR3 1942018498533031808 2024.6866 0.56 0.34
Gaia DR3 1942384773344557056 2024.6865 0.20 0.13
Gaia DR3 1942384872124424704 2024.6865 0.32 0.22
Gaia DR3 1960631100090155008 2024.6838 0.11 0.10
Gaia DR3 1960631100090155264 2024.6838 0.27 0.20
Gaia DR3 1967282939282261120 2025.4455 0.22 0.15
Gaia DR3 1967283042361454848 2025.4455 0.13 0.11
Gaia DR3 1971147791380289536 2024.6837 0.23 0.16
Gaia DR3 1971148100617936128 2024.6837 0.00 0.00
Gaia DR3 1985752226366174720 2024.6865 0.49 0.33
Gaia DR3 1985752226366174848 2024.6865 0.25 0.17
Gaia DR3 1994221730075916800 2024.6759 0.88 0.58
Gaia DR3 1994221936234345344 2024.6759 1.35 1.07
Gaia DR3 200266181758708736 2024.8649 1.06 0.65
Gaia DR3 200266250476976768 2024.8649 0.36 0.25
The third component, DR3 1815165879238575232, Gaia DR3 2005665584549011328 2025.4400 0.21 0.15
is too far away, DM=4.3
The third component, DR3 2005665584530878848, Gaia DR3 2005665584549011328 2024.6867 0.33 0.23
is too far away, DM=4.4
The third component, DR3 2005665584530878848, Gaia DR3 2005665962510573824 2024.6867 0.40 0.28
is too far away, DM=4.4
Gaia DR3 2021978621937192832 2025.4371 1.04 0.60
Gaia DR3 2021979584009866368 2025.4371 0.42 0.28
Gaia DR3 2021979588354850176 2025.4371 2.05 1.73
Gaia DR3 2029432043779954432 2025.4400 0.83 0.45
Gaia DR3 2029433521248546304 2025.4400 0.10 0.09
Gaia DR3 2032121960300897280 2025.4398 1.10 0.61
Gaia DR3 2047012822409354752 2025.4399 0.52 0.32
Gaia DR3 2047014368597597696 2025.4399 0.45 0.29
Gaia DR3 2051915235490493952 2025.4399 0.30 0.22
Gaia DR3 2051915304209972992 2025.4399 0.30 0.21
Gaia DR3 2056444914159643008 2025.4426 0.49 0.31
Gaia DR3 2056444914159643136 2024.6781 0.50 0.33
Gaia DR3 2056444948519384064 2025.4426 0.51 0.33
Gaia DR3 2062485390533084544 2024.6780 0.30 0.20
Gaia DR3 2062485390533085568 2024.6780 0.32 0.23
Gaia DR3 2066763860841212672 2024.6783 1.48 1.12
Gaia DR3 2066763865140783104 2024.6783 2.93 2.51
Gaia DR3 2067274450853577344 2025.4455 0.70 0.38
Gaia DR3 2067274515272929536 2025.4427 0.29 0.20
Gaia DR3 2076871091425583232 2025.4399 1.29 0.88
Gaia DR3 2076871091425586944 2025.4399 0.91 0.51
Gaia DR3 2078886285778453632 2025.4400 0.73 0.41
Gaia DR3 2078886290079138560 2025.4400 0.59 0.36
Gaia DR3 2079604404307344128 2025.4372 0.20 0.14
Gaia DR3 2079604442969955968 2025.4372 0.82 0.49
Gaia DR3 2086713404118555520 2025.4426 0.37 0.26
Gaia DR3 2086713404118555648 2025.4426 0.28 0.20
Gaia DR3 2087586175833524480 2024.6780 0.35 0.25
Gaia DR3 2087586175833525376 2024.6780 0.23 0.14
Gaia DR3 2091495764301335424 2025.4452 0.11 0.10
Gaia DR3 2091495798661081344 2025.4452 0.13 0.11
Gaia DR3 2098996731404747776 2025.4399 0.53 0.33
Gaia DR3 2098996731404747904 2025.4399 1.47 1.02
Gaia DR3 2098996804424367360 2025.4399 0.37 0.25
Gaia DR3 2099859268216511488 2025.4399 1.14 0.77
Gaia DR3 2099859268216512000 2025.4399 1.41 0.90
Gaia DR3 2100011516217168640 2025.4399 1.46 1.03
Gaia DR3 2100011516217169920 2025.4399 1.49 1.00
Gaia DR3 2100451630105040256 2025.4399 1.26 0.83
Gaia DR3 2100451630105041152 2025.4399 0.38 0.26
Gaia DR3 2100501760964021504 2025.4399 0.37 0.25
Gaia DR3 2100502001482188032 2025.4399 0.41 0.28
Gaia DR3 2104998900960077056 2025.4398 0.00 0.00
Gaia DR3 2104999343337767424 2025.4398 0.60 0.35
Gaia DR3 211413206732904960 2024.8650 0.34 0.23
Gaia DR3 211413206732905600 2024.8650 0.37 0.26
Gaia DR3 2116748243555689728 2025.4397 0.48 0.32
Gaia DR3 2116748247853607168 2025.4397 0.26 0.18
Gaia DR3 2116892318236994816 2025.4398 0.22 0.16
This is a triple system. Gaia DR3 2119132916774654080 2025.4398 0.36 0.25
The third component has no Gaia identification;
[…]
This is a triple system. Gaia DR3 2119132916774654208 2025.4398 1.11 0.73
The third component has no Gaia identification;
[…]
Gaia DR3 2126777988630768000 2025.4400 0.75 0.44
Gaia DR3 2126777992924832128 2025.4400 1.00 0.65
Gaia DR3 2128304179486008704 2025.4426 0.37 0.26
Gaia DR3 2128304183783972992 2025.4426 0.28 0.20
Gaia DR3 2129540099271049216 2025.4426 0.82 0.46
Gaia DR3 2129540103571086592 2025.4426 0.99 0.56
Gaia DR3 2130903047014230528 2025.4426 0.63 0.37
Gaia DR3 2130903223113004032 2025.4426 0.33 0.23
Gaia DR3 2137630172689997824 2025.4372 0.31 0.21
Gaia DR3 2137630172690002176 2025.4372 0.29 0.20
Gaia DR3 2140767319879292160 2025.4372 0.39 0.28
Gaia DR3 2140767560397457536 2025.4453 0.29 0.21
Gaia DR3 2148617626742496512 2025.4371 0.29 0.20
Gaia DR3 2148617729821714688 2025.4371 0.43 0.29
Gaia DR3 2157622725756418816 2025.4398 0.29 0.21
Gaia DR3 2157623451608162688 2025.4398 0.88 0.47
Gaia DR3 2171870747199482496 2024.6864 1.03 0.59
Gaia DR3 2171870751495498752 2024.6864 0.67 0.38
Gaia DR3 2179261496856372224 2024.6864 0.93 0.58
Gaia DR3 2179261496856372352 2024.6864 0.92 0.56
Gaia DR3 2182397334073261568 2024.6782 1.03 0.62
Gaia DR3 2182397338377483264 2024.6782 0.84 0.47
Gaia DR3 2186439383632375296 2024.6863 2.40 1.77
Gaia DR3 2186439482412215808 2024.6863 0.62 0.40
Gaia DR3 219593745044391552 2024.8618 0.37 0.26
Gaia DR3 219605599154126976 2024.8618 0.49 0.32
Gaia DR3 2196841961921480960 2024.6864 0.86 0.53
Gaia DR3 2196841966222594944 2024.6864 1.75 1.40
Gaia DR3 2201661091331626752 2024.6867 0.25 0.18
Gaia DR3 2201661297490052096 2024.6867 0.24 0.17
Gaia DR3 2201834466266276864 2024.6867 0.23 0.16
Gaia DR3 2201834470572701952 2024.6867 0.28 0.20
Gaia DR3 222299432702445184 2024.8618 1.45 1.07
Gaia DR3 222299436999652096 2024.8618 1.48 1.13
Gaia DR3 2242465891977167872 2025.4426 0.24 0.17
Gaia DR3 2242465891977204992 2025.4426 0.43 0.30
Gaia DR3 229737942400057728 2024.8618 0.59 0.36
Gaia DR3 229737942400695424 2024.8618 1.75 1.35
Gaia DR3 2503402743097074304 2024.8618 1.01 0.58
Gaia DR3 2503402747391938688 2024.8618 0.83 0.45
Gaia DR3 2512615551025883136 2024.6868 0.95 0.54
Gaia DR3 2512615555321101696 2024.6868 0.31 0.22
Gaia DR3 2514529598906714880 2024.8618 0.52 0.33
Too weak Gaia DR3 2514541414361083648 2024.8618 >2.>2.. >2.>2.
Gaia DR3 2524864836409161728 2024.6868 0.36 0.26
Gaia DR3 2529474126591876096 2024.6815 0.41 0.28
Gaia DR3 2529661799483080192 2024.6815 0.29 0.21
Gaia DR3 2531509253895359744 2024.6868 0.23 0.17
Gaia DR3 2531509253895360000 2024.6868 0.29 0.21
Gaia DR3 2555905080453468544 2024.6815 0.30 0.21
Gaia DR3 2555905080453468800 2024.6815 0.26 0.19
Gaia DR3 2556421335522520576 2024.6816 0.29 0.21
Gaia DR3 2556421610401138176 2024.6816 0.44 0.30
Gaia DR3 2565584802867037696 2024.6868 0.19 0.13
Gaia DR3 2565584837226776576 2024.6868 0.21 0.14
Gaia DR3 2572437745309855488 2024.6869 0.23 0.14
Gaia DR3 2572437749605558272 2024.6869 0.35 0.24
Gaia DR3 2573278051366910336 2024.6869 0.16 0.12
Gaia DR3 2573278120086386432 2024.6869 0.16 0.12
Triple system, variable Gaia DR3 2581429521337662464 2024.6868 0.26 0.18
Triple system, variable Gaia DR3 2581429521338139264 2024.6868 0.28 0.20
Gaia DR3 2586004584926025600 2024.6869 0.30 0.21
Gaia DR3 2586004589220994560 2024.6869 0.28 0.20
Gaia DR3 2611384547405130752 2024.6840 0.34 0.24
Gaia DR3 2611384547405183488 2024.6840 0.32 0.23
Gaia DR3 2621082995876646400 2024.6840 0.31 0.22
Gaia DR3 2621082995876646528 2024.6840 0.37 0.26
Gaia DR3 2638615189818036608 2024.6840 0.13 0.11
Gaia DR3 2638615189818036736 2024.6840 0.13 0.11
Gaia DR3 2642922251741817344 2024.6840 0.51 0.32
Gaia DR3 2642922286101533696 2024.6840 0.56 0.34
Gaia DR3 266993755737826176 2024.8650 1.07 0.65
Gaia DR3 266993755740665600 2024.8650 1.09 0.68
Gaia DR3 2683565561622947584 2024.6839 0.48 0.32
Gaia DR3 2683565630342923520 2024.6839 0.24 0.17
Gaia DR3 2688137537130343936 2024.6838 0.17 0.12
Gaia DR3 2688137537130451456 2024.6838 0.22 0.16
Gaia DR3 2722918766408972288 2024.6839 0.34 0.24
Gaia DR3 2722918796473758720 2024.6839 0.38 0.26
Gaia DR3 2725049173267026560 2024.6838 0.17 0.12
Gaia DR3 2725049173267027968 2024.6838 0.29 0.21
Gaia DR3 2731572610114153856 2024.6840 0.19 0.14
Gaia DR3 2731572610114154112 2024.6840 0.24 0.18
Gaia DR3 2738818666619511808 2024.6868 0.32 0.22
Gaia DR3 2738818666619512064 2024.6868 0.35 0.24
Gaia DR3 2739496687336986112 2024.6841 0.20 0.14
Gaia DR3 2739498199165473792 2024.6841 0.25 0.18
Gaia DR3 2741770924059034624 2024.6814 0.48 0.30
Gaia DR3 2741770958420224000 2024.6814 0.21 0.15
Gaia DR3 2743364563085132800 2024.6841 1.73 1.42
Gaia DR3 2743364563085132928 2024.6841 1.77 1.44
Gaia DR3 2752042939643757952 2024.6868 0.23 0.16
Gaia DR3 2752042939643758080 2024.6868 0.22 0.15
Gaia DR3 2763789469038658432 2024.6841 0.23 0.16
Gaia DR3 2763789469039961472 2024.6841 0.17 0.12
Gaia DR3 2799550942998190592 2024.6815 0.00 0.00
Gaia DR3 2799550947293531648 2024.6814 0.42 0.28
Gaia DR3 280403536989156224 2024.8650 0.32 0.23
Too weak Gaia DR3 280403541287570176 2024.8650 >2.>2. >2.>2.
Gaia DR3 2813290096701231104 2024.6840 0.77 0.39
Gaia DR3 2813290100998398464 2024.6840 0.75 0.40
Gaia DR3 2833895636096202752 2024.6840 0.32 0.22
Gaia DR3 2833895636096692224 2024.6757 0.28 0.21
Gaia DR3 2836313908842872704 2024.6866 0.29 0.19
Gaia DR3 2836313908842872832 2024.6866 0.25 0.17
Gaia DR3 2837991381336616320 2024.6867 0.45 0.31
Gaia DR3 2837991454349643136 2024.6867 0.45 0.30
Gaia DR3 2847357575430420992 2024.6815 0.26 0.18
Gaia DR3 2847357575430421120 2024.6815 0.65 0.38
Gaia DR3 2849031405789315328 2024.6814 0.20 0.13
Gaia DR3 2849031410085904640 2024.6814 0.33 0.23
Gaia DR3 2860343426229962752 2024.6814 0.24 0.16
Gaia DR3 2865611564396936192 2024.6866 0.29 0.21
Gaia DR3 2865611564396936320 2024.6866 0.35 0.24
Gaia DR3 2868784063464098048 2024.6866 0.78 0.49
Gaia DR3 2868784136478476032 2024.6866 0.20 0.13
Gaia DR3 2874196929143032192 2024.6815 1.22 0.83
Gaia DR3 2874196929143032448 2024.6815 1.20 0.77
Gaia DR3 2879160021551346176 2024.6866 0.52 0.34
Gaia DR3 2879160021551346688 2024.6866 0.43 0.29
Gaia DR3 293553975230799104 2024.6869 0.31 0.21
Gaia DR3 293554074014451200 2024.6869 0.21 0.13
Gaia DR3 2989629982019538304 2024.8621 2.52 1.85
Gaia DR3 2989630016379276160 2024.8621 2.23 1.86
Gaia DR3 3002343871195786496 2024.8622 0.38 0.27
Gaia DR3 3002531548383047296 2024.8622 0.37 0.26
Gaia DR3 3015129546453825920 2024.8622 1.38 0.95
Gaia DR3 3015130405447132288 2024.8621 1.34 0.89
Gaia DR3 3033070586928188416 2024.8624 0.47 0.32
Gaia DR3 3033070655647665536 2024.8624 1.47 1.16
Gaia DR3 3033419235197542912 2024.8624 0.74 0.43
Gaia DR3 3033442909056369792 2024.8624 0.63 0.39
Gaia DR3 3033693700785580800 2024.8624 0.95 0.56
Too weak Gaia DR3 3033693700785585152 2024.8624 >2.>2. >2.>2.
Gaia DR3 3035042522369773824 2024.8624 1.27 0.87
Gaia DR3 3035042526672692096 2024.8624 1.12 0.83
Gaia DR3 3042299715729341056 2024.8624 1.03 0.63
Gaia DR3 3042300093686461312 2024.8624 0.95 0.57
Gaia DR3 3046204150242468480 2025.2173 0.58 0.37
Gaia DR3 3046204180298475264 2025.2173 0.43 0.30
Gaia DR3 3048356684771809536 2024.8624 0.37 0.27
Too weak Gaia DR3 3048357539467185280 2024.8624 >2.>2. >2.>2.
Gaia DR3 3061488839331542912 2024.8624 0.90 0.52
Gaia DR3 3061488839334653696 2024.8624 0.91 0.54
Gaia DR3 3075884367113894400 2024.8625 1.57 1.30
Gaia DR3 3075884367113894528 2024.8625 1.22 0.83
Gaia DR3 3079739250589041024 2024.8625 1.70 1.30
Gaia DR3 3079739254881696256 2024.8625 1.18 0.68
Gaia DR3 3084652044636398976 2024.8625 0.35 0.25
Gaia DR3 3084652044636399104 2024.8625 0.38 0.27
Gaia DR3 3103862814877943296 2024.8622 0.38 0.27
Gaia DR3 3103862814877944064 2025.2174 0.45 0.31
Gaia DR3 3107726150841412224 2024.8624 1.54 1.31
Gaia DR3 3107726155141310080 2024.8624 1.52 1.25
Gaia DR3 3155908468358227968 2024.8624 1.20 0.85
Gaia DR3 3155908575735398400 2024.8624 1.01 0.67
Gaia DR3 3158878734598549376 2025.2173 0.48 0.31
Gaia DR3 3158878734598549504 2024.8623 0.63 0.37
Gaia DR3 3158926322836178816 2025.2173 0.42 0.28
Gaia DR3 3158926322836178944 2024.8623 0.95 0.49
Gaia DR3 3170300942420466176 2025.2173 0.28 0.20
Gaia DR3 3170303755624614528 2024.8623 0.48 0.31
Gaia DR3 3170303759920420224 2024.8623 0.33 0.23
Gaia DR3 3170394607068638336 2025.2173 0.30 0.22
Gaia DR3 3174712286151938944 2024.8621 1.19 0.80
Gaia DR3 3174712354871414912 2024.8621 0.97 0.58
Gaia DR3 3194331181363949312 2024.8619 0.96 0.55
Gaia DR3 3194331284443164288 2024.8619 0.86 0.48
Too weak Gaia DR3 3194720786437148032 2024.8619 >2.>2. >2.>2.
Gaia DR3 3194720889516362880 2024.8619 1.08 0.72
Gaia DR3 3201928153876092416 2024.8620 1.78 1.46
Gaia DR3 3201928153876683648 2024.8620 1.97 1.60
Gaia DR3 3208097204381858304 2024.8622 0.78 0.43
Gaia DR3 3208097582339739136 2024.8622 0.71 0.40
Gaia DR3 3208920321978859008 2024.8622 0.43 0.30
Gaia DR3 3208920326274769664 2024.8622 1.48 1.07
Too weak Gaia DR3 3211638593895777792 2024.8622 >2.>2. >2.>2.
Gaia DR3 3211638593896898944 2024.8622 0.42 0.28
Gaia DR3 3213811602532945152 2024.8622 1.48 1.13
Gaia DR3 3213811641189319680 2024.8622 1.52 1.14
Gaia DR3 3216741186843662720 2024.8622 2.00 1.60
Gaia DR3 3216741186843662848 2024.8622 2.88 2.44
Too weak Gaia DR3 3223175082210855936 2024.8622 >2.>2. >2.>2.
Gaia DR3 3223175082210856192 2024.8622 1.37 0.88
Gaia DR3 3233274852427241216 2024.8620 0.51 0.33
Too weak Gaia DR3 3233274955506455936 2024.8620 >2.>2. >2.>2.
Gaia DR3 3237157051889659904 2024.8622 0.70 0.40
Gaia DR3 3237157159265190400 2024.8622 0.83 0.46
Gaia DR3 3239798461074336512 2024.8649 0.68 0.39
Gaia DR3 3239798461074336640 2024.8649 0.79 0.43
Gaia DR3 3242368702645853184 2024.8649 0.83 0.45
Gaia DR3 3242368706941778688 2024.8649 0.77 0.43
Gaia DR3 3251884739561945472 2024.8620 2.90 2.34
Gaia DR3 3251884739561945600 2024.8620 0.93 0.54
Gaia DR3 3258912371210481152 2024.8620 0.72 0.41
Gaia DR3 3258912439929957376 2024.8620 1.19 0.83
Gaia DR3 3266980170921153920 2024.8618 1.18 0.74
Too weak Gaia DR3 3266980243936341248 2024.8618 >2.>2. >2.>2.
Gaia DR3 3270823243234194432 2024.8618 1.30 0.93
Gaia DR3 3270823243234194944 2024.8618 1.09 0.73
Gaia DR3 3279804084274429440 2024.8620 2.52 2.10
Gaia DR3 3279804844484336128 2024.8620 0.61 0.36
Gaia DR3 3285744612456150016 2024.8620 0.32 0.23
Gaia DR3 3285755951169814528 2024.8620 0.42 0.28
Gaia DR3 3288572968680438528 2024.8621 0.22 0.15
Gaia DR3 3288572968680438912 2024.8621 0.20 0.14
Gaia DR3 3290322734060874752 2024.8649 0.63 0.37
Gaia DR3 3290322738355773696 2024.8649 0.81 0.42
Too weak Gaia DR3 3290540643520088448 2024.8649 >2.>2. >2.>2.
Gaia DR3 3290540746599303296 2024.8649 0.63 0.37
Gaia DR3 3308694065827291648 2024.8621 0.47 0.31
Gaia DR3 3308694065827292160 2024.8621 0.71 0.39
Gaia DR3 3338164860103556352 2024.8649 0.40 0.28
Gaia DR3 3338164860104174592 2024.8649 1.23 0.73
Gaia DR3 3344333085975753216 2024.8623 0.63 0.37
Gaia DR3 3344333292134182400 2024.8623 1.06 0.62
Gaia DR3 3346799805954385408 2024.8649 0.48 0.31
Gaia DR3 3346799805954387456 2024.8649 0.19 0.14
Gaia DR3 3360175090027418496 2024.8624 0.44 0.29
Gaia DR3 3360175090027418880 2024.8624 1.80 1.48
Gaia DR3 3365948002813451648 2024.8623 1.40 1.00
Gaia DR3 3365948002813451776 2024.8623 1.45 1.08
Gaia DR3 3371529368651025664 2024.1278 0.40 0.28
Gaia DR3 3371529467432172672 2024.1279 0.48 0.33
Gaia DR3 3399607116051404032 2024.8649 0.59 0.36
Too weak Gaia DR3 3399608589223938816 2024.8649 >2.>2. >2.>2.
Gaia DR3 3400637667683964544 2024.8649 0.55 0.34
Gaia DR3 3400637702043702656 2024.8649 0.41 0.27
This is a triple system. Gaia DR3 3402090259984528768 2024.1278 0.40 0.28
The third component is DR3 3402090466140560128
This is a triple system. Gaia DR3 3402090466142958464 2024.1278 0.51 0.33
The third component is DR3 3402090466140560128
Gaia DR3 3410692942038997120 2024.8621 0.39 0.27
Gaia DR3 3410692942038997760 2024.8621 0.43 0.29
Gaia DR3 3417608762800814080 2024.8650 0.32 0.23
Gaia DR3 3417608767098501120 2024.8650 0.37 0.26
Gaia DR3 3418915433586276096 2024.8649 0.33 0.24
Gaia DR3 3418915433588729856 2024.8649 1.06 0.56
Gaia DR3 3451266742171824640 2024.8650 0.20 0.15
Gaia DR3 3451267120128948992 2024.1278 0.19 0.15
Gaia DR3 3593154766362734336 2025.2179 0.31 0.23
Gaia DR3 3593154766362734464 2025.2179 0.31 0.23
Gaia DR3 3609320267350909696 2025.4392 0.47 0.32
Gaia DR3 3609320503573212416 2025.4392 0.53 0.36
Gaia DR3 3609413309226560768 2025.4392 1.15 0.85
Gaia DR3 3609413485320886528 2025.4392 0.35 0.26
Gaia DR3 3618102204160590080 2025.4446 0.45 0.32
Gaia DR3 3618102204160590208 2025.2181 0.27 0.20
Gaia DR3 3627968323859543680 2025.4446 0.34 0.26
Gaia DR3 3627968328155313024 2025.4446 1.40 1.06
Gaia DR3 3629449713915122688 2025.4446 0.25 0.19
Gaia DR3 3629449713915123072 2025.4446 1.00 0.65
Gaia DR3 3636691509812473856 2025.4392 0.26 0.18
Gaia DR3 3636714977513779712 2025.4392 0.31 0.22
Gaia DR3 3639520621950395776 2025.4447 0.12 0.10
Gaia DR3 3639520621950395904 2025.4447 0.12 0.10
The third component has no Gaia identification; Gaia DR3 3641315024925921024 2025.4393 0.27 0.20
its separation from DR3 3641315024925921024 is 0.482
The third component has no Gaia identification; Gaia DR3 3641315024927514880 2025.4393 0.69 0.40
its separation from DR3 3641315024925921024 is 0.482
Gaia DR3 3666150003300751744 2025.4448 0.22 0.16
Gaia DR3 3666150003300991744 2025.4448 0.25 0.18
Gaia DR3 367020421620737536 2024.6815 0.28 0.20
Gaia DR3 367020421621893120 2024.6815 0.26 0.19
Gaia DR3 3677417214346066048 2025.2180 0.34 0.25
Gaia DR3 3677417214346902144 2025.2180 0.58 0.37
Gaia DR3 3680089852236243584 2025.2180 0.23 0.17
Gaia DR3 3680089852236243712 2025.2180 0.25 0.19
Gaia DR3 3686259727375524864 2025.4392 0.43 0.29
Gaia DR3 3686261307923490688 2025.4392 0.25 0.18
Gaia DR3 3717756234385505792 2025.4446 0.19 0.14
Gaia DR3 3717756337464721152 2025.4364 0.76 0.50
Gaia DR3 3718581417862169344 2025.4364 0.35 0.26
Gaia DR3 3718581520941384320 2025.4447 0.41 0.29
Gaia DR3 372066046121799168 2024.6870 0.00 0.00
Gaia DR3 372066076185339648 2024.6869 0.26 0.18
Gaia DR3 3724521696934640384 2025.4447 0.20 0.14
Gaia DR3 3724521701229940096 2025.4447 0.34 0.24
Gaia DR3 3738783878870503424 2025.4365 0.34 0.25
Gaia DR3 3738783878870503552 2025.4365 0.77 0.53
Gaia DR3 3739054324371407232 2025.4447 0.43 0.31
Gaia DR3 3739054324371407488 2025.4447 0.24 0.18
Gaia DR3 374529119673672576 2024.6870 0.38 0.25
Gaia DR3 374529123966850944 2024.6870 0.29 0.20
Gaia DR3 3768566346037170304 2025.2178 0.34 0.25
Gaia DR3 3768566350332212352 2025.2178 0.34 0.25
Gaia DR3 377244848968006016 2024.6816 0.26 0.18
Gaia DR3 377245055126435072 2024.6816 0.22 0.15
Gaia DR3 3792739899447945216 2024.1285 0.73 0.52
Gaia DR3 3792740006822516608 2024.1285 1.10 0.89
Gaia DR3 3793106419072038272 2025.2178 0.26 0.19
Gaia DR3 3793107930900527616 2025.2178 0.20 0.14
Gaia DR3 3838561363635304448 2024.8679 0.35 0.25
Gaia DR3 3838561363635829632 2024.8679 0.34 0.24
Gaia DR3 3841514200895729152 2024.8679 0.37 0.24
Gaia DR3 3841514200898362368 2024.8679 0.33 0.23
Gaia DR3 3890860179670959104 2025.2178 0.25 0.18
Gaia DR3 3890860183966486656 2025.2178 0.24 0.17
The third component is DR3 3895404602964088448. Gaia DR3 3895404602962524544 2025.2179 0.20 0.15
No parallax information.
The third component is DR3 3895404602964088448. Gaia DR3 3895404602964117504 2025.2179 0.32 0.23
No parallax information.
Gaia DR3 3907465249087614976 2025.2179 0.50 0.32
Too weak Gaia DR3 3907466035066193664 2024.1286 >2.>2. >2.>2.
Gaia DR3 3930274446007361664 2025.4365 0.21 0.16
Gaia DR3 3930274613510165504 2025.4365 0.17 0.13
This is a multi-component system. Gaia DR3 3936935283153115008 2025.4419 0.24 0.16
One of the components is DR3 3936935287448996864.
This is a multi-component system. Gaia DR3 3936947278997113216 2025.4419 1.03 0.81
One of the components is DR3 3936935287448996864.
Gaia DR3 3945118265299248128 2025.4447 0.14 0.11
Gaia DR3 3945118643256370688 2025.4447 0.16 0.12
Gaia DR3 3974007108685133056 2024.1285 1.00 0.70
Gaia DR3 3974007585425593472 2025.2179 0.68 0.40
Gaia DR3 3975566491051012480 2025.2179 0.25 0.18
Gaia DR3 3975566491051012608 2025.2179 0.25 0.18
Gaia DR3 4007902062872083456 2024.1285 0.41 0.30
Gaia DR3 4007902269030513920 2024.1285 0.95 0.70
Gaia DR3 4025108384759142272 2025.2179 0.60 0.37
Gaia DR3 405028575795255808 2024.6870 0.26 0.19
Gaia DR3 405028683170219392 2024.6870 0.31 0.22
Gaia DR3 4103210548364197376 2025.4452 0.23 0.17
Gaia DR3 4103210582723957120 2025.4452 0.31 0.24
Gaia DR3 418549201560407040 2024.6815 0.45 0.32
Gaia DR3 418549205864071296 2024.6815 0.43 0.30
Gaia DR3 4197023904013455872 2025.4371 0.19 0.13
Gaia DR3 4197023904014523776 2025.4371 0.16 0.12
Gaia DR3 4218533748765026560 2025.4373 0.32 0.23
Gaia DR3 4218533959216594176 2025.4373 0.77 0.44
Gaia DR3 4230699329529382400 2025.4373 0.28 0.20
Gaia DR3 4230699363889120128 2025.4373 0.18 0.13
Gaia DR3 4231853855392895104 2024.6783 0.94 0.48
Gaia DR3 4231853855392895360 2024.6783 1.24 0.70
Gaia DR3 4232150409999693056 2024.6783 0.42 0.29
Gaia DR3 4232150410001183872 2024.6783 0.42 0.29
Gaia DR3 4243504314963260672 2024.6780 0.36 0.25
Gaia DR3 4243504314963261696 2024.6780 1.10 0.68
Gaia DR3 4245247040891396480 2024.6782 1.53 1.23
Gaia DR3 4245247247049828480 2024.6782 0.41 0.29
Gaia DR3 4249024035142729728 2024.6782 1.19 0.74
Gaia DR3 4249024035142730368 2024.6782 0.38 0.26
Gaia DR3 428811394564909568 2024.6814 0.24 0.17
Gaia DR3 428811428924644352 2024.6814 0.32 0.23
Gaia DR3 4292414749731991808 2025.4371 0.31 0.21
Gaia DR3 4292415883603395456 2025.4371 0.46 0.30
Gaia DR3 4335277282059878656 2025.4343 0.53 0.36
Gaia DR3 4335277282066850048 2025.4343 0.56 0.37
Gaia DR3 4350763323519241600 2025.4343 0.63 0.39
Gaia DR3 4350763323519241984 2025.4343 0.80 0.45
Gaia DR3 4353764371788098944 2025.4369 0.36 0.26
Gaia DR3 4353764436210206080 2025.4369 0.26 0.19
Gaia DR3 4368065856969737728 2025.4396 0.20 0.15
Gaia DR3 4368065921391001600 2025.4396 0.77 0.42
Gaia DR3 4385319393432940544 2025.4395 0.81 0.44
Gaia DR3 4385319427793947648 2025.4395 0.32 0.22
Gaia DR3 4390274857979921408 2025.4396 0.47 0.31
Gaia DR3 4390274857979922432 2025.4396 0.26 0.19
Gaia DR3 4395522998779760128 2025.4367 0.29 0.21
Gaia DR3 4395523033138822656 2025.4367 0.29 0.21
Gaia DR3 4404091664690806016 2025.4340 0.23 0.17
Gaia DR3 4404091664691143168 2025.4340 0.28 0.20
Gaia DR3 4407541313703639424 2025.4369 1.07 0.69
Gaia DR3 4407544268641139712 2025.4369 1.12 0.74
Gaia DR3 441620979895171840 2024.8619 0.30 0.22
Gaia DR3 441620979901788544 2024.8619 0.64 0.39
Gaia DR3 4430185034123000960 2025.4341 0.26 0.19
Gaia DR3 4430185068482324864 2025.4367 0.24 0.17
Gaia DR3 4435683451255623808 2025.4395 0.24 0.17
Gaia DR3 4435689739087756800 2025.4395 0.11 0.10
Gaia DR3 4441920942064076288 2025.4395 0.47 0.31
Gaia DR3 4441920946360132096 2025.4395 0.37 0.25
Gaia DR3 4445849153513150080 2025.4450 0.25 0.18
Gaia DR3 4445849157808061312 2025.4450 0.25 0.18
Gaia DR3 4451719690908056192 2025.4341 0.28 0.20
Gaia DR3 4451731819895701376 2025.4341 1.02 0.64
Gaia DR3 4454413455738305792 2025.4341 0.60 0.38
Gaia DR3 4454413460037125888 2025.4341 0.62 0.38
Gaia DR3 4466613022581649152 2025.4343 0.78 0.48
Gaia DR3 4466613022581649664 2025.4343 0.64 0.41
Gaia DR3 4471562916556377856 2025.4397 0.27 0.20
Gaia DR3 4471658131695403264 2025.4397 0.77 0.41
Gaia DR3 4474801218817508864 2025.4370 0.72 0.41
Gaia DR3 4474801257476653952 2025.4370 0.34 0.24
The third component is DR3 4493544082437142784 , Gaia DR3 4493544082437142400 2025.4370 0.84 0.49
no parallax data available.
The third component is DR3 4493544082437142784 , Gaia DR3 4493544185516358656 2025.4370 0.84 0.47
no parallax data available.
Gaia DR3 4502416488439872128 2025.4370 0.24 0.17
Gaia DR3 4502416488439872256 2025.4370 0.26 0.18
Gaia DR3 4513425898394508288 2025.4398 0.31 0.22
Gaia DR3 4513425932752582144 2025.4398 0.42 0.28
Gaia DR3 4513771449965175296 2025.4398 0.60 0.36
Gaia DR3 4513771832248074240 2025.4398 1.31 0.92
Gaia DR3 4541750383049979904 2025.4396 0.38 0.26
Gaia DR3 4541753303627759104 2025.4396 0.24 0.17
Gaia DR3 4544016820111855360 2025.4370 0.20 0.14
Gaia DR3 4544017129349496832 2025.4370 0.32 0.22
Gaia DR3 4558397710654524288 2025.4344 0.26 0.20
Gaia DR3 4558397710655578368 2025.4344 0.38 0.28
Gaia DR3 4571298692894669824 2025.4344 0.77 0.44
Gaia DR3 4571298692894670336 2025.4344 0.91 0.50
Gaia DR3 4572807566444752128 2025.4395 0.32 0.22
Gaia DR3 4577271961615916032 2025.4370 0.33 0.23
Gaia DR3 4577272270853565952 2025.4370 0.21 0.15
Gaia DR3 4584470704757139968 2025.4397 0.46 0.30
Gaia DR3 4584470739116876800 2025.4397 0.35 0.23
Gaia DR3 4589844945796230016 2025.4452 0.17 0.12
Gaia DR3 4589850065397245824 2025.4452 0.30 0.21
Gaia DR3 459661938487951360 2024.8617 0.49 0.32
Gaia DR3 459661938487952000 2024.8617 0.98 0.56
Gaia DR3 4599117333509619200 2025.4397 1.12 0.63
Gaia DR3 4599118119486450688 2025.4397 0.70 0.37
Gaia DR3 4599984504586131456 2025.4397 0.23 0.17
Gaia DR3 4599984642025088128 2025.4397 0.14 0.11
Gaia DR3 4600068239268380928 2025.4370 0.00 0.00
Gaia DR3 4600068239268381312 2025.4370 0.27 0.19
Gaia DR3 460621193663868160 2024.8617 0.34 0.25
Gaia DR3 460621296743080960 2024.8617 1.24 0.84
Gaia DR3 4724021644644224 2024.8617 1.23 0.86
Gaia DR3 4725155516009856 2024.8617 1.15 0.77
Gaia DR3 5110662129034491904 2024.8619 1.02 0.62
Gaia DR3 5110662129034492032 2024.8619 0.61 0.37
Gaia DR3 5114544745110388736 2024.8619 0.70 0.41
Gaia DR3 5114547700047887360 2024.8619 0.88 0.50
Gaia DR3 5166931251492281472 2024.8619 1.24 0.89
Too weak Gaia DR3 5166931595089502208 2024.8619 >2.>2. >2.>2.
Gaia DR3 5169322066511702016 2024.8619 0.51 0.33
Gaia DR3 5169322070807079552 2024.8619 0.77 0.45
Too weak Gaia DR3 5181911234131665920 2024.8619 >2.>2. >2.>2.
Gaia DR3 5181911238426924288 2024.8619 0.42 0.29
Gaia DR3 53720557585606912 2024.8621 1.14 0.73
Gaia DR3 53720561882110336 2024.8621 1.04 0.62
Gaia DR3 5737419559713009024 2024.8625 2.01 1.69
Gaia DR3 5737419559713009152 2024.8625 1.11 0.69
Gaia DR3 5741344919302959104 2024.8679 0.24 0.18
Gaia DR3 5741344919302959360 2025.2123 0.29 0.23
Gaia DR3 5741345125461458944 2024.8679 0.25 0.19
Gaia DR3 5741345125461459200 2025.2123 0.26 0.20
Gaia DR3 5755402484702384128 2024.8625 0.42 0.29
Gaia DR3 5755402720924702720 2024.8625 0.76 0.43
Gaia DR3 5760475695776013568 2024.8625 1.44 1.02
Gaia DR3 5760475695777562880 2024.8625 1.37 1.00
Gaia DR3 579529081235551360 2024.8679 0.19 0.12
Gaia DR3 579529081235551488 2024.8679 0.18 0.12
Gaia DR3 594453852070106624 2024.8679 0.19 0.11
Gaia DR3 594453852070765696 2025.2176 0.25 0.18
Gaia DR3 601946577137602816 2024.8625 2.29 1.90
Gaia DR3 601946955094723968 2024.8625 0.77 0.43
Gaia DR3 619889232512875648 2024.8679 0.25 0.18
Gaia DR3 619890709981625344 2024.8679 0.28 0.20
Gaia DR3 6321593106912540288 2025.4340 0.53 0.36
Gaia DR3 6321593111208064384 2025.4340 0.42 0.30
Gaia DR3 6321929729270985472 2025.4339 0.24 0.19
Gaia DR3 6321929733564266752 2025.4339 0.18 0.12
Gaia DR3 633627316769035264 2024.8679 0.40 0.27
Gaia DR3 633627321064388096 2024.8679 0.33 0.23
Gaia DR3 639517366199833344 2024.8679 0.30 0.21
Gaia DR3 639517366199833472 2024.8679 0.27 0.19
Gaia DR3 644549800855341184 2024.8680 0.25 0.18
Too weak Gaia DR3 644549869574817792 2024.8680 >2.>2. >2.>2.
Gaia DR3 644936691509435520 2024.8680 0.33 0.23
Gaia DR3 644936691509435648 2025.2177 0.37 0.26
Gaia DR3 648508863053616384 2025.2177 0.28 0.20
Gaia DR3 648508867348963840 2024.8680 0.60 0.36
Too weak Gaia DR3 656657519823458688 2024.8652 >2.>2. >2.>2.
Gaia DR3 656657622902671488 2024.8652 0.66 0.37
Gaia DR3 661127824863058176 2024.1279 0.24 0.18
Gaia DR3 661128752575993600 2024.8652 0.28 0.20
Too weak Gaia DR3 668946623847575936 2024.8652 >2.>2. >2.>2.
Gaia DR3 668947001804697472 2024.8652 0.45 0.30
Gaia DR3 675639449909479040 2024.8652 1.28 0.83
Gaia DR3 675639449911182336 2024.8652 1.30 0.86
Gaia DR3 680661782802091392 2025.2175 0.58 0.34
Gaia DR3 680662573076074752 2024.8652 0.32 0.22
Gaia DR3 6840365615137220608 2024.6838 0.29 0.22
Gaia DR3 6840365718216434688 2024.6838 0.22 0.16
Gaia DR3 6903444700303516160 2024.6782 0.46 0.32
Gaia DR3 6903444700303516800 2024.6782 0.95 0.50
Gaia DR3 6906039925703317248 2024.6781 0.29 0.21
Gaia DR3 6906039925703317376 2024.6781 0.29 0.21
Gaia DR3 6909250362214709248 2024.6839 0.38 0.27
Gaia DR3 6909250362214915072 2024.6839 0.35 0.25
Gaia DR3 6912178464759019520 2024.6839 0.38 0.27
Gaia DR3 6912178464759020544 2024.6839 0.45 0.32
Gaia DR3 695440421670598400 2024.8680 0.31 0.22
Gaia DR3 695440421670598656 2024.8680 0.32 0.21
Gaia DR3 710817744619067520 2024.8652 1.50 1.24
Gaia DR3 710817744619067648 2024.8652 0.40 0.26
Gaia DR3 758958211973432704 2025.2179 0.26 0.19
Gaia DR3 758958929232265472 2025.2179 0.43 0.29
Gaia DR3 77161217776670208 2024.8617 0.30 0.21
Gaia DR3 77161222072044288 2024.8617 0.28 0.19
Gaia DR3 787833483266126080 2025.2123 0.34 0.25
Gaia DR3 787833551986183168 2025.2123 0.31 0.22
Gaia DR3 796311542548505984 2024.8680 0.27 0.19
Gaia DR3 796311954865365888 2024.1279 1.19 0.89
Gaia DR3 803485581043936768 2025.2123 0.41 0.28
Gaia DR3 803485615403673728 2025.2123 0.30 0.22
Gaia DR3 817937454502566528 2024.8680 0.41 0.28
Gaia DR3 817937458796755200 2024.8680 0.53 0.34
Gaia DR3 820692212165053056 2024.8680 0.19 0.12
Gaia DR3 820692216460834944 2025.2177 0.20 0.14
Gaia DR3 853328573431352576 2025.2178 0.39 0.28
Gaia DR3 853328573431352704 2025.2178 0.32 0.24
Gaia DR3 87111458705996800 2024.8617 0.97 0.58
Too weak Gaia DR3 87111905382846976 2024.8617 >2.>2. >2.>2.
Gaia DR3 883478861596266880 2024.8623 0.75 0.41
Gaia DR3 883478865892528128 2024.8623 1.08 0.69
Gaia DR3 898728439334815616 2024.8650 0.58 0.35
Gaia DR3 898728439334815744 2024.8650 0.39 0.26
Gaia DR3 899627393169280384 2024.8651 0.42 0.26
Gaia DR3 899627393169280512 2024.8651 0.44 0.29
Gaia DR3 904299046277136768 2024.8652 1.44 0.99
Gaia DR3 904299046277137024 2024.8652 0.54 0.33
Gaia DR3 914241517609344128 2024.8652 0.26 0.17
Gaia DR3 914244399532441472 2024.8652 0.25 0.17
Gaia DR3 916482150508905856 2024.8652 0.19 0.12
Gaia DR3 917982605923642752 2024.8652 0.24 0.16
Gaia DR3 922595430863222144 2024.1279 0.87 0.59
Gaia DR3 922601585552127104 2024.1279 0.28 0.21
Gaia DR3 930712128780483840 2024.8652 1.68 1.36
Gaia DR3 930712133074729856 2024.8652 1.63 1.25
Gaia DR3 937667952870488704 2024.8623 0.23 0.16
Gaia DR3 937667952871051776 2025.2174 0.23 0.17
Gaia DR3 951148377744097024 2024.8651 0.38 0.26
Gaia DR3 951148446463574016 2024.8651 0.35 0.24
Gaia DR3 951621305182882432 2024.8650 2.64 2.10
Gaia DR3 951621305182882944 2024.8650 1.87 1.37
Gaia DR3 953178312430077312 2024.8651 1.54 1.16
Gaia DR3 953178312430645376 2024.8651 0.55 0.34
Too weak Gaia DR3 982674365408534400 2024.8651 >2.>2. >2.>2.
Gaia DR3 982674434128011904 2024.8651 0.54 0.35
Gaia DR3 98692339803443328 2024.6869 0.23 0.15
Gaia DR3 98692614681349248 2024.6869 0.14 0.11
Gaia DR3 990524191236435200 2024.8651 0.48 0.32
Gaia DR3 990524191236435456 2024.8651 0.39 0.27
Gaia DR3 992789150829982848 2024.1278 0.28 0.21
Gaia DR3 992789872384490496 2024.8650 0.29 0.21